### v1

# More Modality, More AI: Exploring Design Opportunities of AI-Based Multi-modal Remote Monitoring Technologies for Early Detection of Mental Health Sequelae in Youth Concussion Patients

Abstract.
Anxiety, depression, and suicidality are common mental health sequelae following concussion in youth patients, often exacerbating concussion symptoms and prolonging recovery. Despite the critical need for early detection of these mental health symptoms, clinicians often face challenges in accurately collecting patients’ mental health data and making clinical decision-making in a timely manner. Today’s remote patient monitoring (RPM) technologies offer opportunities to objectively monitor patients’ activities, but they were not specifically designed for youth concussion patients; moreover, the large amount of data collected by RPM technologies may also impose significant workloads on clinicians to keep up with and use the data. To address these gaps, we employed a three-stage study consisting of a formative study, interface design, and design evaluation. We first conducted a formative study through semi-structured interviews with six highly professional concussion clinicians and identified clinicians’ key challenges in remotely collecting patient information and accessing patient treatment compliance. Subsequently, we proposed preliminary clinician-facing interface designs with the integration of AI-based RPM technologies (AI-RPM), followed by design evaluation sessions with highly professional concussion clinicians. Clinicians underscored the value of integrating multi-modal AI-RPM technologies to support clinicians’ decision-making while emphasizing the importance of customizable interfaces with explainability and multiple responsible design considerations.
Youth concussion, mental health sequelae, clinical decision-making, remote patient monitoring, artificial intelligence
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing Human computer interaction (HCI)$c$s: Human-centered computing HCI design and evaluation methodsccs: Human-centered computing Collaborative and social computing theory, concepts and paradigms
1. Introduction
[…]
The standard of care (SoC) approaches of concussion clinicians include (1) asking patients to recall their at-home behaviors during clinical visits (Silverberg et al., 2020; Brent and Max, 2017) and (2) instructing patients or parents to complete self-evaluation questionnaires at home, such as Generalized Anxiety Disorder-7 (GAD-7) and Patient Health Questionnaire-9 (PHQ-9) (Carson et al., 2021; Silverberg et al., 2020). However, these approaches are not without limitations: it could be challenging for concussion patients to comprehensively and accurately recall their previous behaviors in a weekly clinic visit (Frissa et al., 2016; Van den Bergh and Walentynowicz, 2016); patients may be able to complete questionnaires on a regular basis at home, but concussion clinicians can only access the questionnaire results when patients return to the clinics in person, which still delays clinicians’ decision-making.
Previous work has proposed different Remote Patient Monitoring (RPM) technologies to collect patients’ at-home information in various scenarios (Chung et al., 2016; Olen and Dechert-Crooks, 2017; Barnes et al., 2024; van Ierssel et al., 2023; Toresdahl et al., 2021; Tirosh et al., 2024) For example, mobile health (mHealth) applications have been proposed in which patients can report their physiological measurements to care provider teams in real-time. Wearable technologies, such as the ActiGraph (Yang et al., 2020) and Philips Actiwatch (Donahue et al., 2024), have been explored to measure the daily steps and sleep patterns of adult concussion patients. Despite their relative success in capturing patients’ activity data, wearable-based RPM approaches are not without limitations: mHealth applications could lead to cognitive burdens for patients with low digital literacy and brain injuries (Giebel et al., 2024; Purdy, 2023); extensive amount of time and effort is required to possess the massive amount of multi-modal data collected by wearable devices (Ginsburg et al., 2024; Baig et al., 2017). These limitations could result in low compliance among the patients and delays in accessing collected data from clinicians, which further affect timely diagnosis and intervention.
Recent exploration in the HCI community demonstrates the promising potential of leveraging innovative AI technologies for RPM (AI-RPM)to address the aforementioned limitations of traditional RPM approaches. For instance, conversational agents (CAs) powered by large language models (LLMs) enable asynchronous patient-care provider communication to verbally collect patients’ health conditions through regular health check-in conversations (Yang et al., 2024). LLMs have also been explored to organize and summarize key information from massive RPM data on clinician-facing dashboard (Wu et al., 2024). Meanwhile, another line of research explored using AI prediction algorithms with RPM data to assist clinicians’ decision-making in their clinical workflow (Yin et al., 2024).
Despite the promising opportunities of AI-RPM systems, they could lead to additional difficulties and burdens if not appropriately designed. Recent studies discovered that AI prediction models without faithful and reliable explanations cannot provide helpful information (Antoniadi et al., 2021) or actionable insights to clinicians (Rane et al., 2023; Zhang et al., 2024), lose clinician trust (Zhang et al., 2020; Chanda et al., 2024), and even result in human-AI competition (Zhang et al., 2024) Moreover, in high-stakes, high-risk, and uncertain clinical scenarios, clinicians follow strict clinical guidelines and specifications in their daily workflow. AI systems that are in conflict with clinical specifications cannot be used and even cause potential harm to stakeholders (Romero-Brufau et al., 2020).
In contrast, human-centered AI (HCAI) design guidelines (Shneiderman, 2022; Wang et al., 2023b) underscores the importance of engaging stakeholders in the system design process and understanding stakeholders’ workflow, in particular, the needs and challenges they encounter. By following HCI design guidelines, AI systems can be seamlessly integrated into stakeholders’ workflow where stakeholders can confidently understand, trust, and use AI-powered systems in their current workflow. A number of studies have demonstrated the effectiveness of human-centered AI systems in medical and healthcare scenarios to facilitate the needs of diverse stakeholders, including older adults (Hao et al., 2024), cancer patients (Wu et al., 2024), and clinicians (Wang et al., 2023b; Fogliato et al., 2022; Yang et al., 2019; Zhang et al., 2024). Nevertheless, how to employ HCAI guidelines in the design of AI-RPM systems has not been studies yet, and little is known about the challenges concussion clinicians face in remote monitoring youth patients and how AI-RPM technologies could be of practical benefit.
In this work, our objective is to investigate the current workflow of concussion clinicians with respect to the challenges in remote monitoring and decision making of youth concussion patients. In particular, we focus on how clinicians collect patients’ mental health-related information, what decisions they need to make, and what the difficulties are in their current workflow. In addition, we want to explore the design opportunities of AI-RPM technologies in supporting clinicians’ needs for remote monitoring and decision-making during their practice. In this study, we seek to explore two main research questions (RQs):
  • RQ1: What challenges and needs do concussion clinicians face in remote monitoring of mental health symptoms and decision making with multi-modal RPM data for youth concussion patients?
  • RQ2: What are the design considerations needed for AI-RPM technologies to support concussion clinicians in their workflow?
We approached to address these RQs through a three-stage study design: first, we conducted a formative study of semi-structured interviews with six highly professional concussion clinicians and derived a suite of design considerations of AI-RPM technologies that could support clinicians’ need for remote monitoring and clinical decision-making. We came up with preliminary designs of clinician-facing interfaces with different AI-RPM technologies based on design considerations. Subsequently, we engaged the same group of clinicians in a design interface evaluation study to further collect feedback and insights on the design considerations of different AI-RPM technologies. We consolidate the design feedback and revise the preliminary interface design to come up with a complete clinician-facing system design interface, along with the design considerations, as the final artifacts.
This paper makes the following key contributions: (1) we identified the challenges and needs that concussion clinicians encounter during their clinical practice with youth concussion patients, particularly with respect to patient mental health sequelae; (2) we derive a suite of design considerations and visualization feedback to empower advanced AI-RPM technologies to support remote patient monitoring and clinicians’ decision-making; and (3) we deliver a user-centered interface design of a concussion clinician-oriented dashboard system supported by multiple AI-RPM technologies, where the initial interface design and refinements are grounded in users’ needs and feedback.
2. Related Work
We review prior research on the negative influence youth have due to the mental health sequelae in post-concussion, AI-RPM technologies in clinical settings, as well as designing AI-powered systems for clinical decision support.
2.1. Challenges Related to Mental Health Sequelae Following Concussion
Mental health sequelae, such as anxiety, depression, and suicidality, are common among youth patients (ages 11–17) following a concussion (Stein et al., 2019; Gornall et al., 2021; Ledoux et al., 2022). Among these youth concussion patients, 70–80% of them experience longer recovery periods, normally more than 28 days (Scorza and Cole, 2019; Cuff et al., 2022). Studies have already shown the importance of integrating the assessment of mental health conditions into clinical follow-up visits (Gornall et al., 2021; Russell et al., 2023). Monitoring youth concussion patients’ mental health-related information enables clinicians to timely identify any abnormalities in the patient’s mental health and make mental health-related decisions, such as referring the patient to a mental health expert for further assessment or intervention. Therefore, early identification of mental health sequelae in the initial stages of concussion recovery is crucial for reducing the risk of worsening mental health issues in youth concussion patients and minimizing the negative impact on their concussion recovery.
However, there are challenges when it comes to monitoring youth concussion patients’ mental health. Most youth concussion patients are advised to rest at home after a concussion (Silverberg and Mikolić, 2023), their mental health sequelae-related symptoms, such as anxiety, often manifest in their everyday environments, particularly at home, which makes concussion clinicians difficult to detect within clinical settings. Although clinicians can inquire about mental health sequelae-related symptoms during clinical visits, recall bias (Frissa et al., 2016; Van den Bergh and Walentynowicz, 2016) may prevent patients from providing complete and accurate information to concussion clinicians. Moreover, clinical follow-up visits for youth concussion patients often lack consistency. Although it is recommended to conduct follow-ups every three days after a concussion (Agency, 2024), one study showed that actual follow-up rates vary significantly, with the lowest being only 13.2% (Ramsay et al., 2023). Insufficient follow-up prevents concussion clinicians from collecting information to accurately assess the youth patient’s mental health sequelae timely, making it difficult to make mental health-related decisions such as conducting further assessments or making a referral to a mental health expert. Therefore, it is essential to remotely monitor and collect behavioral data from youth concussion patients outside of clinical settings to assist clinicians in the early detection of mental health sequelae.
2.2. AI-RPM technologies in Clinical Settings
Remote Patient Monitoring (RPM) includes primarily wearable devices and home-use medical equipment that collect patient health data outside of clinical settings (Mendel et al., 2024). These systems transmit monitored patient data to healthcare providers for real-time monitoring and assessment, thereby supporting clinicians in their decision-making (Catalyst, 2018; Baig et al., 2017; of Health et al., 2008). Wearable devices, for example, serve as valuable tools for remotely monitoring patient data. Existing studies have leveraged wearable technology to track and collect key physiological data from concussion patients during their recovery period, such as physical activity levels, for recovery progress assessments (Tirosh et al., 2024). Additionally, researchers have investigated the relationship between post-concussion symptoms and physical activity in young patients. For instance, Yang et al. (2020) utilized ActiGraph, a smartwatch-based device, to monitor patients’ physical activity and collect movement-related data, such as step count.
Furthermore, researchers in the Human-Computer Interaction (HCI) and Computer-Supported Cooperative Work (CSCW) fields have started exploring mobile health (mHealth) technologies, which use mobile devices to collect patient information, facilitate self-management, and enhance remote communication between patients and clinicians (Corwin et al., 2024; Nyapathy and Arriaga, 2019; Salamah et al., 2021; West et al., 2018). For instance, El-Rashidy et al. (2021) indicated that mHealth can improve diagnostic efficiency and reduce healthcare costs for patients. RPM systems can display collected data via user interfaces for patient review (Griggs et al., 2018) or integrate the data into electronic health record (EHR) systems (Dinh-Le et al., 2019). However, given the significant time constraints that clinicians already face (Häikiö et al., 2020), along with their existing workload, the additional data generated by RPM systems could further exacerbate these challenges.
With recent advances in AI, AI-driven tools have the potential to assist clinicians by analyzing and summarizing large volumes of patient data and presenting information in a more interpretable format, thus reducing the cognitive load in clinical workflows (Häikiö et al., 2020). Research suggests that combining subjective patient perceptions with quantifiable metrics enhances the value and interpretability of patient data (Häikiö et al., 2020). For instance, voice assistants powered by large language models (LLMs) can interact with patients using natural language, capture subjective symptom experiences, and visualize them on a dashboard for clinicians (Yang et al., 2024). AI can also be leveraged for predictive analytics to help forecast patient health outcomes (Zhang et al., 2024; Pathak, 2024; Collins and Moons, 2019). Dabek et al. (2022) developed an AI model to predict the risk of mental health sequelae following a concussion, achieving 88.2% accuracy in predicting outcomes within 90 days post-injury. Despite these advancements, clinicians often remain skeptical about the accuracy of AI-generated predictions (Bohr and Memarzadeh, 2020), which often arises from the lack of transparency in AI decision-making. Providing clinicians with a clear understanding of AI’s limitations and actionable insights can enhance their understanding of AI outputs (Liao et al., 2020), thus increasing confidence in AI-assisted decision-making. Recent studies emphasize that AI explanations should align with clinicians’ cognitive models (Sendak et al., 2020; Sivaraman et al., 2023) to enhance trust and acceptance of AI-generated recommendations. Additionally, supporting AI predictions with evidence-based references and citations may further help clinicians validate AI outputs and increase confidence in AI-assisted decision-making. AI has the potential to support remote patient monitoring (RPM) systems by enhancing the efficiency of data collection and processing while supporting clinical decision-making. However, poor system design may bring new challenges and workload for clinicians. Therefore, integrating well-thought-out design with AI-RPM systems is important to ensure that they enhance rather than hinder clinicians’ work efficiency.
2.3. Designing AI-Powered Systems for Clinical Decision Support
In recent years, with the rapid advancement of artificial intelligence (AI) technology, human-computer interaction (HCI) researchers have increasingly explored how AI can better support clinicians in making clinical decisions. Their research spans various medical domains, including cancer diagnosis (Cai et al., 2019b, a; Denekamp et al., 2007), cardiotoxicity detection (Ahmed et al., 2024), diabetic retinopathy screening (Beede et al., 2020), and rehabilitation assessment (Lee et al., 2021, 2020).
Despite these advancements, the integration of AI-driven decision support systems into clinical workflows remains challenging. One major issue is that these systems often fail to fully consider the needs of stakeholders during the design phase, limiting their practical usability (Cai et al., 2019b; Green and Chen, 2019; Khairat et al., 2018; Lee et al., 2020; Romero-Brufau et al., 2020) and reducing clinicians’ trust (Yang et al., 2016; Elwyn et al., 2013). As a result, AI-driven systems may not effectively assist clinicians in making faster and more accurate decisions (Antoniadi et al., 2021). To address this gap, existing research has sought to incorporate human-centered design principles into AI-powered system development. Studies have shown that involving stakeholders early in the design process and establishing continuous feedback loops (Abdulaal et al., 2021) can enhance the AI system’s integration into clinical workflows (Wang et al., 2023b), which allows AI-powered systems to be tailored to the needs of different stakeholders, ultimately improving the usability of AI system. For example, Yang et al.(Sendak et al., 2020) conducted multiple discussions with clinicians to understand the challenges they face during meetings, collected their feedback, and iteratively refined the design of an AI-powered system based on this input. Similarly, Lee et al. (2020) engaged seven therapists throughout the system design process, which ultimately increased their willingness to use and satisfaction with the system. These studies highlight the importance of human-centered design in AI-driven systems.
Therefore, it is essential to develop a deep understanding of the challenges and needs of clinicians and actively involve them in the design workflow to fully leverage the potential of AI technologies to help clinicians with concussions detect mental health sequelae and make informed clinical decisions.
3. Formative Study
To better understand the workflow of clinicians treating youth concussion patients, we conducted a formative study focusing on three aspect: (1) how clinicians gather youth patients’ mental health-related information, (2) the decisions that clinicians need to make, and (3) the challenges they face in their current workflow.
3.1. Study Participants and Procedure
Table 1. Demographics of Participants in Our Formative Study
P# Gender Department Job Title Year of Practice
P1 Female Pediatric Sports Medicine Pediatric Sports Medicine Physician 16 years
P2 Female Complex Concussion Clinic Neuropsychologist 12 years
P3 Female Concussion Clinics Concussion Clinician 19 years
P4 Male Sports Medicine Division Chief 30 years
P5 Male Concussion Clinics Concussion Clinician 16 years
P6 Male Sports Medicine Non-Operative Sports Medicine Doctor 10 years
The key expert stakeholders in the targeted scenarios are concussion clinicians with prior experience treating youth concussion patients with mental health symptoms. These highly specialized concussion professionals work in very fast-paced environments and are often overloaded with numerous complex patient cases, which makes their time very limited. For this reason, we recruited six U.S.-based concussion clinicians. Recruitment was facilitated through professional networks within related clinical fields of domain experts in the research team via convenience sampling (Sedgwick, 2013). Six concussion clinicians were recruited as participants for the formative study, where each participant engaged in a semi-structured interview of 35-45 minutes conducted remotely via Zoom. Table 1 provides details on the demographics of the participants and their clinical experience. This study was reviewed and approved by first author’s Institution Institutional Review Board (IRB) and the study complied with the approved procedure for ethical research practices of human subject protection.
At the beginning of each interview, we first asked the participant to recall a recent case where they encountered a youth concussion patient with mental health sequelae. They described how they performed clinical practices to detect mental health sequelae and executed follow-up decision making. While participants were sharing their experiences, we prompted participants to tell more about what type(s) of patient clinical data they collected, used or expected to support the identification of mental health sequelae, both in and outside of clinical settings. After that, participants shared their experience with existing tools, if any, for collecting patient data, monitoring patient status, and detecting mental health sequelae. Participants shared their feedback and insights on the benefits and limitations of existing tools and the expectations of novel tools that could potentially help them. During the interview, we asked participants to refrain from disclosing any personally identifiable information (PII). The complete interview protocol is provided in Appendix A.
All interviews were audio-recorded and transcribed with participants’ consent. We employed the inductive thematic analysis approach (Braun and Clarke, 2012, 2019) to derive key themes. Two researchers first independently coded one session of the interview transcripts to identify meaningful segments and categorize segments into codes. The researchers then discussed individual codes and resolved any discrepancies, developed consensus codes, and applied the resulting codes to the remaining data.
3.2. Findings
Our qualitative analysis provided a comprehensive understanding of concussion clinicians’ workflow with youth concussion patients, and derived three major challenges that concussion clinicians encountered during the practices: (1) collecting patients’ mental health information outside the clinic; (2) communicating effectively with patients and their family members; (3) Assessing patients’ compliance with clinical recommendations.
3.2.1. Clinical Workflow of Concussion Clinicians
The majority of youth concussion patients are athletes where, and concussions are commonly caused by head injuries during sports participation. After they had a head injury, the youth patients underwent an initial assessment by primary care physicians. According to P4, primary care physicians conducted basic evaluations and symptom scoring based on international guidelines. If specialized care is needed, patients are referred to specialized concussion clinicians.
The clinical tasks of concussion clinicians primarily include seeing new concussion patients and following up with existing patients, which typically follow a similar routine. When patients visit the clinic, concussion clinicians assess the severity of concussion symptoms with standard scales as suggested by clinical guidelines. In addition, concussion clinicians will conduct interview sessions with patients to collect more in-depth information about their health conditions.
”And I guess the only way that I currently assess my patients in that way is by using the symptom score that is published in the international guidelines, at least as of 2017. That’s the symptom score that I use. And then interview. I will very commonly ask about Sleep habits. I will very commonly ask about nutrition habits. I am very aggressive in making I guess my patients go to school.” (P3)
For the majority of concussion patients, the best treatment available is just to keep rest both physically and cognitively. Thus, concussion clinicians will often engage in a conversation with patients after finishing the assessment to provide personalized lifestyle recommendations and self-evaluation methodologies that would be best suited for concussion recovery. For example, one participant (P5) stated that they often provide patients with symptom logs and suggested patients track their daily experiences. Patients will be asked to return the symptom logs during follow-up appointments, which typically occur after one week. Symptom logs are crucial for concussion clinicians because they will heavily rely on the symptom logs to get a thorough understanding of how patients underwent the past week and how the progression of recovery from concussion is. If symptoms persist beyond 28 days, further interventions are initiated, such as referrals to neuropsychologists or specialized concussion clinics (P3, P5, P6).
”Yeah, after we see them (patients) for the initial visit, we generally send them home with the log and ask them to fill it out every day based on how they felt, and then bring it back when they come back for their follow up appointment so that we can track how their symptoms have been going over time.” (P5)
Concussion clinicians (P1, P2, P3, P5, P6) emphasized that mental health concerns are highly prevalent among youth concussion patients. Thus, mental health assessments are an integral part of the concussion assessment process. Concussion clinicians (P6, P4) typically review the patient’s mental health history, including prior diagnoses, current treatments, and medication use. In addition, clinicians leverage standardized questionnaires, such as the Patient Health Questionnaire-9 (PHQ-9) (Levis et al., 2019) and the Generalized Anxiety Disorder 7-item Scale (GAD-7) (Mossman et al., 2017), to identify potential symptoms of mental health issues and the severity of symptoms. Concussion clinicians (P4, P6) also observe signals of emotional distress during the conversation with patients to detect potential mental health issues and will follow up with additional mental health screening if such signals are presented. Clinician (P6) reported that there is a high portion of concussion patients develop anxiety or depression following their injury. In such cases, clinicians (P4) often recommend cognitive behavioral therapy (CBT) interventions as the intervention strategy for the mental health issues mentioned above. However, if mental health symptoms are found to impede concussion recovery, concussion clinicians will need to collaborate with neurologists to determine the intervention plan, such as prescribing appropriate medications. For patients exhibiting severe mental health symptoms, clinicians (P2, P3) will require the patients to complete an additional suicide risk assessment. If no immediate suicide risk is detected, clinicians typically recommend continued rest at home with ongoing symptom monitoring. However, the concussion clinician (P2) mentioned that 10% of youth concussion patients develop suicidal thoughts after a concussion. In such cases, clinicians refer patients to mental health specialists, such as psychiatrists or psychologists.
”In general, of people who I see with like symptoms of like anxiety, depression, things like that after concussion. It’s probably about 40 to 50 percent.” (P6)
3.2.2. Clinicians’ Challenges in Tracking Patients’ Mental Health Information Outside of Clinics
As mentioned above, the first 28 days following a concussion diagnosis is a critical period of time for youth concussion patients’ recovery. However, patients only visit the concussion clinic every one or two weeks, while spending the majority of time outside the clinic, more specifically, outside of clinical supervision. This gap creates significant challenges for concussion clinicians, who (P2, P5) emphasize the importance of monitoring a wide range of mental health symptom indicators during this recovery period. For instance, clinicians (P1, P2, P3, P4, P5) would focus on collecting detailed information about sleep patterns (e.g., sleep hours and wake-up times), and physical activity levels. P3 noted that the sleep data has not really been integrated into concussion statements. Moreover, clinicians (P2, P3, P6) are closely interested in youth patients’ academic performance, such as whether they completed assignments on time, and social behavioral patterns, such as whether they rest at home consistently and maintain social connections rather than withdrawing socially. These physical and cognitive behaviors could provide critical insights for clinicians to assess and reflect on youth patients’ mental health status during concussion recovery.
”…if the worry is getting to the point where they (patients) can’t do their homework… Or if they aren’t leaving the house… not sleeping, really feeling terrible, and it seems to be consuming them… Then I would say that might like cross the line from just typical worry into like, okay, this is more of a mental health anxiety, you know, diagnosis type of thing.” (P2)
Furthermore, P4 stated that during clinical visits, the clinician mainly relies on nonverbal cues in the clinical environment to identify youth patients’ mental health conditions. If any abnormalities are detected, youth patients would be asked directly about mental health-related symptoms, such as suicidal tendencies. Mental health-related information is crucial, as mental health issues impact the clinician’s ability to adjust treatment plans in a timely manner. Moreover, in some cases, youth concussion patients may experience impairment in functioning or panic attacks, and the concussion clinicians will then focus on understanding what they had been through and what events triggered the episode to more accurately evaluate patients’ condition as well as design a more appropriate concussion recovery plan to avoid these negative events happening again, as suggested by P2.
”……I think, having more information about really what was happening at the time, that they had a symptom. What was triggering, that how that we know more real time information could be helpful.” (P2)
Despite the critical need for timely health data of patients at home, concussion clinicians (P2, P5) reported significant challenges in collecting accurate and reliable information about patients’ mental health outside the clinic. Ideally, clinicians would benefit from having real-time access to patients’ health data while they are at home. However, such a system is currently unavailable. Instead, clinicians must rely on patients’ recall of their living experiences during clinical visits to reconstruct what happened during the recovery period. Although these discussions could help inform clinical decision-making, patients’ memory recall is often incomplete or inaccurate, especially when clinical visits span several weeks. This lack of reliable data collection methodology can hinder concussion clinicians’ understanding of the patient’s recovery progress and delay timely mental health interventions, such as referrals or additional assessments. Once youth concussion patients leave the clinic, clinicians have no way to obtain this information in a timely manner.
”…more real-time information could be helpful. Because by the time we see them (patients) in the clinic that could be you know, weeks later they might not even remember that that (mental health sequelae-related symptoms) had happened…… It’s hard for anyone to remember what they did, you know, weeks before.” (P2)
Youth concussion patients currently have some remote communication options, such as making phone calls with the provider team, to report their health status regularly to clinicians when they are outside of clinic. However, clinicians (P2, P3) reported several limitations with phone communication. First, youth concussion patients rarely initiate calls to their concussion clinicians. Moreover, even if youth concussion patients make a phone call, clinicians may not always be in the office or available to answer immediately, which can delay the transmission of critical health information.
An alternative option is using MyChart, a patient portal that allows youth concussion patients to access their medical records and send messages to their healthcare providers. Concussion clinicians (P4) strongly recommend patients use MyChart to share health-related information, particularly when youth concussion patients are reluctant to make direct phone calls. Compared to phone communication, MyChart offers the advantage of asynchronous messaging, enabling clinicians to review patient information and respond at their convenience.
Despite its potential benefits, the adoption of MyChart is limited in current clinical scenarios. Concussion clinicians (P2, P6) noted that youth patients rarely use the messaging feature in MyChart and rarely take the initiative to contact clinicians. Instead, parents of these youth patients are more likely to use MyChart to communicate with clinicians, which impedes the information flow between patients and clinicians.
”One in 10 [patients] maybe, or even one in 20 [patients] might send a message with a question.” (P6)
Furthermore, clinicians caution against relying on MyChart for emergencies, especially when patients experience suicidal tendencies. Clinicians (P2) warned that their busy schedules prevent them from checking messages in real time, which could lead to delays in critical interventions and put patients at risk.
”Yeah, it’s not really immediate. You know, cause they (patients) might send it at night, and then we don’t maybe don’t see it till the next afternoon, and you know. We usually, you know, tell them if it’s an emergency, don’t use that method because we don’t really know when we’re gonna see them or be able to respond to it.” (P2)
3.2.3. Clinicians’ Challenges in Communicating with Patients and Families About Mental Health Issues
Communication challenges persist not only outside the clinics but also during in-person consultations. During in-person visits, parents often accompany their children. However, some patients may feel uncomfortable to disclose sensitive mental health issues. One clinician (P1) described a case where a patient with severe anxiety disclosed self-harm and suicidal thoughts only after requesting to speak privately, away from their parents’ presence.
“I stepped out of the I had the mom stay in the room……And I said, ’Is there something wrong? Is there something going on?’ He said, ’Yeah, I need to talk to you privately.’” (P1)
Clinicians suspect that a major reason that patients choose not to disclose their mental health status is their fear of being labeled as having a mental illness. Being labeled as having a mental illness can affect how they are perceived socially, which could potentially isolate them from their peers. Some patients may also struggle with the societal expectation of their physical recovery, especially if youth patients are athletes who are expected to perform at a high level of competition.
This highlights the sensitivity of mental health disclosures is intertwined with many societal factors that may not be directly related to mental health symptoms.
Moreover, some concussion clinicians (P1, P4, P5) reported that sometimes both youth concussion patients and their parents may deliberately withhold information about the patient’s mental health. In one case (P4), a youth concussion patient was a successful athlete, and the parents downplayed their child’s mental health symptoms during clinical visits. The parents may expect to expedite the child’s recovery process so that they can return to their athletic career sooner. These parents may worry that prolonged rest due to mental health sequelae could hinder their child’s athletic performance and career trajectory. However, when mental health issues are not disclosed and addressed in a timely manner, the issues could get worse and lead to serious consequences, such as self-harm or an increased risk of suicidal behavior. The consequences of severe mental health issues could lead to long-term physical and cognitive impairment of the patients as well as hinder an athlete’s ability to return to sports.
”Their (patients) parents are very vested in the child’s athletic career……the parents come in and there’s obviously an agenda to minimize their symptoms. Or you know, just try not to buck the system, if it’s gonna limit their (patients) activity.” (P4)
To detect undisclosed mental health issues, clinicians rely on a combination of strategies. One approach is to build trust and collaborate directly with parents. Another is to observe the patient’s behavior during consultations. For example, clinicians (P1, P4) can observe patients’ behaviors, such as signs of anxiety or a persistently quiet and disengaged attitude, during clinical visits. Moreover, worsened concussion symptoms such as insomnia and low activity levels may indicate underlying issues beyond the concussion itself, such as mental health sequelae. One clinician (P3) stated that if a youth patient’s concussion symptoms do not improve within four weeks after a concussion, it may indicate underlying mental health issues that are contributing to a prolonged recovery period. However, both observing patients during in-person visits and asking about their sleep and activity levels at home require clinicians to gather this information during the clinical visit. Once the patient returns home, clinicians can hardly access this information.
”Because we expect 96 percent of concussions to be healed by the four-week point. And so if it’s not healed yet, then it’s probably not the concussion, right? It’s probably either a headache syndrome or a mental health concern.” (P3)
3.2.4. Clinicians’ Challenges in Assessing Patients’ Compliance with Clinical Recommendations
At the end of each clinical visit, concussion clinicians typically provide patients with recommendations related to physical activity or sleep. These clinical recommendations are intended to serve as the concussion treatment plan to support both physical and mental recovery. However, understanding whether patients adhere to these recommendations is a persistent challenge.
Concussion clinicians (P2, P4) rely on patients’ self-reports to assess compliance during the time period between the last visit and the current one, despite these reports are often unreliable. For instance, some patients may claim to have followed the recommendations, but their responses to mental health questionnaires reveal high levels of anxiety or depression, indicating limited or no improvement.
“What we recommended, if they(youth patients) did it or not, they might not have done it… … We don’t know if they’re following them or not. Or they come in, and they say they did, but we don’t know if they did.” (P2)
Concussion clinicians (P) noted that there is currently no effective way to track patients’ compliance with clinical recommendations outside the clinic. Clinicians (P4) mentioned that they often ask patients’ parents about youth patients’ adherence to recommendations but found that information from family members is also hardly reliable. One reason could because some parents are frequently away from home due to work commitments, which leaves them with limited knowledge of patients’ daily health-related information. In such cases, clinicians may reach out to other family members for a closer understanding of the patient’s mental health conditions when they are at home. However, making such phone calls consumes a considerable amount of time in clinicians’ busy schedules, and the information collected through these calls is often limited.
“So I think sometimes the parent comes in with a child who isn’t around them that much at home. Like maybe it’s the working parent, and they’re just sort of the driver and they don’t have a lot of information…… So sometimes I can’t get good information ’cause they’re not engaged as much as I would, you know, would hope.” (P4)
4. Initial System Design
In our formative study, We found that clinicians have a need for remotely monitor the health-related information of youth concussion patients at home, particularly regarding their mental health. Additionally, we found that clinicians face challenges when making mental health-related decisions, particularly during clinical visits, where it can be difficult to gather mental health information from patients and their families through communication. However, clinicians rely on this information to guide their clinical decisions and treatment plans.
Building on these identified challenges and needs, we proposed AI-RPM technology that could potentially assist clinicians in remotely monitoring patients’ at-home data and addressing decision-making difficulties. We then visualized a possible system interface based on these technologies and detailed the specific functions of the system. The goal of our system design is to explore whether concussion clinicians find the technologies useful in remotely collecting youth patient mental health-related information outside of clinics and supporting clinicians’ challenges associated with mental health-related decision-making.
4.1. Applicable AI-RPM Technology
4.1.1. Wearables in Tracking Patients’ Sleep Pattern and Physical Activities
Based on Section 3.2.2, clinicians expressed a desire to monitor patients’ at-home quantitative data, such as sleep and exercise patterns, as these factors are closely related to concussion recovery and mental health status. However, they currently lack effective ways to access real-time health data from patients at home. In addition, as noted in Section 3.2.4, clinicians frequently ask patients whether they adhere to clinical recommendations, such as maintaining proper sleep and exercise routines. However, clinicians often struggle to verify the accuracy of patients’ self-reported information.
To address these challenges, we propose that wearable devices such as smartwatches could help clinicians collect real-time sleep and exercise data from patients and provide these data to clinicians for timely review. Wearables offer a non-invasive and natural way to monitor patients’ health metrics without significantly disrupting their daily routines. As discussed in Section 2.2, wearable devices can continuously monitor patients’ sleep and physical activity data in real-time (Dias and Paulo Silva Cunha, 2018; Baig et al., 2017; Bate et al., 2023). We believe that patients’ sleep patterns and physical activity data collected through wearables could enable clinicians to remotely access at-home patient data, track adherence to sleep and exercise recommendations, and ultimately make more informed mental health-related clinical decisions.
4.1.2. LLM-Powered Conversational Agents in Supporting Patients Self-Report Mental Health Symptoms
In Section 3.2.2, clinicians expressed a need to understand patients’ experiences with symptoms such as panic attacks, including what specifically triggered them. However, patients often struggle to accurately recall these details, and such subjective experiences cannot be easily quantified using standardized scales or medical instruments. Additionally, based on Section 3.2.3, concussion clinicians reported that youth concussion patients sometimes conceal their mental health issues from their parents. It was only during a private conversation with the clinician that the patient disclosed the truth about their mental health issues.
Given this challenge, we hypothesize that LLM-powered conversational agents (CAs) (Yang et al., 2024; Mahmood et al., 2023; Chan et al., 2024) could help clinicians collect mental health-related information effectively. CAs can be deployed in patients’ homes to engage in natural language conversations, ask relevant questions, and record dialogue content. AI can then analyze historical conversations to summarize the content of the conversation, which allows clinicians to quickly gain insights into their mental well-being. One study found that children perceive conversational agents as warmer and more reliable than their parents (Van Brummelen et al., 2023). Therefore, we speculate that CAs could encourage patients to disclose their mental health issues, enable AI to analyze these conversations, extract key information, and securely transmit relevant insights to clinicians while preserving privacy. As a result, clinicians could review patients’ historical conversations with CAs, gaining insights into their subjective experiences, emotional states, and mental health-related life events—ultimately supporting a more accurate assessment of patients’ mental health.
4.1.3. AI Risk Prediction in Supporting Mental Health Sequelae Detection
As discussed in Section 3.2.3, clinicians reported that patients and their parents sometimes deliberately conceal mental health conditions, making it difficult for clinicians to develop appropriate treatment plans. In addition, clinicians noted that the severity of mental health issues in a patient influences treatment plans and subsequent clinical decisions by clinicians, such as whether a referral is necessary. Given these two challenges, we believe that AI-powered risk prediction for disease progression could provide valuable support to clinicians. AI has the capability to analyze patients’ health data and use algorithms to predict the likelihood of developing a condition over a certain period (Collins and Moons, 2019). Recent studies have already explored AI-driven risk prediction for mental health sequelae following a concussion and have found that AI predictions can achieve high accuracy (Dabek et al., 2022). By leveraging AI risk score prediction, clinicians may no longer have to rely solely on patients’ self-reported mental health status and daily behaviors when making assessments. Moreover, if patients or their parents choose to withhold information, AI could help clinicians detect potential mental health sequelae in a timely manner. If a patient is flagged as having a high risk of developing mental health sequelae, the clinician can proactively schedule follow-up visits and make necessary mental health-related clinical decisions. Therefore, we hypothesize that AI-driven risk prediction has the potential to assist concussion clinicians in detecting mental health sequelae more effectively.
Figure 1. The initial system design draft for evaluation and gathering information from concussion clinicians: (A) AI risk score prediction for mental health sequelae, (B) patient self-report symptoms, (C) patient sleep and physical activities.
4.2. Preliminary UI Design of the System
Based on the potentially applicable technologies that could support concussion clinicians in addressing these challenges, we designed three modules for our initial system design, as shown in Figure 1: (1) AI risk score prediction for mental health sequelae (Fig. 1.A), (2) patient self-reported symptoms (Fig. 1.B), and (3) patient sleep and physical activity tracking (Fig. 1.C). During our UI design process, we adhered to Human-Centered AI (HCAI) principles to ensure a user-friendly interface. Our goal with this design is to visualize the system architecture and interface based on the previously proposed technologies, which makes our system design more intuitive for clinicians. Then, we can use the initial system design to gather clinicians’ perceptions and feedback on visual representation, data types, and technologies.
4.2.1. The AI Risk Score Prediction Module
The first module is AI Risk Score Prediction(Fig 1.A). The module shows the probability of mental health sequelae within one year with a range from 0% to 100%, where the higher probability score signifies a higher probability of experiencing mental health sequelae for clinicians to review. Feature importance scores are also provided to provide additional explanations for clinicians to decide whether to trust or not to trust the AI-predicted score. This module also offers text-based explanations to help clinicians understand the sources of the factors that contribute to the risk score. At the bottom is the risk score chart. Clinicians can hover over any day on the risk score chart to view the likelihood that young patients develop mental health sequelae at that specific point in time. Additionally, clinicians can select different dates to review patient data from various time periods. This module aims to enable clinicians to remotely monitor youth patients’ mental health sequelae without requiring in-person consultations, which supports concussion clinicians in making timely, informed decisions based on patients’ mental health conditions.
4.2.2. The Self-Report Symptoms Module
The second module is the self-reported symptoms (Figure 1.B). On the right side of the module, a history of conversations between youth concussion patients and conversational agents (CAs) is available for concussion clinicians to review. Based on the conversation history, AI detects the symptoms that are mentioned by patients and highlights categories that are relevant to mental health sequelae, such as anxiety, depression, and suicidal ideation. The categories are displayed on the left side of the panel. If a patient self-reports symptoms within a specific category, the corresponding indicator changes from green to red, which reminds clinicians to review the entry. To facilitate efficient navigation, clinicians can click on a category name or use the search bar to enter keywords, which prompts the right-side module to display the relevant conversation excerpts. Additionally, clinicians can select different dates to review patient data from various time periods. The primary goal of this module is to allow concussion clinicians to quickly assess patients’ mental health symptoms at home, extending support beyond clinical settings.
4.2.3. The Sleep and Physical Activities Module
Lastly, we designed the sleep and physical activities module (Fig. 1.B). The module presents sleep-related information that concussion clinicians want to track for their patients at home, including sleep hours, light sleep, and wake-up time, as well as activity-related data such as step count. When a clinician selects any category, the module displays the patient’s relevant data from the past week. Clinicians can also select different dates to review patient data from other time periods. Additionally, the module provides the source of the data, which is through wearable sensors, such as smartwatches. The goal of this module is to assist concussion clinicians in monitoring youth concussion patients’ sleep and physical activity at home, addressing the current gap in existing clinical systems that lack this information.
5. Design Interface Evaluation
In the previous section, we introduced the potential of AI-RPM technologies to assist clinicians in remotely collecting patient information and supporting mental health-related decision-making. We then designed a system that integrates these technologies, focusing on how the collected data is presented and utilized. Based on the initial system design, we conducted a design interface evaluation (Savage, 1996) to gather feedback from the six original clinicians who represent our target end-users. The goal of this section is to explore concussion clinicians’ perceptions of the system—specifically, whether they find these technologies useful for remote patient monitoring and how the system might better support them in addressing challenges related to mental health decision-making.
5.1. Procedure
We conducted user reviews with the same participants from our formative study. Due to the time constraints of concussion clinicians, each user review was limited to 10 minutes. We conducted a Zoom screen-sharing session to present our three modules to participants, introducing the data and functionalities included in each module. Based on each module, we then collected verbal feedback from participants regarding data type, visualization, interaction, usability, and potential use cases. Our research team documented their responses through note-taking. After the session, we collected participants’ feedback and applied axial coding to identify key themes. We then identified design modifications based on the themes. Next, we iterated on our initial system design following design modifications to come up with the final version of our system design. Our goal is to ensure that the final system design shows the potential to effectively support concussion clinicians in addressing the challenges of detecting mental health sequelae. The detailed interview protocol for our design interface evaluation is provided in Appendix B.
5.2. Findings
In this section, we presented the feedback of concussion clinicians on the initial system design, including the features they found helpful, their concerns and points of confusion, and suggested improvements to the system.
5.2.1. AI Risk Score Prediction as Supportive Evidence for Clinicians in Patient Communication
Concussion clinicians (P1, P5, P6) reported that if the AI risk score prediction module provides accurate data, it can be useful in assisting clinicians with treatment planning and making timely adjustments to clinical decisions, such as issuing referrals or a set of follow-up visits.
”That [the AI risk score prediction] might be something helpful early on for them to say, Hey, this is somebody who probably gonna benefit from seeing another provider earlier rather than later.” (P5)
Moreover, P1 mentioned that the AI Risk Score Prediction module is straightforward and has the potential to effectively support clinicians in communicating with patients and their parents when concerns arise about patients’ mental health sequelae during clinical visits. By directly presenting the risk scores of potential mental health outcomes, youth patients and their families gain a clearer understanding of both the current mental health state and the possibility of potential future mental health sequelae, particularly for youth patients and families who are hesitant to discuss mental health issues with concussion clinicians.
”I think it’s pretty simple, and if I was able to show that to a family… maybe something as simple as this to show [the patient’] mom… ’I’m concerned about your son… There’s a 50% risk that may develop [mental health sequelae]. We should really be in touch with it.’ Yes, I think this is great.” (P1)
5.2.2. Explanation of AI Results and Data Sources as Supportive Information for Clinicians’ Understanding
However, concussion clinicians expressed confusion regarding the data sources and the interpretation of results in certain modules. First, they raised concerns about the AI Risk Prediction module. When we presented this module, despite the interface providing key contributing data and a brief explanation of data sources, clinicians still found it difficult to understand the meaning of the risk score. They were also unsure whether the score was derived from other modules or from self-reported questionnaires administered to patients, such as the Generalized Anxiety Disorder 7-item Scale (GAD-7) and the Patient Health Questionnaire-9 (PHQ-9). Additionally, clinicians expressed confusion about the self-reported symptoms module. Specifically, P2 and P4 mentioned that they were unclear about how self-reported symptoms were collected.
”So it’s just my confusion. So when it says self-reported symptoms that’s based on like a questionnaire that we’ve given them about their depression and anxiety level? Or this is what AI predicts that they have a 50 percenty chance of like developing depression based on that?” (P2)
5.2.3. CAs and Wearables as Additional Valuable Data Sources Outside the Clinic
Concussion clinicians (P3, P6) expressed optimism about the potential of the self-reported symptoms module. They believed that patients might be more willing to share information with a voice assistant than directly with a clinician, making this module a valuable additional data source. One clinician mentioned that they would like to quickly review self-reported symptom data before each clinical visit to better understand youth patients’ current conditions. At the same time, P6 noted a potential challenge: if patients have a voice-based conversational agent at home, they might place it in a corner and forget about it. To address this, they suggested that CAs should initiate daily conversations proactively to collect patient information more effectively.
”I guess from a perspective of maybe they would tell this device things that they wouldn’t tell me. Possibly I could see that.” (P6)
Among the three modules in our system design, concussion clinicians (P3, P4, P5) particularly appreciated the sleep and physical activity module. They highlighted that such data is currently missing from electronic health record (EHR) systems, yet wearable-generated data on sleep and physical activity could serve as key indicators for monitoring adolescent concussion patients’ mental health and recovery progress. P3 believed that this information could help monitor patients’ health on a weekly basis and assess whether they have followed clinical recommendations. However, P6 raised a concern that if a patient already owns a smartwatch, wearing a second device for physiological monitoring might be inconvenient. Additionally, having two devices could serve as a constant reminder of their injury, potentially reinforcing the perception that they have not yet fully recovered. Additionally, P3 suggested that it would be best if patients did not see the data collected by the smartwatch to avoid reminding them of their condition.
”This one (the Sleep and Physical Activities panel) to me might be the most helpful [one]… I think that’s always hard [to get].” (P5)
”So I really like the sleep and physical activity, but I would suggest that the patients not see that data that it only goes to the physician. Because I don’t want it to cause them anxiety. So yeah, I really like that.” (P3)
5.2.4. Variations in Data Detail and Presentation Types for Clinicians’ Preferences
Concussion clinicians exhibited varying preferences regarding the type and level of detail in the data presented within the sleep and physical activity module. Some clinicians preferred more detailed information than what was currently displayed. For instance, clinicians (P2, P3, P5) suggested including details related to napping under the sleep pattern section, such as whether the patient took naps and their duration. Additionally, clinicians (P1, P3) emphasized the importance of heart rate in assessing concussion recovery and monitoring mental health sequelae. Specifically, they expressed interest in viewing daily resting heart rate during sleep over the course of a week. They noted that during recovery, a patient’s heart rate tends to be lower, indicating a calm state, and it often correlates with obtaining eight hours of quality sleep. In contrast, if a patient experiences fragmented or restless sleep, their resting heart rate increases. By analyzing these patterns, clinicians could better assess whether a patient is experiencing anxiety.
”If you (patients) are having a light, irritable kind of sleep, I would expect the resting heart rate to be in the seventies or eighties. [It means] maybe they (patients) are anxious. Maybe there’s something else going on. ” (P1)
Moreover, in the self-report symptoms module, clinicians (P3, P5) suggested they want to include dietary information, such as appetite or nutrition, as they noted that when youth patients experience depression or high levels of anxiety, both the patients and their parents often report decreased appetite. Regarding the symptoms part of the module, instead of focusing solely on whether youth exhibit symptoms of mental health sequelae, concussion clinicians (P3, P5) emphasized that they are more concerned with the duration and severity of these symptoms. The severity and duration of symptoms have different impacts on clinicians’ treatment plans and decision-making, which potentially helps clinicians distinguish between general emotional issues and mental health sequelae following a concussion.
”…persistent or non-improving symptoms. That’s what I look for.”(P5)
Additionally, P3 expressed a new need for a feature that could indicate how adherence to recommendations impacts recovery rates, suggesting that such insights could enhance clinical decision-making.
”I don’t know if my patients are following my recommendations and knowing if they are and how that affects their recovery would be really interesting.” (P3)
5.2.5. Clinicians’ Concerns About Emergency Situation and Time Constraints
In the self-reported symptoms module, concussion clinicians placed significant emphasis on alerts related to suicidality. P1 emphasized the need for immediate intervention when suicidal tendencies emerge in youth concussion patients. However, they also highlight a critical gap in current clinical practice—the absence of established protocols for managing or intervening in such situations. When the user interface flags a potential suicide risk, clinicians report uncertainty about how to respond effectively due to a lack of concrete, actionable guidelines. Failure to intervene in a timely manner could result in missed opportunities for intervention and serious legal or ethical consequences. Additionally, even when clinicians recognize signs of suicidal ideation, managing emergency interventions without triggering resistance from parents or youth patients remains a highly sensitive and complex challenge. Furthermore, P4 reported that they face a dilemma: The youth generation experiences frequent mood swings and may sometimes exaggerate their feelings and symptoms. If a youth patient reports self-harm intentions that turn out to be overstated, clinicians worry about overreacting to false alarms. Striking the right balance between taking reports seriously and avoiding unnecessary escalations is a persistent concern in their practice. Clinicians (P2, P4) generally believe that parents should be brought into the loop when such situations arise.
”So I think this is when you need to loop the paradigm and say, hey, your child has been identified through our software. You know, cause I think, especially since we’re dealing with minors, you’re gonna have to get the parents loop too.” (P4)
Moreover, clinicians (P2, P5) expressed concerns about their already heavy workloads, stating that they have little time to thoroughly review such detailed data. These clinicians noted that having dedicated staff to help filter and process data would greatly reduce their burden. Without such support, they often skim through relevant data quickly on the day of the appointment or shortly before, just to gain a basic understanding of the youth concussion patient’s condition.
”I think that’s gonna that would be really tedious to try to monitor and create extra work for us.”(P5)
Regarding the newly introduced data and user interface modules, concussion clinicians (P1, P2) expressed concerns about the difficulty of finding time to process the data thoroughly. Clinicians already feel overwhelmed by the excessive features within current electronic health record (EHR) systems and their tasks. P1 highlighted that processing large amounts of data between tightly scheduled appointments—sometimes up to 20 patients per day—makes it nearly impossible to dedicate sufficient time and energy for in-depth analysis. Although monitoring systems can collect vast amounts of valuable data, clinicians often lack the capacity to fully review this information within their workflows. Given their heavy workload, the time clinicians can dedicate to each patient is extremely limited—sometimes as little as 15 minutes per appointment. As a result, concussion clinicians must rely on brief time windows, such as the moments before a scheduled meeting, to extract key information and quickly assess the condition of youth concussion patients.
”I would need to look at it on that day of the visit… Honestly, we are quite overwhelmed with the EHR. And all the different features and everything that we’re asked to do with it.”(P1)
6. Refined System Design
Based on the design interface evaluation, clinicians acknowledged the value of AI-RPM technologies which can potentially help them remotely collect and monitor patient data. In addition, clinicians found that the data, features, and presentation methods in our initial system interface could support their mental health-related decision-making. At the same time, they provided valuable feedback and suggestions based on the initial system design. In this section, we present design considerations derived from the design interface evaluation and use them to refine our initial system design. We then provide a refined system design in this section. The goal of presenting this refined system design is to visually illustrate the design considerations more effectively and provide insights for future studies.
6.1. Design Considerations
6.1.1. Explanation of Data Sources and Module Relationships
Through design interface evaluation, we found that while concussion clinicians considered the risk score useful for health decision-making and communication with patients and their families, they also raised concerns. Specifically, they were unsure how the score was generated and how it connected to other modules. To address this issue, we recommend using visualizations to illustrate the relationships between AI results and other data sources. This would enable clinicians to intuitively understand how various factors influence the risk score, potentially enhancing their trust in the system. One clinician explicitly mentioned wanting to see the relationship between clinical recommendations and patient recovery. So we suggest visualizing this by displaying the correlation between the risk score and exercise recommendation compliance.
6.1.2. Personalized Time Selection
One clinician found the AI-generated risk score unhelpful because the initial UI displayed a one-year risk score, whereas she typically treats youth patients for only four weeks. If a patient does not recover within this period, they are referred for further evaluation. Therefore, we believe clinicians would benefit from viewing AI model risk scores over different time spans to meet their varying needs and align with their workflow. To accommodate different clinical needs, we propose adding a time range selector to the module. This feature would allow clinicians to customize the risk score display based on their specific decision-making timeframes, offering personalized support for different clinical workflows.
Furthermore, for the sleep and activities module in Section 5.2.3, a clinician expressed a preference for viewing data on a weekly basis. To align with this preference, we recommend that the module default to displaying seven days’ worth of data while also providing the option to explore detailed data for specific dates. This would allow clinicians to closely examine anomalies, particularly when irregular patterns appear on certain days.
6.1.3. Providing Information on Other Stakeholders to Support Collaboration
Concussion clinicians noted that in the self-reported symptoms module, the most critical yet challenging situation for clinicians is detecting suicidality in patients. However, they expressed uncertainty about how to remotely respond if a patient exhibits self-harm tendencies. Additionally, clinicians raised concerns about the authenticity of such reports, as youth patients may sometimes exaggerate their symptoms. Clinicians emphasized that if an emergency occurs, parents should be involved, as they are considered key stakeholders in managing such situations. In response to this concern, we propose integrating an emergency contact module within this UI module, which allows clinicians to contact parents when necessary. Since parents may have quicker access to their children and be able to take immediate action, this feature could enhance response efficiency. Additionally, clinicians expressed a need to track the number of days symptoms persist, as this directly reflects a patient’s recovery progress and informs mental health-related decisions such as giving a referral. We recommend that when CAs engage in conversations with patients and their parents, they also document their observations on the duration of mental health symptoms. Then from the design side, we propose adding a data point in the module to display symptom duration.
In addition to monitoring the patient’s emotional state, clinicians often ask parents about behavioral changes. To enhance data comprehensiveness and objectivity, we suggest incorporating parent-VA (virtual assistant) conversation data into this module, offering an additional perspective on the patient’s condition. Finally, to reduce confusion and improve the module’s usability for clinicians, we need to clearly define how the data is collected. Ensuring transparency in data sources will help clinicians trust and effectively utilize this module in their practice.
6.1.4. Balancing Additional Data Presentation and Cognitive Load
In the initial system design, certain key data types that are highly important to clinicians were not covered. As section 5.2.4 mentioned, concussion clinicians emphasized the need for heart rate during sleep, nap times, wake-up time, and sleep time rather than just total sleep hours. To address this, we recommend displaying these additional data points within the module. However, given clinicians’ concerns about data overload—especially as they noted that existing electronic health record (EHR) systems have already increased their workload—we propose a customizable data selection feature. Within the system interface, clinicians can personalize their data view based on their specific needs and workload at different times, rather than being presented with all data at once. Additionally, we believe it is essential to prioritize the most relevant information in the default display. Specifically, the most critical data will be shown by default, while supplementary explanatory details will be secondary and only accessible when needed. This design approach aims to reduce cognitive load, ensuring that clinicians can quickly access key insights without being overwhelmed by excessive information. By minimizing cognitive burden, this method enables more efficient and personalized analysis.
6.1.5. Integrating Related Data into a Unified System Interface
In our formative study, clinicians mentioned using GAD-7 and PHQ-9 to monitor patients’ anxiety and depression. Additionally, as Section 5.2.2 mentioned, clinicians noted that they use these questionnaires to patients again during the design interface evaluation, which makes us realize that we could integrate all the information clinicians want to see into the same system interface. Based on this, we propose adding a questionnaire results module to the system interface, displaying patients’ completed questionnaires before their clinical visits. This integration would allow clinicians to access all relevant data in a single interface before a consultation, saving time, improving data review efficiency, and enabling a faster and more evidence-based assessment of patients’ mental health status. Importantly, these questionnaire results differ from the anxiety and depression statuses identified in the self-reported symptoms module. They provide a more structured and standardized assessment. By combining subjective self-reported symptoms with objective questionnaire scores, clinicians can gain a more comprehensive understanding of patients’ mental health conditions, which would support more confident mental health-related decision-making, even when patients are outside the clinical setting.
6.2. Visualization of Refined System Design
In this version, we considered the possibility of seamlessly integrating the system design with existing electronic health record (EHR) systems in the future. Therefore, we referenced the design of current EHR systems to ensure consistency. To align with the visual and interaction patterns of EHR systems, we incorporated a patient list on the left side of the interface. A red dot next to a patient’s name indicates a potential risk of mental health sequelae, making it easier for clinicians to identify at-risk patients at a glance. The final system design consists of five modules, which aim to enhance system usability and support clinical decision-making. In the following sections, we will explain the data, functionality, potential interactions, and purpose of each module in detail. The final system design interface is shown in Fig. 2.
Figure 2. The final system design is based on feedback from concussion clinicians gathered through design interface evaluation: (A) patient basic information, (B) patient sleep and physical activities, (c) patient self-report symptoms, (D) patient questionnaire result, and (E) AI risk score prediction for mental health sequelae.
6.2.1. The Basic Patient Information Module
To better integrate seamlessly with the existing EHR system, we have added a basic patient information section in the upper-left module of our system (Figure 2A). This section includes key details such as the date of injury, the number of days since the concussion occurred, the youth patient’s mental health history, and medication usage. These pieces of information provide clinicians with essential insights into the patient’s condition, including the time elapsed since the injury and their mental health background. In particular, the latter two details—mental health history and psychiatric medication usage—are important for concussion clinicians when detecting patients’ mental health sequelae.
6.2.2. The Sleep and Physical Activities Module
In this module (Fig.2 B) concussion clinicians can hover over the question mark icon next to the module title to view its data source, which comes from wearable devices. Below the title, there are options to view data from the past week or the most recent day, with the default display set to one week of data. The module includes three selectable tabs: Sleep Data, Physical Activities, and Recommendation Compliance. The displayed information updates dynamically based on the selected tab. In the Sleep Data tab, clinicians can access heart rate, nap duration, sleep and wake-up times, and total sleep hours. They can select one or multiple metrics based on their specific needs. Different data types are displayed in distinct colors to help clinicians visually distinguish between them more easily. When hovering over a data point in the chart, the exact value for that day is displayed, providing detailed insights into the patient’s sleep patterns.
6.2.3. The Self-Report Symptoms Module
On the right side of the module (Fig.2 C) title, concussion clinicians can hover over an icon to view the data source, which comes from CAs. The module provides two options for selecting a time range: the past week and the last 24 hours. Below the time selector, the interface is divided into two sections. On the left side, there are two tabs: one displaying VA-patient conversation history and the other showing VA-parent conversation history. Clinicians can expand the conversation panel by clicking the button in the top right corner to focus on historical dialogues. On the right side, the Symptoms Reported section visualizes symptom-related information. If the patient or parent has mentioned any symptoms, a colored indicator dot appears next to the corresponding symptom category, with the color reflecting severity—red for severe symptoms, and orange for moderate symptoms. If the patient reports no symptoms, the corresponding dot turns green. To the right of each category, the symptom duration is displayed, indicating how long the reported issue has persisted. Clinicians can also click AI Summary to view a detailed breakdown of the category, providing additional context to support their assessment. At the bottom of the Symptoms Reported section, we provide call buttons for contacting the patient or parent, which ensures that clinicians can quickly intervene in case of an emergency.
6.2.4. The Questionnaires Result Module
In this module (Fig. 2 D), the title includes an option to view data source information, which comes from patient-administered questionnaires at home. These questionnaires include GAD-7 for anxiety, PHQ-9 for depression, and PCSS for concussion severity. The middle section of the chart presents the scores for each questionnaire along with corresponding explanations to help clinicians interpret the results. The module also displays a comparison with previous scores, providing clinicians with a sense of the patient’s recovery trend. Additionally, clinicians can click ”Show Details” to view a breakdown of individual section scores within each questionnaire for a more in-depth assessment.
6.2.5. The AI Risk Score Prediction Module
The final module (Fig. 2 F) presents the AI-generated prediction of a patient’s risk of developing mental health sequelae. Concussion clinicians can hover over the question mark icon next to the title to view information about the source of the risk score. Based on the current predictive capabilities of the AI model, this module offers four different time range options, with the default set to risk predictions within four weeks. Below the time selector, clinicians can view the risk score chart for the selected period and hover over specific time points to examine the risk score at each point. At the bottom of the module, we display the features contributing to the risk score, along with the influence of each contributing factor, represented as a percentage. Visually, if a feature has a high contribution to the risk score, we provide a visual alert. The dot next to the feature title changes color based on the severity of the symptoms associated with that feature, with red indicating severe symptoms and orange indicating moderate severity. This severity level is also reflected in the percentage value on the right, using corresponding colors to enhance clarity. Additionally, the visual representation of each contributing module is color-matched to its corresponding section in the UI. This alignment helps clinicians intuitively understand the connections between different modules and their impact on the AI risk score.
7. Discussion
In this work, we identified multiple challenges that concussion clinicians face when treating youth concussion patients with mental health sequelae, including the difficulties of collecting patients’ information outside the clinics, communicating with patients and their families, as well as making timely clinical decision-making about concussion treatment based on mental health symptoms. Further, we proposed multiple AI-RPM technologies accompanied by AI prediction models to support concussion clinicians’ clinical workflow and iteratively refined a clinician-oriented EMR dashboard user interface to consolidate detailed feedback on information representations, interactions, customizability, and multiple responsible AI considerations. In this section, we delve into how the aforementioned challenges arise and explore how they can potentially be addressed through AI-RPM technologies. Additionally, we discuss the role of AI-RPM systems in supporting clinicians’ decision-making processes. Finally, we talk about concerns regarding responsibility in the potential use of AI-RPM systems and identify design opportunities to tackle these challenges.
7.1. Remote Monitoring of Patient Mental Health-Related Nonclinical Data
7.1.1. Concussion Clinicians’ Need of Monitoring Patients’ Mental Health Symptoms
A unique challenge faced by concussion clinicians, which was identified in our work, is that clinicians find it critical but difficult to assess the severity of patients’ mental health symptoms when they are outside the clinic. For concussion patients, there is a significant overlap between concussion symptoms and mental health symptoms, and these two illnesses could influence each other’s recovery progress. Although concussion clinicians do not possess expertise in the diagnosis and treatment of mental health issues, their clinical decision-making for concussion treatment is closely correlated with patients’ mental health symptoms. If patients exhibit severe mental health symptoms, concussion clinicians need to adjust the concussion treatment plan and the scheduling for follow-up visits accordingly and refer the patients to mental health specialists for professional intervention.
In clinical scenarios, it is not uncommon to have multiple illnesses intertwined with each other. For instance, a recent work discovered that cancer oncologists need to be constantly aware of patients’ cardiotoxicity during cancer treatment and collaborate with cardiologists to adjust the cancer treatment plan (Wu et al., 2024). Nevertheless, these two scenarios are fundamentally different. Based on our formative study, concussion clinicians do not collaborate with mental health specialists, and neither do they diagnose mental health disorders. Concussion clinicians, in most of their clinical practices, only need to detect whether patients’ mental health issues surpass a certain level of severity threshold. If the mental health issues are severe, concussion clinicians will refer patients to mental health specialists for mental health intervention. Otherwise, concussion clinicians will not do the referral but take the patient’s mental health status as a factor in the design of concussion recovery recommendations.
We believe the integrated design of leveraging AI-RPM technologies to collect patients’ mental health-related data outside the clinic and then developing AI-based predictive models to assess patients’ mental health symptoms could provide actionable insights into concussion clinicians’ current workflow. As suggested by Bennett et al. (2012); Zhang et al. (2024), clinicians expect AI predictive models to provide actionable insights, which is critical to ensure AI models’ usefulness in clinicians’ workflow. Future research could further explore the design of AI-RPM technologies with respect to what other modalities of patients’ data could be beneficial for mental health assessment and further develop AI models to integrate such data, along with contextual information like medical histories and behavioral patterns for more reliable predictions with sufficient explainabilities.
7.1.2. Remote Monitoring of Objective Patient Data
Our study revealed significant challenges in how concussion clinicians assess mental health issues in youth concussion patients. Youth patients often avoid openly discussing mental health concerns due to the stigma of being labeled with mental health disorders. Consequently, clinicians are often limited to collecting patient information through mental health assessment questionnaires or direct observation of patients’ behaviors during clinical visits. However, if patients are unwilling to communicate openly, the validity of questionnaires can be compromised, and observation during clinical visits can be left as the only viable option. However, direction observation is inherently limited from the geographical perspective. Once patients leave the clinic, concussion clinicians can hardly track patients’ mental health status in a timely and effective manner. Even if patients are willing to faithfully and openly discuss their mental health status, patients’ memories could be fragmented and inaccurate, especially for sleeping patterns.
We suggest that RPM technologies, particularly wearable devices like smartwatches, could help address the aforementioned limitations by providing clinicians with real-time access to key physiological metrics. In particular, these physiological metrics are objective measurements that can faithfully reflect patients’ mental health status. Prior research demonstrated the feasibility and advantages of wearable in monitoring objective patient information outside the clinic (Mendel et al., 2024; Frazier et al., 2023; Su et al., 2019; Wu et al., 2024). Wearable devices could monitor objective data such as sleep patterns, physical activity levels, and resting heart rate, which are critical signals for tracking and assessing patients’ mental well-being. For instance, elevated anxiety levels may be inferred through reduced physical activity or disrupted sleep patterns, while a stable resting heart rate could indicate recovery progress. Concussion clinicians could use such data for clinical decision-making, like scheduling timely follow-up visits or initiating referrals to mental health specialists as needed.
An additional benefit provided by wearable devices is that they can monitor patients’ status non-intrusively to be seamlessly integrated into patients’ daily routines. This non-intrusive characteristic minimizes potential cognitive burdens added to the patients, where patients will only need to remember to charge the wearables regularly.
7.1.3. Remote Monitoring of Subjective Patient Data
Clinicians in our study emphasized that the severity and duration of symptoms such as anxiety, depression, or even suicidal ideation are critical factors in assessing whether youth concussion patients are experiencing mental health sequelae. However, these symptoms are more of subjective experiences that can hardly be captured using wearables because the data can be collected by wearables are standardized objective physiological measurements.
To address this gap, we propose integrating advanced generative AI with RPM technology to gather patients’ subjective data through text-based or voice-based conversational assistants (CAs). These CAs are supported by large language models (LLMs) that could engage in free-form natural language conversations with patients. Compared with traditional RPM approaches, natural conversation significantly reduces the requirement of patients’ technical literacy to engage with the devices, enables flexible patient self-reporting, and can be further designed to conduct regular health check-ins through pre-set question lists. Recently, a number of works have demonstrated the promising opportunities of LLM-based CA to support various patient populations in healthcare scenarios (Bartle et al., 2023, 2022; Wang et al., 2023a; Simpson et al., 2020; Ma et al., 2024; Yang et al., 2023), for example, to support older adults with chronic diseases (Yang et al., 2024).
Using AI-driven analytics, key psychological risk factors such as suicidal thoughts could be automatically detected. LLMs could then generate a summary report of the patient’s mental health for the day and transmit it remotely to the clinicians’ EMR system. By combining subjective data from CAs with objective information collected by wearables, clinicians could gain a more comprehensive understanding of the patients’ mental health in their daily lives. Subjective information often provides contextual insights for objective data, which enables clinicians to make more accurate judgments. For example, if wearable data shows that a patient had very low physical activity on a given day, the CA might reveal through its conversation that the patient’s mental health was stable and the low activity was due to a family gathering at home. This additional context would allow the clinician to assess the situation more accurately, avoiding unnecessary concern over what might otherwise appear to be a worrisome pattern in objective metrics. By leveraging the strengths of both data types, our proposed AI-RPM approach could enhance the quality and precision of mental health assessments in youth concussion care.
7.1.4. Patient Privacy Consideration With AI-RPM Technologies
Lastly, when designing such an AI-RPM system, it is essential to underscore protecting patient privacy and ensuring data security. This issue has been widely discussed in many studies and is recognized as a core challenge in AI system design (Gawankar et al., 2024; Bala et al., 2024; Korobenko et al., 2024). However, in our research context, the target population comprises youth patients aged 11–17, whose privacy needs are uniquely sensitive. The privacy of youth patients is not only a concern for the patients themselves but is also critical to their parents, who often play a key role in the medical process. Parents typically expect to have transparency and informed consent regarding how their child’s data is collected and used (Sisk et al., 2020; Haley et al., 2024). Therefore, finding ways to protect patients’ privacy while building and maintaining parental trust in the data collection and usage process is a critical issue that requires careful consideration and balance. Future research could explore methods to effectively safeguard youth health data while simultaneously building trust and participation among parents.
7.2. Multi-Stakeholder Collaboration in the Design of Personalized and Accountable AI-RPM Systems
7.2.1. Collaboration With Clinicians in Supporting Clinical Workflow
Our findings highlight that while clinicians are open to using AI-RPM systems, they are concerned about the potential for information overload. Clinicians are already overwhelmed by the variety of data and tasks within existing EHR systems. Introducing another AI system with additional data could significantly increase their workload, potentially reducing the number of available weekly appointment slots and further straining already limited healthcare resources. In our study, clinicians generally believe they do not have the time to process such detailed information in their daily work. However, clinicians still expressed willingness to check the data collected by the AI-RPM systems before each clinical visit because they believe such data could support them better prepare for patient consultations.
To address this concern, AI-RPM systems need to be designed in collaboration with clinicians to prioritize personalized and context-relevant data presentation. One approach to achieving this is through layered information presentation, where critical data is prioritized and displayed prominently, while secondary information remains accessible but hidden by default. For example, clinicians could be shown a summary of essential metrics, such as daily activity levels, sleep patterns, and key symptom trends, while additional context or raw data could be available on demand (e.g., through interactive elements on the interface). Designers must collaborate closely with clinicians to determine which data points should be prioritized and how they should be presented to support efficient clinical decision-making.
Clinicians in our study also emphasized the importance of defining thresholds for symptom severity, particularly in distinguishing between mild, moderate, and severe symptoms. Some concussion clinicians mentioned that they might not be particularly concerned about mild anxiety, but severe or prolonged symptoms would require their attention. Without clear definitions, AI systems may either underreport or overemphasize symptoms, leading to delayed interventions or unnecessary concerns. Clinicians are the domain experts who can help designers bridge the gap with their knowledge and expertise. By working with clinicians to establish these thresholds, AI-RPM systems can more effectively categorize symptoms, ensuring that clinicians are alerted only when necessary. Such a collaborative approach will not only reduce cognitive overload but also build trust and confidence in the system’s reliability, which will allow clinicians to rely on the AI system as a supportive tool rather than a burden.
Ultimately, the goal of engaging concussion clinicians in the collaborative design process is to enable the system for clinicians to access the most relevant patient information quickly, improve the efficiency of their workflows, and make informed decisions without being overwhelmed by data. By integrating clinicians’ preferences and needs into the system design, AI-RPM systems can enhance their usability and effectiveness, promoting better human-AI collaboration in clinical practice.
7.2.2. Collaboration with Parents in Managing Emergency Situations
In the proposed design section of self-reported symptoms, there is a ”suicidal ideation” alert module which clinicians believe would be of significant benefit to alert clinicians for any unexpected emergency. However, clinicians expressed considerable concerns about this feature. Clinicians questioned the accuracy of the ”suicidal ideation” alerts, not solely in terms of the technical capabilities of the AI system but more importantly regarding the reliability of the self-reported data. In our research context, clinicians pointed out that at the developmental stage, youth patients may often express emotions with exaggeration. For example, youth patients might say things like ”I want to hurt myself” or ”I have suicidal thoughts,” which might not necessarily reflect their actual intentions. Such statements could result in a false alarm when captured by the AI system and misclassified as a ”high-risk alert.” The implications of such false alarms are multifaceted.
First, responding to a high-risk alert would require clinicians to dedicate considerable time and resources to contact the patient and address the situation, even if it turns out to be unwarranted. False alarms could lead to a waste of a clinician’s valuable time and critical medical resources that could be allocated to other patients. Second, if false alarms occur frequently, clinicians could gradually lose trust in the system, which eventually leads to the abandonment of such a system. Another concern expressed by clinicians is the ambiguity of responsibility that AI systems introduce in medical decision-making. Previous studies have highlighted the phenomenon where clinicians, as final decision-makers, have full responsibility for medical outcomes, while AI systems themselves are not held accountable. This issue was also raised in our study. Specifically, if the AI system triggers an emergency alert (e.g., indicating a patient might have suicidal ideation), clinicians face multiple constraints in responding. These constraints might include a lack of resources or time to address the alert promptly, or situations where clinicians fail to review the AI system’s information in a timely manner. Any subsequent negative outcome, such as a life-threatening incident, raises the question of who should be held accountable.
We believe that in a high-risk and high-uncertainty medical scenario where the patients are young, it’s a key challenge to ensure the factuality of emergency alerts and find a balance between handling alerts and clinicians’ responsibilities throughout the development lifecycle of an AI system. Addressing this challenge requires collaboration among a broader set of stakeholders. For instance, involving the youth patient’s family in handling emergencies could play a crucial role. If the AI system detects an emergency, such as suicidal ideation, the system should notify not only the clinician but also the patient’s family immediately because family members could take appropriate action in cases where the clinician cannot respond promptly. For example, family members could calm the patient, take them to a healthcare facility, or contact mental health professionals for intervention. Additionally, clinicians who receive such alerts could collaborate with the family by providing timely professional guidance. Such notification and collaboration mechanisms would not only enable families to gain a more comprehensive understanding of the patient’s condition but also reduce risks arising from delayed clinical responses. Engaging parents in the emergency care-providing process ensures patient safety while alleviating some of the pressure on clinicians who might be constrained by limited time or resources.
To address ambiguity in responsibility for medical decisions, collaborative efforts involving hospital administrators, legal experts, and researchers are needed to develop clear guidelines on accountability in AI-supported clinical scenarios. Stakeholders should clearly define roles and accountability in emergencies triggered by AI systems to address the potential liability concerns highlighted in existing studies (Goodman and Flaxman, 2017). An addition benefits would be helping clinicians feel more secure when using AI systems in their clinical workflow.
Future research should explore mechanisms to promote effective collaboration between clinicians, parents, AI system developers, HCI designers, and other stakeholders. The multi-stakeholder collaboration will be essential for designing accountable and trustworthy AI-RPM systems that can handle high-risk scenarios while ensuring patient safety and clinician confidence.
7.3. AI as a Collaborative Decision-Making Partner For Patient Care Outside the Clinic
Our study highlights that concussion clinicians found certain AI-supported metrics—such as physical activity, sleep patterns, and nutritional intake—useful. The metrics about physical activities serve as critical indicators that clinicians interpret in relation to potential mental health sequelae. For instance, disruptions in sleep or drastic reductions in physical activity may signal the need for closer monitoring or further evaluation of youth concussion patients. However, some clinicians expressed reluctance to trust AI-generated probabilistic scores unless they were accompanied by high levels of accuracy and interpretability. Instead, they prefer using the AI system as a data collection and consolidation tool but leave clinical decisions to human experts.
This finding emphasizes the importance of designing AI-powered systems that do not attempt to replace clinical judgment but rather complement it by providing structured and reliable data that clinicians need to support their decision-making. As identified in previous CSCW studies on human-AI collaboration in high-stakes environments, trust and interpretability are key components influencing whether end-users effectively adopt AI systems in their workflows (Liao et al., 2020; Amershi et al., 2019). Our study explores how these principles apply in the specific context of youth concussion care, where clinicians balance multiple sources of uncertainty, patient variability, and time-sensitive decisions with respect to patients’ mental health conditions.
Our findings resonate with ongoing discussions in the CSCW community regarding the design of collaborative systems that support human decision-making while preserving user agency. Existing works emphasize that human-centered AI systems should aim to act as “intelligent assistants” rather than autonomous decision-makers (Yang et al., 2019; Zhang et al., 2020). In particular, enabling users to maintain control over critical decisions is critical in clinical settings where over-reliance on AI-generated recommendations could lead to adverse outcomes if predictions are incorrect or poorly contextualized. Our work contributes new insights on how AI-RPM systems can be effectively integrated into clinical workflows without introducing additional cognitive burdens or reducing clinician autonomy. Specifically, our study extends the understanding of human-AI collaboration by demonstrating how AI can facilitate patient care outside the clinic through continuous monitoring of non-clinical patient data to bridge the long-existing patient care gaps between clinical visits. While prior HCI and CSCW research has explored AI’s role in inpatient or acute care settings (Liao et al., 2020), we shift the focus to outpatient care and remote monitoring as an emerging important healthcare area.
Our study highlights the need for AI system design approaches that prioritize transparency, explainability, and collaboration across multiple stakeholders. Our findings suggest that AI-generated insights should be designed as part of an interactive process, where clinicians can not only access key metrics but also contextualize them with their broader understanding of a patient’s condition. For instance, future work could investigate interactive dashboards that allow clinicians to adjust or refine AI predictions based on their observations, creating a feedback loop between human expertise and machine learning outputs. The human-AI collaboration paradigm should not be a static and standardized format, instead, such collaboration should be seen as dynamic and iterative (Muller and Weisz, 2022; Fragiadakis et al., 2024).
For the information presentation strategies to support clinicians’ professional workflow, our study highlights the potential for collaborative data presentation strategies that tailor information to the specific needs of different stakeholders, including clinicians, patients, and caregivers. As suggested by prior HCI and CSCW studies on personalized information delivery (Amershi et al., 2019), layering information hierarchically and providing context-specific views can reduce information overload while ensuring that users have access to relevant data when needed. In the context of AI-RPM systems, we propose to display summary statistics for routine monitoring while offering detailed visualizations or explanations for cases requiring further investigation.
Our study not only provides practical recommendations for designing AI-RPM systems but also contributes to broader CSCW discussions on human-AI collaboration in high-stake and high-risk domains. Future research could explore collaborative design methodologies that involve clinicians, patients, and caregivers throughout the development process, ensuring that AI systems meet the diverse needs of all stakeholders.
Moreover, addressing trust concerns and ensuring interpretability remain key areas for future exploration (Barnes and Hutson, 2024; Mendel et al., 2024; Prahl and Van Swol, 2017; Dietvorst et al., 2015; Wang et al., 2019). The CSCW community could investigate how interactive system features, such as real-time feedback or explanatory interfaces, influence user trust and decision-making in long-term monitoring scenarios. Additionally, as AI-RPM systems become more prevalent, interdisciplinary collaborations involving legal, ethical, and technical perspectives will be necessary to establish guidelines for responsible deployment and use.
8. Limitations and Future Work
This study identifies the challenges and needs of concussion clinicians when detecting mental health sequelae among youth concussion patients. It also presents design considerations and a proposed UI to support clinicians in this process. However, the study has several limitations. First, the sample size of our participants was limited. Six concussion clinicians participated in both the formative study and heuristic evaluation. However, recruiting these experts was challenging due to their demanding workloads (Wang et al., 2021; Jin et al., 2020) and the specialized expertise required to diagnose youth concussion patients with mental health sequelae. As a result, we conducted our study with highly relevant experts in this specific domain. Previous research has also employed a similar number of expert participants in related studies (Cai et al., 2019b; Beede et al., 2020; Jacobs et al., 2021; Yang et al., 2024; Zhang et al., 2024). Additionally, our recruited clinicians came from different hospitals and regions across the United States, enhancing sample diversity. Future research should expand the participant pool to include a broader range of clinicians across different geographic areas to improve the generalizability of the findings. Second, the final UI design was designed based on feedback from concussion clinicians during the heuristic evaluation. However, the final design has not yet been evaluated by concussion clinicians to assess its effectiveness and usability. Future research should incorporate participatory design methodologies to further refine the UI, ensuring it aligns more closely with clinicians’ workflows and needs. Third, our study focused on UI design rather than developing an interactive prototype. The primary goal of our UI design was to explore the feasibility and provide insights for further design refinements before actual development, thereby saving resources and time for both researchers and clinicians. Future research should investigate clinicians’ pain points and needs when interacting with a functional system developed. Finally, our study scenario specifically examined the detection of mental health sequelae after concussions among youth and explored the potential of AI-RPM in this context. Future studies could explore the broader application of AI-RPM in various clinical settings and understand how AI-driven systems can assist clinicians in early detection and intervention across different medical domains.
9. Conclusion
We explored the challenges and needs faced by concussion clinicians related to mental health issues in youth concussion patients and proposed design considerations for the use of AI-RPM technologies in this high-risk, high-stakes context. Our study makes two key contributions to the CSCW field. First, we identified three core challenges clinicians encounter when collecting and monitoring mental health-related information and making mental health-related decisions: (1) collecting patients’ mental health information outside the clinic; (2) communicating effectively with patients and their family members; and (3) assessing patients’ compliance with clinical recommendations. Second, we proposed a series of opportunities for AI-RPM technologies to address these challenges, as well as design considerations for AI-powered systems to collect multimodal, non-clinical data that supports clinicians in making informed mental health-related decisions. To our knowledge, this is the first study in the CSCW field to focus on the intersection of concussion clinicians and youth concussion patients in the context of mental health-related decision-making, providing a foundation for designing AI-RPM systems in similar scenarios.
Appendix A APPENDIX: INTERVIEW SCRIPTS FOR FORMATIVE STUDY
We begin with a brief introduction of our research and study procedure.
Background
(1) Could you please tell me your years of practice, department, and specialization?
(2) How many patients do you see each year? how many of them have mental health symptoms? How do you handle them (maybe they don’t need to refer)?
A Recent Experience
(1) Can you think of a recent time when you had a young concussion patient (11-17 years old) and thought they might have mental health issues? How did you find out?
Information Collection and Monitoring
(1) (If not mentioned in the previous question) For that patient in your story, before the final referral, what symptom-related information makes you think it might be a mental health issue? How did you deal with it?
(2) (if they mention any at-home symptoms) What kind of symptom-related information do you find helpful when patients are resting at home (smartwatch and smart speaker)? (if not mentioned) Do you need any at-home symptoms? (e.g., patients’ mental status or sleep patterns at home )
(3) (if not mentioned) How do their parents get involved in this process?
Challenges and Workaround Strategies
(1) What challenges do you have when it comes to collecting symptom-related information from patients, managing that information, or making mental health decisions (e.g., further mental health testing or referring to a mental health expert)?
Technology (Expectations and Concerns)
(1) Do you use any existing tools or features in EHR (e.g., EPIC) in the youth patients’ mental health decision-making? If so, what are they and how do you like them?
Appendix B APPENDIX: INTERVIEW SCRIPTS FOR DESIGN INTERFACE EVALUATION
(1) This section (Fig.1) shows the probability of mental health sequelae within the next 15 days with a range from 0% to 100%, where the higher probability score signifies a higher probability of experiencing mental health sequelae. Feature importance scores show the contributing factors to the probability. The risk score chart below shows the risk scores for the past ten days, today, and the next 15 days. We’d love your comments. What do you think about it? What other information or visualization would you like to see?
(2) (Less familiar, need extra emphasis) In this section, we will use a smart speaker to have conversations with patients. The conversation history is shown on the right. On the left is the symptom overview section. If patient-reported symptoms toward a particular category, the dot for that category will change from green to red as a highlight for your review. You can click on the category name or use the search bar to type in keywords, then the corresponding conversation content will be shown on the right. What do you think about it? What other information or visualization would you like to see?
(3) (More familiar, need less time) The data in this section is collected from patients’ smartwatches. You can click on other data buttons to view corresponding recorded data in detail. What do you think about it? What other information or visualization would you like to see?
(4) Above these three sections, what other information would you need to help make mental health decisions?
Closing Question (1) Is there anything else you’d like to share or a final question to ask us? Let us know or reach out via follow-up email.

### v2

@@
 Subsequently, we proposed preliminary clinician-facing interface designs with the integration of AI-based RPM technologies (AI-RPM), followed by design evaluation sessions with highly professional concussion clinicians.
-Clinicians underscored the value of integrating multi-modal AI-RPM technologies to support clinicians’ decision-making while emphasizing the importance of customizable interfaces with explainability and multiple responsible design considerations.
+Clinicians underscored the value of integrating multi-modal AI-RPM technologies to support their decision-making while emphasizing the importance of customizable interfaces through collaborative design and multiple responsible design considerations.
 Youth concussion, mental health sequelae, clinical decision-making, remote patient monitoring, artificial intelligence
 copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing Human computer interaction (HCI)$c$s: Human-centered computing HCI design and evaluation methodsccs: Human-centered computing Collaborative and social computing theory, concepts and paradigms
+Figure 1.
+Refined AI-RPM system designed to support concussion clinicians in mental health monitoring: (A) patient list, (B) patient basic information, (C) sleep and physical activities, (D) self-report symptoms, (E) questionnaire result, and (F) AI risk score prediction for mental health sequelae.
 1. Introduction
 […]
-The standard of care (SoC) approaches of concussion clinicians include (1) asking patients to recall their at-home behaviors during clinical visits (Silverberg et al., 2020; Brent and Max, 2017) and (2) instructing patients or parents to complete self-evaluation questionnaires at home, such as Generalized Anxiety Disorder-7 (GAD-7) and Patient Health Questionnaire-9 (PHQ-9) (Carson et al., 2021; Silverberg et al., 2020).
-However, these approaches are not without limitations: it could be challenging for concussion patients to comprehensively and accurately recall their previous behaviors in a weekly clinic visit (Frissa et al., 2016; Van den Bergh and Walentynowicz, 2016); patients may be able to complete questionnaires on a regular basis at home, but concussion clinicians can only access the questionnaire results when patients return to the clinics in person, which still delays clinicians’ decision-making.
-Previous work has proposed different Remote Patient Monitoring (RPM) technologies to collect patients’ at-home information in various scenarios (Chung et al., 2016; Olen and Dechert-Crooks, 2017; Barnes et al., 2024; van Ierssel et al., 2023; Toresdahl et al., 2021; Tirosh et al., 2024) For example, mobile health (mHealth) applications have been proposed in which patients can report their physiological measurements to care provider teams in real-time.
-Wearable technologies, such as the ActiGraph (Yang et al., 2020) and Philips Actiwatch (Donahue et al., 2024), have been explored to measure the daily steps and sleep patterns of adult concussion patients.
-Despite their relative success in capturing patients’ activity data, wearable-based RPM approaches are not without limitations: mHealth applications could lead to cognitive burdens for patients with low digital literacy and brain injuries (Giebel et al., 2024; Purdy, 2023); extensive amount of time and effort is required to possess the massive amount of multi-modal data collected by wearable devices (Ginsburg et al., 2024; Baig et al., 2017).
-These limitations could result in low compliance among the patients and delays in accessing collected data from clinicians, which further affect timely diagnosis and intervention.
+The standard of care (SoC) approaches of concussion clinicians include: (1) asking patients to recall their at-home behaviors during clinical visits (Silverberg et al., 2020; Brent and Max, 2017), and (2) instructing patients or parents to complete self-evaluation questionnaires at home, such as Generalized Anxiety Disorder-7 (GAD-7) and Patient Health Questionnaire-9 (PHQ-9) (Carson et al., 2021; Silverberg et al., 2020).
+However, these approaches are not without limitations: it could be challenging for concussion patients to comprehensively and accurately recall their previous behaviors in a clinic visit (Frissa et al., 2016; Van den Bergh and Walentynowicz, 2016); patients may be able to complete questionnaires on a regular basis at home, but concussion clinicians can only access the questionnaire results when patients return to the clinics in person, which still delays clinicians’ decision-making.
+Previous work has proposed different Remote Patient Monitoring (RPM) technologies to collect patients’ at-home information in various scenarios (Chung et al., 2016; Olen and Dechert-Crooks, 2017; Barnes et al., 2024; van Ierssel et al., 2023; Toresdahl et al., 2021; Tirosh et al., 2024).
+For example, mobile health (mHealth) applications have been proposed in which patients can report their physiological measurements to care provider teams in real time.
+Wearable devices, such as the ActiGraph (Yang et al., 2020) and Philips Actiwatch (Donahue et al., 2024), have been explored to measure the daily steps and sleep patterns of adult concussion patients.
+Despite their relative success in capturing patients’ activity data, these traditional RPM approaches are not without limitations: mHealth applications could lead to cognitive burdens for patients with low digital literacy and brain injuries (Giebel et al., 2024; Purdy, 2023); extensive amount of time and effort is required to process the massive amount of multi-modal data collected by wearable devices (Ginsburg et al., 2024; Baig et al., 2017).
+These limitations may reduce patient compliance and lead to clinicians delaying collected data access, which further affects timely diagnosis and intervention.
 Recent exploration in the HCI community demonstrates the promising potential of leveraging innovative AI technologies for RPM (AI-RPM) to address the aforementioned limitations of traditional RPM approaches.
-For instance, conversational agents (CAs) powered by large language models (LLMs) enable asynchronous patient-care provider communication to verbally collect patients’ health conditions through regular health check-in conversations (Yang et al., 2024).
-LLMs have also been explored to organize and summarize key information from massive RPM data on clinician-facing dashboard (Wu et al., 2024).
+For instance, conversational agents (CAs) powered by large language models (LLMs) enable asynchronous patient-provider communication to verbally collect patients’ health conditions through regular health check-in conversations (Yang et al., 2024).
+LLMs have also been explored to organize and summarize key information from massive RPM data on clinician-facing dashboards (Wu et al., 2024).
 Meanwhile, another line of research explored using AI prediction algorithms with RPM data to assist clinicians’ decision-making in their clinical workflow (Yin et al., 2024).
 Despite the promising opportunities of AI-RPM systems, they could lead to additional difficulties and burdens if not appropriately designed.
-Recent studies discovered that AI prediction models without faithful and reliable explanations cannot provide helpful information (Antoniadi et al., 2021) or actionable insights to clinicians (Rane et al., 2023; Zhang et al., 2024), lose clinician trust (Zhang et al., 2020; Chanda et al., 2024), and even result in human-AI competition (Zhang et al., 2024) Moreover, in high-stakes, high-risk, and uncertain clinical scenarios, clinicians follow strict clinical guidelines and specifications in their daily workflow.
-AI systems that are in conflict with clinical specifications cannot be used and even cause potential harm to stakeholders (Romero-Brufau et al., 2020).
-In contrast, human-centered AI (HCAI) design guidelines (Shneiderman, 2022; Wang et al., 2023b) underscores the importance of engaging stakeholders in the system design process and understanding stakeholders’ workflow, in particular, the needs and challenges they encounter.
-By following HCI design guidelines, AI systems can be seamlessly integrated into stakeholders’ workflow where stakeholders can confidently understand, trust, and use AI-powered systems in their current workflow.
+Recent studies discovered that AI prediction models without faithful and reliable explanations cannot provide helpful information (Antoniadi et al., 2021) or actionable insights to clinicians (Rane et al., 2023; Zhang et al., 2024), lose clinician trust (Zhang et al., 2020; Chanda et al., 2024), and even result in human-AI competition (Zhang et al., 2024).
+Moreover, in high-stakes, high-risk, and uncertain clinical scenarios, clinicians follow strict clinical guidelines and specifications in their daily workflow.
+AI systems misaligned with clinical specifications may be unusable and even cause potential harm to stakeholders (Romero-Brufau et al., 2020).
+In contrast, human-centered AI (HCAI) design guidelines (Shneiderman, 2022; Wang et al., 2023b) underscore the importance of engaging stakeholders in the system design process and understanding stakeholders’ workflow, in particular, the needs and challenges they encounter.
+By following HCAI design guidelines, AI systems can be seamlessly integrated into stakeholders’ workflow where stakeholders can confidently understand, trust, and use AI-powered systems in their current workflow.
 A number of studies have demonstrated the effectiveness of human-centered AI systems in medical and healthcare scenarios to facilitate the needs of diverse stakeholders, including older adults (Hao et al., 2024), cancer patients (Wu et al., 2024), and clinicians (Wang et al., 2023b; Fogliato et al., 2022; Yang et al., 2019; Zhang et al., 2024).
-Nevertheless, how to employ HCAI guidelines in the design of AI-RPM systems has not been studies yet, and little is known about the challenges concussion clinicians face in remote monitoring youth patients and how AI-RPM technologies could be of practical benefit.
-In this work, our objective is to investigate the current workflow of concussion clinicians with respect to the challenges in remote monitoring and decision making of youth concussion patients.
+Nevertheless, how to employ HCAI guidelines in the design of AI-RPM systems has not been studied yet, and little is known about the challenges concussion clinicians face in remote monitoring youth patients and how AI-RPM technologies could be of practical benefit.
+In this work, our objective is to investigate the current workflow of concussion clinicians with respect to the challenges in remote monitoring and decision-making of young concussion patients.
 In particular, we focus on how clinicians collect patients’ mental health-related information, what decisions they need to make, and what the difficulties are in their current workflow.
 In addition, we want to explore the design opportunities of AI-RPM technologies in supporting clinicians’ needs for remote monitoring and decision-making during their practice.
-In this study, we seek to explore two main research questions (RQs):
-  • RQ1: What challenges and needs do concussion clinicians face in remote monitoring of mental health symptoms and decision making with multi-modal RPM data for youth concussion patients?
-  • RQ2: What are the design considerations needed for AI-RPM technologies to support concussion clinicians in their workflow?
-We approached to address these RQs through a three-stage study design: first, we conducted a formative study of semi-structured interviews with six highly professional concussion clinicians and derived a suite of design considerations of AI-RPM technologies that could support clinicians’ need for remote monitoring and clinical decision-making.
-We came up with preliminary designs of clinician-facing interfaces with different AI-RPM technologies based on design considerations.
-Subsequently, we engaged the same group of clinicians in a design interface evaluation study to further collect feedback and insights on the design considerations of different AI-RPM technologies.
-We consolidate the design feedback and revise the preliminary interface design to come up with a complete clinician-facing system design interface, along with the design considerations, as the final artifacts.
-This paper makes the following key contributions: (1) we identified the challenges and needs that concussion clinicians encounter during their clinical practice with youth concussion patients, particularly with respect to patient mental health sequelae; (2) we derive a suite of design considerations and visualization feedback to empower advanced AI-RPM technologies to support remote patient monitoring and clinicians’ decision-making; and (3) we deliver a user-centered interface design of a concussion clinician-oriented dashboard system supported by multiple AI-RPM technologies, where the initial interface design and refinements are grounded in users’ needs and feedback.
+To answer these questions, we conducted a formative study of semi-structured interviews with six highly professional concussion clinicians and derived a suite of design considerations of AI-RPM technologies that could support clinicians’ needs and clinical decision-making.
+We then came up with preliminary system designs based on design considerations.
+Subsequently, we engaged the same group of clinicians in an evaluation study to further collect feedback and insights on the design.
+Finally, we consolidate the feedback and revise the design to come up with a complete clinician-facing system design as the final artifact.
+This paper makes the following key contributions: (1) we identified the challenges and needs that concussion clinicians encounter during their clinical practice with youth concussion patients, particularly with respect to patient mental health sequelae; (2) we derive a suite of design considerations and visualization feedback to empower advanced AI-RPM technologies to support remote patient monitoring and clinicians’ decision-making; and (3) we deliver a user-centered system design of a concussion clinician-oriented dashboard system supported by multiple AI-RPM technologies, where the preliminary system design and refinements are grounded in users’ needs and feedback.
 2. Related Work
-We review prior research on the negative influence youth have due to the mental health sequelae in post-concussion, AI-RPM technologies in clinical settings, as well as designing AI-powered systems for clinical decision support.
-2.1. Challenges Related to Mental Health Sequelae Following Concussion
-Mental health sequelae, such as anxiety, depression, and suicidality, are common among youth patients (ages 11–17) following a concussion (Stein et al., 2019; Gornall et al., 2021; Ledoux et al., 2022).
-Among these youth concussion patients, 70–80% of them experience longer recovery periods, normally more than 28 days (Scorza and Cole, 2019; Cuff et al., 2022).
-Studies have already shown the importance of integrating the assessment of mental health conditions into clinical follow-up visits (Gornall et al., 2021; Russell et al., 2023).
-Monitoring youth concussion patients’ mental health-related information enables clinicians to timely identify any abnormalities in the patient’s mental health and make mental health-related decisions, such as referring the patient to a mental health expert for further assessment or intervention.
-Therefore, early identification of mental health sequelae in the initial stages of concussion recovery is crucial for reducing the risk of worsening mental health issues in youth concussion patients and minimizing the negative impact on their concussion recovery.
-However, there are challenges when it comes to monitoring youth concussion patients’ mental health.
-Most youth concussion patients are advised to rest at home after a concussion (Silverberg and Mikolić, 2023), their mental health sequelae-related symptoms, such as anxiety, often manifest in their everyday environments, particularly at home, which makes concussion clinicians difficult to detect within clinical settings.
-Although clinicians can inquire about mental health sequelae-related symptoms during clinical visits, recall bias (Frissa et al., 2016; Van den Bergh and Walentynowicz, 2016) may prevent patients from providing complete and accurate information to concussion clinicians.
-Moreover, clinical follow-up visits for youth concussion patients often lack consistency.
-Although it is recommended to conduct follow-ups every three days after a concussion (Agency, 2024), one study showed that actual follow-up rates vary significantly, with the lowest being only 13.2% (Ramsay et al., 2023).
-Insufficient follow-up prevents concussion clinicians from collecting information to accurately assess the youth patient’s mental health sequelae timely, making it difficult to make mental health-related decisions such as conducting further assessments or making a referral to a mental health expert.
-Therefore, it is essential to remotely monitor and collect behavioral data from youth concussion patients outside of clinical settings to assist clinicians in the early detection of mental health sequelae.
+Based on the prior research, we first identified the challenges related to mental health sequelae after a concussion.
+Then, we provided a brief overview of current AI-RPM technologies in clinical settings, as well as AI-based system design for clinical decision support.
+2.1. Challenges in Post-Concussion Mental Health Sequelae
+Mental health sequelae, such as anxiety, depression, PTSD, and suicidality, are common among youth concussion patients (ages 11–17) (Stein et al., 2019; Gornall et al., 2021; Ledoux et al., 2022).
+Yang et al.
+(Yang et al., 2015) studied how concussions impacted the mental health of 71 collegiate athletes and found that 14 reported symptoms of depression, 24 experienced symptoms of anxiety, and 10 reported both depression and anxiety.
+Another study showed that youth concussion patients tend to have twice the risk of suicide compared to their peers who have never had a concussion (Fralick et al., 2019).
+Moreover, mental health sequelae can prolong youth concussion patients’ concussion recovery (Silverberg et al., 2020; Iverson and Lange, 2003), and hinder their ability to attend school and participate in normal social activities (Iadevaia et al., 2015), which can be an important way to maintain mental well-being (Zamfir, 2015).
+Thus, it is important to integrate the assessment of mental health conditions into clinical follow-up procedures (Gornall et al., 2021; Russell et al., 2023).
+The assessment enables concussion clinicians to identify abnormalities in youth patient mental health conditions and conduct intervention, such as conducting a further assessment or referring the patient to a mental health expert.
+However, there are challenges when it comes to accessing youth concussion patients’ mental health conditions.
+Most youth patients are advised to rest at home after a concussion (Silverberg and Mikolić, 2023).
+Their mental health symptoms, such as anxiety, changes in sleep patterns, and disinhibition (Brent and Max, 2017), often manifest outside of clinics, which makes it difficult for concussion clinicians to observe these symptoms in clinical settings.
+Although clinicians can inquire about mental health symptoms during clinical visits, recall bias (Frissa et al., 2016; Van den Bergh and Walentynowicz, 2016) may prevent patients from providing complete and accurate information to concussion clinicians.
+Self-report questionnaires like the GAD-7 and PHQ-9 are available for anxiety and depression screening (Brent and Max, 2017).
+However, clinicians often do not receive the results until the patient’s next clinical visit.
+Moreover, clinical follow-up visits for youth concussion patients are often insufficient and lack consistency.
+Although it is recommended to conduct follow-ups every three days after a concussion (Agency, 2024),  Ramsay et al.
+(2023) found that actual follow-up rates vary significantly, with the lowest being only 13.2%.
+Insufficient follow-up prevents concussion clinicians from timely collecting mental health information and assessing the youth patient’s mental health conditions.
+Accordingly, it’s difficult for concussion clinicians to make timely mental health-related decisions.
+Therefore, novel approaches are needed to support clinicians in remote monitoring the mental health symptoms of youth concussion patients on time.
 2.2. AI-RPM Technologies in Clinical Settings
-Remote Patient Monitoring (RPM) includes primarily wearable devices and home-use medical equipment that collect patient health data outside of clinical settings (Mendel et al., 2024).
+RPM includes primarily wearable devices and home-use medical equipment that collect patient health data outside of clinical settings (Mendel et al., 2024).
 These systems transmit monitored patient data to healthcare providers for real-time monitoring and assessment, thereby supporting clinicians in their decision-making (Catalyst, 2018; Baig et al., 2017; of Health et al., 2008).
@@
 (2020) utilized ActiGraph, a smartwatch-based device, to monitor patients’ physical activity and collect movement-related data, such as step count.
-Furthermore, researchers in the Human-Computer Interaction (HCI) and Computer-Supported Cooperative Work (CSCW) fields have started exploring mobile health (mHealth) technologies, which use mobile devices to collect patient information, facilitate self-management, and enhance remote communication between patients and clinicians (Corwin et al., 2024; Nyapathy and Arriaga, 2019; Salamah et al., 2021; West et al., 2018).
+Furthermore, researchers in the Human-Computer Interaction (HCI) and Computer-Supported Cooperative Work (CSCW) fields have started exploring mHealth applications that facilitate self-management, and enhance remote communication between patients and clinicians (Corwin et al., 2024; Nyapathy and Arriaga, 2019; Salamah et al., 2021; West et al., 2018).
 For instance, El-Rashidy et al.
@@
 RPM systems can display collected data via user interfaces for patient review (Griggs et al., 2018) or integrate the data into electronic health record (EHR) systems (Dinh-Le et al., 2019).
-However, given the significant time constraints that clinicians already face (Häikiö et al., 2020), along with their existing workload, the additional data generated by RPM systems could further exacerbate these challenges.
-With recent advances in AI, AI-driven tools have the potential to assist clinicians by analyzing and summarizing large volumes of patient data and presenting information in a more interpretable format, thus reducing the cognitive load in clinical workflows (Häikiö et al., 2020).
-Research suggests that combining subjective patient perceptions with quantifiable metrics enhances the value and interpretability of patient data (Häikiö et al., 2020).
-For instance, voice assistants powered by large language models (LLMs) can interact with patients using natural language, capture subjective symptom experiences, and visualize them on a dashboard for clinicians (Yang et al., 2024).
-AI can also be leveraged for predictive analytics to help forecast patient health outcomes (Zhang et al., 2024; Pathak, 2024; Collins and Moons, 2019).
+However, given the significant time constraints that clinicians already face (Häikiö et al., 2020), the additional data generated by RPM systems could further exacerbate these challenges.
+Recently, AI-driven tools have shown the potential to assist clinicians in reducing the cognitive load in clinical workflows by analyzing and summarizing large volumes of patient data or by presenting information in a more interpretable format  (Häikiö et al., 2020).
+For instance, voice assistants powered by LLMs can interact with patients using natural language, capture subjective symptom experiences, and visualize them on a dashboard for clinicians (Yang et al., 2024).
+Despite AI having the potential to enhance today’s RPM systems, poor system design of such AI-RPM may introduce new challenges; Thus, we aim to explore the best design practices to ensure these systems can enhance rather than hinder clinicians’ work efficiency.
+2.3. Designing AI-based Systems for Clinical Decision Making
+Recently, HCI researchers have had an increasing interest in the clinical-AI decision making research topic, and their works span various medical domains, such as cancer diagnosis (Cai et al., 2019b, a; Denekamp et al., 2007), cardiotoxicity early detection (Ahmed et al., 2024), diabetic retinopathy screening (Beede et al., 2020), and rehabilitation assessment (Lee et al., 2021, 2020).
+Among some of these works, AI was leveraged for predictive analytics to help forecast patient health outcomes (Zhang et al., 2024; Pathak, 2024; Collins and Moons, 2019).
 Dabek et al.
-(2022) developed an AI model to predict the risk of mental health sequelae following a concussion, achieving 88.2% accuracy in predicting outcomes within 90 days post-injury.
-Despite these advancements, clinicians often remain skeptical about the accuracy of AI-generated predictions (Bohr and Memarzadeh, 2020), which often arises from the lack of transparency in AI decision-making.
-Providing clinicians with a clear understanding of AI’s limitations and actionable insights can enhance their understanding of AI outputs (Liao et al., 2020), thus increasing confidence in AI-assisted decision-making.
-Recent studies emphasize that AI explanations should align with clinicians’ cognitive models (Sendak et al., 2020; Sivaraman et al., 2023) to enhance trust and acceptance of AI-generated recommendations.
-Additionally, supporting AI predictions with evidence-based references and citations may further help clinicians validate AI outputs and increase confidence in AI-assisted decision-making.
-AI has the potential to support remote patient monitoring (RPM) systems by enhancing the efficiency of data collection and processing while supporting clinical decision-making.
-However, poor system design may bring new challenges and workload for clinicians.
-Therefore, integrating well-thought-out design with AI-RPM systems is important to ensure that they enhance rather than hinder clinicians’ work efficiency.
-2.3. Designing AI-Powered Systems for Clinical Decision Support
-In recent years, with the rapid advancement of artificial intelligence (AI) technology, human-computer interaction (HCI) researchers have increasingly explored how AI can better support clinicians in making clinical decisions.
-Their research spans various medical domains, including cancer diagnosis (Cai et al., 2019b, a; Denekamp et al., 2007), cardiotoxicity detection (Ahmed et al., 2024), diabetic retinopathy screening (Beede et al., 2020), and rehabilitation assessment (Lee et al., 2021, 2020).
-Despite these advancements, the integration of AI-driven decision support systems into clinical workflows remains challenging.
-One major issue is that these systems often fail to fully consider the needs of stakeholders during the design phase, limiting their practical usability (Cai et al., 2019b; Green and Chen, 2019; Khairat et al., 2018; Lee et al., 2020; Romero-Brufau et al., 2020) and reducing clinicians’ trust (Yang et al., 2016; Elwyn et al., 2013).
-As a result, AI-driven systems may not effectively assist clinicians in making faster and more accurate decisions (Antoniadi et al., 2021).
-To address this gap, existing research has sought to incorporate human-centered design principles into AI-powered system development.
-Studies have shown that involving stakeholders early in the design process and establishing continuous feedback loops (Abdulaal et al., 2021) can enhance the AI system’s integration into clinical workflows (Wang et al., 2023b), which allows AI-powered systems to be tailored to the needs of different stakeholders, ultimately improving the usability of AI system.
-For example, Yang et al.(Sendak et al., 2020) conducted multiple discussions with clinicians to understand the challenges they face during meetings, collected their feedback, and iteratively refined the design of an AI-powered system based on this input.
+(2022) developed an AI model with 88.2% accuracy for predicting the risk of mental health sequelae within 90 days post-concussion.
+Despite these advancements, the integration of AI-based decision support systems into clinical workflows remains challenging.
+One major issue is that these systems often fail to consider the needs of stakeholders during the design phase, limiting the system’s usability (Cai et al., 2019b; Green and Chen, 2019; Khairat et al., 2018; Lee et al., 2020; Romero-Brufau et al., 2020).
+Moreover, clinicians remain skeptical about the accuracy of AI-generated predictions, which affects their trust in the system (Bohr and Memarzadeh, 2020; Yang et al., 2016; Elwyn et al., 2013).
+As a result, AI-based systems may not effectively assist clinicians in making faster and more accurate decisions (Antoniadi et al., 2021).
+To address this gap, existing research has sought to incorporate human-centered design principles into AI-based system development.
+Studies have shown that involving stakeholders early in the design process and establishing continuous feedback loops (Abdulaal et al., 2021) can enhance the integration of AI systems into clinical workflows (Wang et al., 2023b).
+As a result, AI-based systems can be better aligned with stakeholder needs, ultimately improving their usability.
+For example, Yang et al.(Sendak et al., 2020) conducted multiple discussions with clinicians to understand the challenges they face during meetings, collected their feedback, and iteratively refined the design of an AI-based system based on this input.
 Similarly, Lee et al.
-(2020) engaged seven therapists throughout the system design process, which ultimately increased their willingness to use and satisfaction with the system.
-These studies highlight the importance of human-centered design in AI-driven systems.
-Therefore, it is essential to develop a deep understanding of the challenges and needs of clinicians and actively involve them in the design workflow to fully leverage the potential of AI technologies to help clinicians with concussions detect mental health sequelae and make informed clinical decisions.
+(Lee et al., 2020) involved seven therapists throughout the design process, which led to greater satisfaction and increased adoption of the system.
+These prior works inspired us to begin our project with a formative study to understand concussion clinicians’ challenges and needs.
+Then, we continued to involve them in the design and development process of the AI-RPM system for the mental health sequelae detection scenario.
 3. Formative Study
-To better understand the workflow of clinicians treating youth concussion patients, we conducted a formative study focusing on three aspect: (1) how clinicians gather youth patients’ mental health-related information, (2) the decisions that clinicians need to make, and (3) the challenges they face in their current workflow.
+To better understand the workflow of clinicians treating youth concussion patients with mental health issues, we conducted a formative study focusing on three main aspects: (1) how clinicians gather youth patients’ mental health-related information, (2) what decisions they need to make, and (3) what the difficulties are in their current workflow.
 3.1. Study Participants and Procedure
@@
 P6 Male Sports Medicine Non-Operative Sports Medicine Doctor 10 years
-The key expert stakeholders in the targeted scenarios are concussion clinicians with prior experience treating youth concussion patients with mental health symptoms.
-These highly specialized concussion professionals work in very fast-paced environments and are often overloaded with numerous complex patient cases, which makes their time very limited.
-For this reason, we recruited six U.S.-based concussion clinicians.
+The key stakeholders in our scenarios are concussion clinicians with experience treating youth concussion patients with mental health symptoms.
+These highly specialized concussion professionals work in fast-paced environments and are often overloaded with numerous complex patient cases, which makes their time very limited.
+For this reason, we were only able to recruit six U.S.-based concussion clinicians who were available to participate in our study.
 Recruitment was facilitated through professional networks within related clinical fields of domain experts in the research team via convenience sampling (Sedgwick, 2013).
-Six concussion clinicians were recruited as participants for the formative study, where each participant engaged in a semi-structured interview of 35-45 minutes conducted remotely via Zoom.
+Each participant engaged in a semi-structured interview of 35-45 minutes conducted remotely via Zoom.
 Table 1 provides details on the demographics of the participants and their clinical experience.
-This study was reviewed and approved by first author’s Institution Institutional Review Board (IRB) and the study complied with the approved procedure for ethical research practices of human subject protection.
-At the beginning of each interview, we first asked the participant to recall a recent case where they encountered a youth concussion patient with mental health sequelae.
-They described how they performed clinical practices to detect mental health sequelae and executed follow-up decision making.
-While participants were sharing their experiences, we prompted participants to tell more about what type(s) of patient clinical data they collected, used or expected to support the identification of mental health sequelae, both in and outside of clinical settings.
-After that, participants shared their experience with existing tools, if any, for collecting patient data, monitoring patient status, and detecting mental health sequelae.
-Participants shared their feedback and insights on the benefits and limitations of existing tools and the expectations of novel tools that could potentially help them.
+This study was reviewed and approved by the first author’s institution’s Institutional Review Board (IRB), and the study complied with the approved procedure for ethical research practices of human subject protection.
+At the beginning of each interview, we first asked the participant to recall a recent case where they encountered a youth concussion patient with mental health symptoms They described how they performed clinical practices, identified mental health symptoms, and executed follow-up decision-making.
+While participants were sharing their experiences, we prompted participants to tell more about what type(s) of patient clinical data they collected, used, or expected to support the identification of mental health sequelae, both in and outside of clinical settings.
+After that, participants shared their experience with existing tools, if any, for collecting patient data, monitoring patient mental health conditions, and identifying mental health sequelae.
+Moreover, participants shared their insights on the benefits and limitations of existing tools and the expectations of novel tools that could potentially help them.
 During the interview, we asked participants to refrain from disclosing any personally identifiable information (PII).
@@
 Two researchers first independently coded one session of the interview transcripts to identify meaningful segments and categorize segments into codes.
-The researchers then discussed individual codes and resolved any discrepancies, developed consensus codes, and applied the resulting codes to the remaining data.
+Then, they discussed individual codes, resolved discrepancies, developed consensus codes, and applied the resulting codes to the remaining transcripts.
+Finally, we iteratively grouped codes into overarching themes and refined these themes through ongoing discussions until we reached a consensus on the final set of themes.
 3.2. Findings
-Our qualitative analysis provided a comprehensive understanding of concussion clinicians’ workflow with youth concussion patients, and derived three major challenges that concussion clinicians encountered during the practices: (1) collecting patients’ mental health information outside the clinic; (2) communicating effectively with patients and their family members; (3) Assessing patients’ compliance with clinical recommendations.
-3.2.1. Clinical Workflow of Concussion Clinicians
-The majority of youth concussion patients are athletes where, and concussions are commonly caused by head injuries during sports participation.
+Our qualitative analysis provided a comprehensive understanding of concussion clinicians’ workflow with youth concussion patients and derived three major challenges that concussion clinicians encountered during the practices: (1) tracking patients’ mental health-related information outside the clinic and identifying its severity; (2) communicating effectively with patients and their family members about patients’ mental health issues; (3) assessing patients’ compliance with clinical recommendations.
+3.2.1. Integrating Mental Health Screening into Concussion Management
+The majority of youth concussion patients are athletes, and concussions are commonly caused by head injuries during sports participation.
 After they had a head injury, the youth patients underwent an initial assessment by primary care physicians.
@@
 When patients visit the clinic, concussion clinicians assess the severity of concussion symptoms with standard scales as suggested by clinical guidelines.
-In addition, concussion clinicians will conduct interview sessions with patients to collect more in-depth information about their health conditions.
-”And I guess the only way that I currently assess my patients in that way is by using the symptom score that is published in the international guidelines, at least as of 2017.
-That’s the symptom score that I use.
+In addition, the clinicians conduct interview sessions with patients to collect more in-depth information about their health conditions.
+”I guess the only way that I currently assess my patients in that way is by using the symptom score that is published in the international guidelines…… That’s the symptom score that I use.
 And then interview.
@@
 I will very commonly ask about nutrition habits.
-I am very aggressive in making I guess my patients go to school.” (P3)
-For the majority of concussion patients, the best treatment available is just to keep rest both physically and cognitively.
-Thus, concussion clinicians will often engage in a conversation with patients after finishing the assessment to provide personalized lifestyle recommendations and self-evaluation methodologies that would be best suited for concussion recovery.
-For example, one participant (P5) stated that they often provide patients with symptom logs and suggested patients track their daily experiences.
+I am very aggressive in making, I guess my patients go to school.” (P3)
+For the majority of concussion patients, the best treatment available is to have a rest both physically and cognitively.
+Thus, concussion clinicians will have a conversation with patients to provide personalized lifestyle recommendations and self-evaluation methodologies for concussion recovery.
+For example, P5 stated that they often provide patients with a symptom log and suggested that patients use it to track their daily activities.
 Patients will be asked to return the symptom logs during follow-up appointments, which typically occur after one week.
-Symptom logs are crucial for concussion clinicians because they will heavily rely on the symptom logs to get a thorough understanding of how patients underwent the past week and how the progression of recovery from concussion is.
+Symptom logs are essential for concussion clinicians to monitor their patients’ weekly health changes and concussion recovery progress.
 If symptoms persist beyond 28 days, further interventions are initiated, such as referrals to neuropsychologists or specialized concussion clinics (P3, P5, P6).
 ”Yeah, after we see them (patients) for the initial visit, we generally send them home with the log and ask them to fill it out every day based on how they felt, and then bring it back when they come back for their follow up appointment so that we can track how their symptoms have been going over time.” (P5)
-Concussion clinicians (P1, P2, P3, P5, P6) emphasized that mental health concerns are highly prevalent among youth concussion patients.
+Concussion clinicians (P1, P2, P3, P5, P6) emphasized that mental health concerns are highly prevalent among youth patients.
 Thus, mental health assessments are an integral part of the concussion assessment process.
@@
 Concussion clinicians (P4, P6) also observe signals of emotional distress during the conversation with patients to detect potential mental health issues and will follow up with additional mental health screening if such signals are presented.
-Clinician (P6) reported that there is a high portion of concussion patients develop anxiety or depression following their injury.
-In such cases, clinicians (P4) often recommend cognitive behavioral therapy (CBT) interventions as the intervention strategy for the mental health issues mentioned above.
+P6 reported that there is a high portion of concussion patients develop anxiety or depression following their injury.
+In such cases, P4 often recommends cognitive behavioral therapy as the intervention strategy for the mental health issues mentioned above.
 However, if mental health symptoms are found to impede concussion recovery, concussion clinicians will need to collaborate with neurologists to determine the intervention plan, such as prescribing appropriate medications.
@@
 If no immediate suicide risk is detected, clinicians typically recommend continued rest at home with ongoing symptom monitoring.
-However, the concussion clinician (P2) mentioned that 10% of youth concussion patients develop suicidal thoughts after a concussion.
+However, P2 mentioned that 10% of youth concussion patients develop suicidal thoughts after a concussion.
 In such cases, clinicians refer patients to mental health specialists, such as psychiatrists or psychologists.
-”In general, of people who I see with like symptoms of like anxiety, depression, things like that after concussion. It’s probably about 40 to 50 percent.” (P6)
-3.2.2. Clinicians’ Challenges in Tracking Patients’ Mental Health Information Outside of Clinics
+“In general, of people who I see with like symptoms of like anxiety, depression, things like that after concussion. It’s probably about 40 to 50 percent.” (P6)
+3.2.2. Clinicians’ Struggle to Track and Assess Mental Health Beyond the Clinic
 As mentioned above, the first 28 days following a concussion diagnosis is a critical period of time for youth concussion patients’ recovery.
 However, patients only visit the concussion clinic every one or two weeks, while spending the majority of time outside the clinic, more specifically, outside of clinical supervision.
-This gap creates significant challenges for concussion clinicians, who (P2, P5) emphasize the importance of monitoring a wide range of mental health symptom indicators during this recovery period.
-For instance, clinicians (P1, P2, P3, P4, P5) would focus on collecting detailed information about sleep patterns (e.g., sleep hours and wake-up times), and physical activity levels.
-P3 noted that the sleep data has not really been integrated into concussion statements.
-Moreover, clinicians (P2, P3, P6) are closely interested in youth patients’ academic performance, such as whether they completed assignments on time, and social behavioral patterns, such as whether they rest at home consistently and maintain social connections rather than withdrawing socially.
-These physical and cognitive behaviors could provide critical insights for clinicians to assess and reflect on youth patients’ mental health status during concussion recovery.
-”…if the worry is getting to the point where they (patients) can’t do their homework… Or if they aren’t leaving the house… not sleeping, really feeling terrible, and it seems to be consuming them… Then I would say that might like cross the line from just typical worry into like, okay, this is more of a mental health anxiety, you know, diagnosis type of thing.” (P2)
-Furthermore, P4 stated that during clinical visits, the clinician mainly relies on nonverbal cues in the clinical environment to identify youth patients’ mental health conditions.
-If any abnormalities are detected, youth patients would be asked directly about mental health-related symptoms, such as suicidal tendencies.
-Mental health-related information is crucial, as mental health issues impact the clinician’s ability to adjust treatment plans in a timely manner.
-Moreover, in some cases, youth concussion patients may experience impairment in functioning or panic attacks, and the concussion clinicians will then focus on understanding what they had been through and what events triggered the episode to more accurately evaluate patients’ condition as well as design a more appropriate concussion recovery plan to avoid these negative events happening again, as suggested by P2.
-”……I think, having more information about really what was happening at the time, that they had a symptom. What was triggering, that how that we know more real time information could be helpful.” (P2)
-Despite the critical need for timely health data of patients at home, concussion clinicians (P2, P5) reported significant challenges in collecting accurate and reliable information about patients’ mental health outside the clinic.
-Ideally, clinicians would benefit from having real-time access to patients’ health data while they are at home.
+This gap creates significant challenges for concussion clinicians, who (P2, P5) emphasize the importance of monitoring a wide range of mental health symptom indicators during the concussion recovery period.
+Specifically, clinicians (P1, P2, P3, P4, P5) mentoined the importance of collecting detailed information about sleep patterns (e.g., sleep hours and wake-up times), and physical activity levels.
+P3 emphasized that the sleep patterns has not really been integrated into clinical systems.
+Moreover, clinicians (P2, P3, P6) are highly interested in youth patients’ academic performance (whether patients completed assignments on time) and social behavioral patterns (whether patients rest at home consistently and maintain social connections rather than withdrawing socially) during concussion recovery time at home.
+These academic and social behaviors provide critical insights for clinicians to assess and reflect on the severity of youth patients’ mental health issues during concussion recovery.
+“From their (patients) activities, feeling [that they have] little interest or in doing activities that they normally would.
+Seeing difficulty or decreased performance in either activities or or school.” (P6)
+Furthermore, P4 stated that to capture mental health-related anomalies, the clinician mainly relies on nonverbal cues during the clinical visits.
+If any abnormalities are detected, youth patients would be asked questions directly about their mental health conditions, such as suicidal tendencies, to identify the severity of their mental health symptoms.
+P5 stated that identifying the severity of the symptoms is crucial for clinicians, as it impacts the clinician’s ability to adjust their treatment plans in a timely manner.
+Moreover, youth concussion patients may experience impairment in functioning or panic attacks, and concussion clinicians will then focus on understanding what they had been through and what events triggered the episode.
+In this way clinicians can evaluate patients’ mental health condition more accurately as well as design a more personalized concussion treatment plan, as suggested by P2.
+“Anybody who has a high report of anxiety, you know, as one of the symptoms or nervousness that we ask on our concussion symptoms log or if they have an underlying diagnosis of an anxiety, disorder, depression.
+And then, you know, maybe a little bit more likely to refer them (patients) earlier on.” (P5)
+Despite the critical need for timely mental health-related information of patients at home, concussion clinicians (P2, P5) reported significant challenges in accurately and reliably collecting the information outside the clinic.
+Ideally, clinicians would benefit from having real-time access to patients’ accurate mental health-related information while patients are at home.
 However, such a system is currently unavailable.
@@
 Once youth concussion patients leave the clinic, clinicians have no way to obtain this information in a timely manner.
-”…more real-time information could be helpful.
+“…more real-time information could be helpful.
 Because by the time we see them (patients) in the clinic that could be you know, weeks later they might not even remember that that (mental health sequelae-related symptoms) had happened…… It’s hard for anyone to remember what they did, you know, weeks before.” (P2)
-Youth concussion patients currently have some remote communication options, such as making phone calls with the provider team, to report their health status regularly to clinicians when they are outside of clinic.
+Youth concussion patients currently have some remote communication options.
+One way is making phone calls with the provider team and reporting patients’ mental health-related information regularly to clinicians when they are outside of the clinic.
 However, clinicians (P2, P3) reported several limitations with phone communication.
@@
 Moreover, even if youth concussion patients make a phone call, clinicians may not always be in the office or available to answer immediately, which can delay the transmission of critical health information.
-An alternative option is using MyChart, a patient portal that allows youth concussion patients to access their medical records and send messages to their healthcare providers.
-Concussion clinicians (P4) strongly recommend patients use MyChart to share health-related information, particularly when youth concussion patients are reluctant to make direct phone calls.
+An alternative option is using MyChart, a patient portal that allows patients to access their medical records and send messages to their healthcare providers.
+Concussion clinicians (P4) strongly recommend that patients use MyChart to share health-related information, particularly when youth concussion patients are reluctant to make direct phone calls.
 Compared to phone communication, MyChart offers the advantage of asynchronous messaging, enabling clinicians to review patient information and respond at their convenience.
@@
 Concussion clinicians (P2, P6) noted that youth patients rarely use the messaging feature in MyChart and rarely take the initiative to contact clinicians.
-Instead, parents of these youth patients are more likely to use MyChart to communicate with clinicians, which impedes the information flow between patients and clinicians.
-”One in 10 [patients] maybe, or even one in 20 [patients] might send a message with a question.” (P6)
+Instead, parents of these youth patients are more likely to use MyChart to communicate with clinicians.
+“One in 10 [patients] maybe, or even one in 20 [patients] might send a message with a question.” (P6)
 Furthermore, clinicians caution against relying on MyChart for emergencies, especially when patients experience suicidal tendencies.
-Clinicians (P2) warned that their busy schedules prevent them from checking messages in real time, which could lead to delays in critical interventions and put patients at risk.
-”Yeah, it’s not really immediate.
+P2 warned that their busy schedules prevent them from checking messages in real time, which could lead to delays in critical interventions and put patients at risk.
+“It’s not really immediate.
 You know, cause they (patients) might send it at night, and then we don’t maybe don’t see it till the next afternoon, and you know.
 We usually, you know, tell them if it’s an emergency, don’t use that method because we don’t really know when we’re gonna see them or be able to respond to it.” (P2)
-3.2.3. Clinicians’ Challenges in Communicating with Patients and Families About Mental Health Issues
+3.2.3. Clinicians’ Struggle to Communicate Mental Health Issues with Patients and Their Families
 Communication challenges persist not only outside the clinics but also during in-person consultations.
 During in-person visits, parents often accompany their children.
-However, some patients may feel uncomfortable to disclose sensitive mental health issues.
-One clinician (P1) described a case where a patient with severe anxiety disclosed self-harm and suicidal thoughts only after requesting to speak privately, away from their parents’ presence.
-“I stepped out of the I had the mom stay in the room……And I said, ’Is there something wrong? Is there something going on?’ He said, ’Yeah, I need to talk to you privately.’” (P1)
-Clinicians suspect that a major reason that patients choose not to disclose their mental health status is their fear of being labeled as having a mental illness.
-Being labeled as having a mental illness can affect how they are perceived socially, which could potentially isolate them from their peers.
-Some patients may also struggle with the societal expectation of their physical recovery, especially if youth patients are athletes who are expected to perform at a high level of competition.
-This highlights the sensitivity of mental health disclosures is intertwined with many societal factors that may not be directly related to mental health symptoms.
-Moreover, some concussion clinicians (P1, P4, P5) reported that sometimes both youth concussion patients and their parents may deliberately withhold information about the patient’s mental health.
-In one case (P4), a youth concussion patient was a successful athlete, and the parents downplayed their child’s mental health symptoms during clinical visits.
-The parents may expect to expedite the child’s recovery process so that they can return to their athletic career sooner.
+However, some youth concussion patients feel uncomfortable disclosing sensitive mental health issues with their parents present.
+P1 described a case where a patient with severe anxiety disclosed self-harm and suicidal thoughts only after requesting to speak privately, away from their parents’ presence.
+“I stepped out of the [room], I had the mom stay in the room…… and I said, ’Is there something wrong? Is there something going on?’ He said, ’Yeah, I need to talk to you privately.’ ” (P1)
+Clinicians suspect that a major reason for this is patients’ fear of mental health stigma (Association, 2023).
+The label of having a mental illness can affect how patients are perceived socially, which could potentially isolate them from their peers.
+Another reason not to disclose mental health issues is the societal expectation toward youth concussion patients.
+Specifically, if youth patients are athletes who are expected to perform at a high level of competition, they may struggle with the societal expectation of their physical recovery.
+Not only do patients withhold their mental health condition, but sometimes their parents also deliberately withhold information about the patient’s mental health condition as reported by concussion clinicians (P1, P4, P5).
+In one case encountered by P4, the youth concussion patient was a successful athlete, and the parents downplayed their child’s mental health symptoms during clinical visits.
 These parents may worry that prolonged rest due to mental health sequelae could hinder their child’s athletic performance and career trajectory.
-However, when mental health issues are not disclosed and addressed in a timely manner, the issues could get worse and lead to serious consequences, such as self-harm or an increased risk of suicidal behavior.
-The consequences of severe mental health issues could lead to long-term physical and cognitive impairment of the patients as well as hinder an athlete’s ability to return to sports.
-”Their (patients) parents are very vested in the child’s athletic career……the parents come in and there’s obviously an agenda to minimize their symptoms.
+So the parents downplayed their child’s mental health symptoms, hoping for their child to return to their athletic career sooner.
+However, when mental health issues are not disclosed and addressed in a timely manner, the issues could get worse and lead to severe mental health issues, such as self-harm or an increased risk of suicidal behavior.
+The consequences of severe mental health issues could lead to long-term physical and cognitive impairment of the patients, which might eventually hinder an athlete’s ability to return to sports in their lives.
+“Their (patients) parents are very vested in the child’s athletic career……the parents come in and there’s obviously an agenda to minimize their symptoms.
 Or you know, just try not to buck the system, if it’s gonna limit their (patients) activity.” (P4)
@@
 Another is to observe the patient’s behavior during consultations.
-For example, clinicians (P1, P4) can observe patients’ behaviors, such as signs of anxiety or a persistently quiet and disengaged attitude, during clinical visits.
+For example, clinicians (P1, P4) observe patients’ behaviors, such as signs of anxiety or a persistently quiet and disengaged attitude, during clinical visits.
 Moreover, worsened concussion symptoms such as insomnia and low activity levels may indicate underlying issues beyond the concussion itself, such as mental health sequelae.
-One clinician (P3) stated that if a youth patient’s concussion symptoms do not improve within four weeks after a concussion, it may indicate underlying mental health issues that are contributing to a prolonged recovery period.
-However, both observing patients during in-person visits and asking about their sleep and activity levels at home require clinicians to gather this information during the clinical visit.
-Once the patient returns home, clinicians can hardly access this information.
-”Because we expect 96 percent of concussions to be healed by the four-week point.
+P3 stated that if a youth patient’s concussion symptoms do not improve within four weeks after a concussion, it may indicate underlying mental health issues that are contributing to a prolonged recovery period.
+However, observing patients’ behavior or asking about their sleep patterns requires patients to be physically present during the clinical visit.
+Once the patients return home, clinicians cannot access this information.
+“Because we expect 96% of concussions to be healed by the four-week point.
 And so if it’s not healed yet, then it’s probably not the concussion, right?
 It’s probably either a headache syndrome or a mental health concern.” (P3)
-3.2.4. Clinicians’ Challenges in Assessing Patients’ Compliance with Clinical Recommendations
-At the end of each clinical visit, concussion clinicians typically provide patients with recommendations related to physical activity or sleep.
-These clinical recommendations are intended to serve as the concussion treatment plan to support both physical and mental recovery.
+3.2.4. Clinicians’ Struggle to Evaluate Patients’ Compliance Beyond the Clinic
+At the end of each clinical visit, concussion clinicians typically provide patients with recommendations related to physical activity levels or sleep schedules.
+These clinical recommendations are intended to serve as the concussion treatment plan to support both physical and mental recovery after a concussion.
 However, understanding whether patients adhere to these recommendations is a persistent challenge.
-Concussion clinicians (P2, P4) rely on patients’ self-reports to assess compliance during the time period between the last visit and the current one, despite these reports are often unreliable.
+Although concussion clinicians (P2, P4) can rely on patients’ self-reports to assess compliance during the time period between the last visit and the current one, these reports are often unreliable.
 For instance, some patients may claim to have followed the recommendations, but their responses to mental health questionnaires reveal high levels of anxiety or depression, indicating limited or no improvement.
-“What we recommended, if they(youth patients) did it or not, they might not have done it… … We don’t know if they’re following them or not.
+“What we recommended, if they (youth patients) [said they] did it or not, they might not have done it… … We don’t know if they’re following them or not.
 Or they come in, and they say they did, but we don’t know if they did.” (P2)
-Concussion clinicians (P) noted that there is currently no effective way to track patients’ compliance with clinical recommendations outside the clinic.
-Clinicians (P4) mentioned that they often ask patients’ parents about youth patients’ adherence to recommendations but found that information from family members is also hardly reliable.
-One reason could because some parents are frequently away from home due to work commitments, which leaves them with limited knowledge of patients’ daily health-related information.
-In such cases, clinicians may reach out to other family members for a closer understanding of the patient’s mental health conditions when they are at home.
-However, making such phone calls consumes a considerable amount of time in clinicians’ busy schedules, and the information collected through these calls is often limited.
-“So I think sometimes the parent comes in with a child who isn’t around them that much at home.
-Like maybe it’s the working parent, and they’re just sort of the driver and they don’t have a lot of information…… So sometimes I can’t get good information ’cause they’re not engaged as much as I would, you know, would hope.” (P4)
-4. Initial System Design
-In our formative study, We found that clinicians have a need for remotely monitor the health-related information of youth concussion patients at home, particularly regarding their mental health.
-Additionally, we found that clinicians face challenges when making mental health-related decisions, particularly during clinical visits, where it can be difficult to gather mental health information from patients and their families through communication.
-However, clinicians rely on this information to guide their clinical decisions and treatment plans.
-Building on these identified challenges and needs, we proposed AI-RPM technology that could potentially assist clinicians in remotely monitoring patients’ at-home data and addressing decision-making difficulties.
-We then visualized a possible system interface based on these technologies and detailed the specific functions of the system.
-The goal of our system design is to explore whether concussion clinicians find the technologies useful in remotely collecting youth patient mental health-related information outside of clinics and supporting clinicians’ challenges associated with mental health-related decision-making.
-4.1. Applicable AI-RPM Technology
-4.1.1. Wearables in Tracking Patients’ Sleep Pattern and Physical Activities
+However, there is currently no effective way to track patients’ compliance outside the clinic.
+P4 mentioned that they often ask patients’ parents about their children’s adherence to the recommendations, but the information provided by the parents may also be unreliable.
+One reason could be that some parents are frequently away from home due to work commitments, which leaves them with limited knowledge of patients’ daily health-related information.
+In such cases, clinicians may reach out to other family members by phone for a closer understanding of the patient’s mental health issues when they are at home.
+However, making such calls consumes a considerable amount of time in clinicians’ busy schedules, and the information collected through these calls is often limited.
+“I think sometimes the parent comes in with a child who isn’t around them (parents) that much at home.
+Like maybe it’s the working parent, and they’re just sort of the driver, and they don’t have a lot of information…… Sometimes I can’t get good information cause they (parents) are not engaged as much as I would, you know, would hope.” (P4)
+4. Preliminary System Design
+In the formative study, we identified clinicians’ urgent need to remotely monitor the mental health of youth concussion patients at home, as well as the challenges they face in clinical decision-making.
+In this section, we explored the design opportunities of AI-RPM technologies in supporting clinicians’ needs and decision-making during their practice.
+Subsequently, we proposed a preliminary system design (Fig.
+2) based on the AI-RPM technologies with detailed features.
+The goal of the design is to explore whether concussion clinicians find the AI-RPM system useful for supporting remote monitoring and decision-making, how the system can be improved, and how they would use it in practice.
+4.1. Design Considerations for AI-RPM Systems
+4.1.1. Wearables for Tracking Patient Physiological Data
 Based on Section 3.2.2, clinicians expressed a desire to monitor patients’ at-home quantitative data, such as sleep and exercise patterns, as these factors are closely related to concussion recovery and mental health status.
-However, they currently lack effective ways to access real-time health data from patients at home.
-In addition, as noted in Section 3.2.4, clinicians frequently ask patients whether they adhere to clinical recommendations, such as maintaining proper sleep and exercise routines.
-However, clinicians often struggle to verify the accuracy of patients’ self-reported information.
-To address these challenges, we propose that wearable devices such as smartwatches could help clinicians collect real-time sleep and exercise data from patients and provide these data to clinicians for timely review.
+In addition, as noted in Section 3.2.4, clinicians frequently ask patients about their clinical recommendation compliance, such as maintaining proper sleep and exercise routines.
+However, clinicians currently lack effective ways to access these at-home data and often struggle to verify the accuracy of patients’ self-reported information during clinical visits.
+To address these challenges, wearable devices such as smartwatches can be used to help clinicians collect real-time sleep and exercise data from patients (Dias and Paulo Silva Cunha, 2018; Baig et al., 2017; Bate et al., 2023).
 Wearables offer a non-invasive and natural way to monitor patients’ health metrics without significantly disrupting their daily routines.
-As discussed in Section 2.2, wearable devices can continuously monitor patients’ sleep and physical activity data in real-time (Dias and Paulo Silva Cunha, 2018; Baig et al., 2017; Bate et al., 2023).
-We believe that patients’ sleep patterns and physical activity data collected through wearables could enable clinicians to remotely access at-home patient data, track adherence to sleep and exercise recommendations, and ultimately make more informed mental health-related clinical decisions.
-4.1.2. LLM-Powered Conversational Agents in Supporting Patients Self-Report Mental Health Symptoms
-In Section 3.2.2, clinicians expressed a need to understand patients’ experiences with symptoms such as panic attacks, including what specifically triggered them.
+Moreover, AI has the potential to process the massive amount of data collected from wearables and provide these data to clinicians for timely review.
+We believe that with access to at-home patient data, clinicians can become aware of patients’ health conditions in real-time, and track patients’ compliance with sleep and exercise recommendations, thus ultimately having the ability to make more informed mental health-related clinical decisions, such as setting up a follow-up meeting for further assessment.
+4.1.2. LLM-based CAs for Patient Self-Reporting in Mental Health
+In Section 3.2.2, clinicians emphasized the need for a comprehensive understanding of patients’ experiences with mental health symptoms, such as panic attacks, to provide effective care.
 However, patients often struggle to accurately recall these details, and such subjective experiences cannot be easily quantified using standardized scales or medical instruments.
 Additionally, based on Section 3.2.3, concussion clinicians reported that youth concussion patients sometimes conceal their mental health issues from their parents.
-It was only during a private conversation with the clinician that the patient disclosed the truth about their mental health issues.
-Given this challenge, we hypothesize that LLM-powered conversational agents (CAs) (Yang et al., 2024; Mahmood et al., 2023; Chan et al., 2024) could help clinicians collect mental health-related information effectively.
+It was only during a private conversation with the clinician that the patient disclosed their mental health status.
+Given this challenge, we hypothesize that LLM-based CAs (Yang et al., 2024; Mahmood et al., 2023; Chan et al., 2024) could help clinicians collect mental health-related information effectively.
 CAs can be deployed in patients’ homes to engage in natural language conversations, ask relevant questions, and record dialogue content.
-AI can then analyze historical conversations to summarize the content of the conversation, which allows clinicians to quickly gain insights into their mental well-being.
+AI can then analyze historical conversations to summarize the key information of the conversation, such as their patient’s suicidal intentions.
+From the perspective of youth patients, the deployment of a CA in their rooms could increase the chance of disclosing their mental health issues, if there are any.
 One study found that children perceive conversational agents as warmer and more reliable than their parents (Van Brummelen et al., 2023).
-Therefore, we speculate that CAs could encourage patients to disclose their mental health issues, enable AI to analyze these conversations, extract key information, and securely transmit relevant insights to clinicians while preserving privacy.
-As a result, clinicians could review patients’ historical conversations with CAs, gaining insights into their subjective experiences, emotional states, and mental health-related life events—ultimately supporting a more accurate assessment of patients’ mental health.
-4.1.3. AI Risk Prediction in Supporting Mental Health Sequelae Detection
-As discussed in Section 3.2.3, clinicians reported that patients and their parents sometimes deliberately conceal mental health conditions, making it difficult for clinicians to develop appropriate treatment plans.
-In addition, clinicians noted that the severity of mental health issues in a patient influences treatment plans and subsequent clinical decisions by clinicians, such as whether a referral is necessary.
-Given these two challenges, we believe that AI-powered risk prediction for disease progression could provide valuable support to clinicians.
-AI has the capability to analyze patients’ health data and use algorithms to predict the likelihood of developing a condition over a certain period (Collins and Moons, 2019).
-Recent studies have already explored AI-driven risk prediction for mental health sequelae following a concussion and have found that AI predictions can achieve high accuracy (Dabek et al., 2022).
-By leveraging AI risk score prediction, clinicians may no longer have to rely solely on patients’ self-reported mental health status and daily behaviors when making assessments.
-Moreover, if patients or their parents choose to withhold information, AI could help clinicians detect potential mental health sequelae in a timely manner.
-If a patient is flagged as having a high risk of developing mental health sequelae, the clinician can proactively schedule follow-up visits and make necessary mental health-related clinical decisions.
-Therefore, we hypothesize that AI-driven risk prediction has the potential to assist concussion clinicians in detecting mental health sequelae more effectively.
+From the perspective of clinicians, gaining such key information allows them to review patients’ historical conversations with CAs, gain accurate information about patients’ mental health-related experiences while preserving privacy, and conduct interventions if necessary.
+Therefore, we speculate that CAs have the potential to act as a communication bridge between clinicians in the hospital and patients at home and provide patients with self-reported symptoms.
+Ultimately, CAs could support clinicians in conducting a more accurate assessment of patients’ mental health and making timely decisions.
+4.1.3. AI-based Risk Prediction for Detecting Mental Health Sequelae
+As we have indicated in Section 4.1.2, clinicians reported that patients and their parents sometimes deliberately conceal mental health conditions.
+In addition, clinicians noted that the severity of mental health symptoms in a patient influences their subsequent clinical decisions, such as treatment plans or providing a referral.
+However, evaluating the severity of mental health symptoms remains challenging.
+Given these two challenges, we believe that AI risk prediction for disease progression could provide valuable support to clinicians in our setting.
+AI has the capability to analyze patients’ health data from the EHR system and utilize algorithms to predict the likelihood of developing specific conditions over a given period (Collins and Moons, 2019).
+In particular, recent studies have already explored AI-driven risk prediction for mental health sequelae following a concussion and have found that AI predictions can attain high accuracy (Dabek et al., 2022).
+By leveraging AI risk score prediction, clinicians may no longer have to rely solely on the conversations with patients during clinical visits to make mental health assessments.
+If patients or their parents choose to withhold information, AI could help clinicians detect potential mental health sequelae in a timely manner.
+Moreover, the severity of the mental health symptoms is normally associated with the AI prediction score.
+The AI prediction score can help clinicians better assess the severity of a patient’s mental health symptoms and make informed decisions accordingly.
 Figure 2.
-The initial system design draft for evaluation and gathering information from concussion clinicians: (A) AI risk score prediction for mental health sequelae, (B) patient self-report symptoms, (C) patient sleep and physical activities.
-4.2. Preliminary UI Design of the System
-Based on the potentially applicable technologies that could support concussion clinicians in addressing these challenges, we designed three modules for our initial system design, as shown in Figure 1: (1) AI risk score prediction for mental health sequelae (Fig.
-1.A), (2) patient self-reported symptoms (Fig.
-1.B), and (3) patient sleep and physical activity tracking (Fig.
-1.C).
-During our UI design process, we adhered to Human-Centered AI (HCAI) principles to ensure a user-friendly interface.
-Our goal with this design is to visualize the system architecture and interface based on the previously proposed technologies, which makes our system design more intuitive for clinicians.
-Then, we can use the initial system design to gather clinicians’ perceptions and feedback on visual representation, data types, and technologies.
+The preliminary system design based on concussion clinicians’ challenges and needs: (A) AI risk score prediction for mental health sequelae, (B) patient self-report symptoms, (C) patient sleep and physical activities.
+4.2. Three Modules of Preliminary System Design Based on the Design Considerations
+Based on the potentially applicable technologies that could support concussion clinicians in addressing their challenges and supporting their decision-making, we designed three modules for our preliminary system, as shown in Fig.
+2.
+We described each module in detail in this section.
+4.2.1. The Sleep and Physical Activities Module
+The sleep and physical activities module (Fig.
+2 B) presents sleep (sleep hours, light sleep, and wake-up time) and activity data (step count) that clinicians want to track.
+These data can be collected by wearables as we discussed on Section 4.1.1 Moreover, clinicians can select any category to review the patient’s relevant data from the past week.
+Clinicians can also select different dates to review patient data from other time periods.
+Additionally, the module provides the source of the data for increasing clinicians’ understanding of the module and their trust towards the system.
+4.2.2. The Self-Report Symptoms Module
+To display the patients’ self-reported symptoms at home, which were collected by CAs as we proposed in Section 4.1.2, we designed the self-reported symptoms (Fig.
+2 B).
+On the right side, a history of conversations between youth concussion patients and CAs is available for concussion clinicians to review.
+Based on the conversation history, AI detects the symptoms that are mentioned by patients and highlights key information that are relevant to mental health sequelae, such as anxiety, depression, and suicidal ideation.
+If a patient self-reports symptoms within a specific category, the corresponding indicator changes from green to red.
+We use the color red to catch clinicians’ attention.
+To facilitate efficient navigation, clinicians can click on a category name or use the search bar to enter keywords, which prompts the right-side module to display the relevant conversation excerpts.
+Similar as the sleep and physical activities module, time selection lets clinicians view data from different periods.
 4.2.3. The AI Risk Score Prediction Module
-The first module is AI Risk Score Prediction(Fig 1.A).
-The module shows the probability of mental health sequelae within one year with a range from 0% to 100%, where the higher probability score signifies a higher probability of experiencing mental health sequelae for clinicians to review.
-Feature importance scores are also provided to provide additional explanations for clinicians to decide whether to trust or not to trust the AI-predicted score.
-This module also offers text-based explanations to help clinicians understand the sources of the factors that contribute to the risk score.
+The AI Risk Score Prediction (Fig.
+2 A) module leverages AI’s ability that we discussed in Section 4.1.3 to show the probability of mental health sequelae.
+The prediction timeframe is within one year with a range from 0% to 100%, where the higher probability score signifies a higher probability of experiencing mental health sequelae for clinicians to review.
+Moreover, feature importance is used to provide data that influence the AI prediction score, which allows clinicians to decide whether to trust the AI-predicted score or not.
+To increase AI explanations and transparency, we provide text-based explanations to help clinicians understand the risk score and its source.
 At the bottom is the risk score chart.
-Clinicians can hover over any day on the risk score chart to view the likelihood that young patients develop mental health sequelae at that specific point in time.
+Clinicians can hover over any day on the chart to view the likelihood of developing mental health sequelae at that specific point in time.
 Additionally, clinicians can select different dates to review patient data from various time periods.
-This module aims to enable clinicians to remotely monitor youth patients’ mental health sequelae without requiring in-person consultations, which supports concussion clinicians in making timely, informed decisions based on patients’ mental health conditions.
-4.2.2. The Self-Report Symptoms Module
-The second module is the self-reported symptoms (Figure 1.B).
-On the right side of the module, a history of conversations between youth concussion patients and conversational agents (CAs) is available for concussion clinicians to review.
-Based on the conversation history, AI detects the symptoms that are mentioned by patients and highlights categories that are relevant to mental health sequelae, such as anxiety, depression, and suicidal ideation.
-The categories are displayed on the left side of the panel.
-If a patient self-reports symptoms within a specific category, the corresponding indicator changes from green to red, which reminds clinicians to review the entry.
-To facilitate efficient navigation, clinicians can click on a category name or use the search bar to enter keywords, which prompts the right-side module to display the relevant conversation excerpts.
-Additionally, clinicians can select different dates to review patient data from various time periods.
-The primary goal of this module is to allow concussion clinicians to quickly assess patients’ mental health symptoms at home, extending support beyond clinical settings.
-4.2.3. The Sleep and Physical Activities Module
-Lastly, we designed the sleep and physical activities module (Fig.
-1.B).
-The module presents sleep-related information that concussion clinicians want to track for their patients at home, including sleep hours, light sleep, and wake-up time, as well as activity-related data such as step count.
-When a clinician selects any category, the module displays the patient’s relevant data from the past week.
-Clinicians can also select different dates to review patient data from other time periods.
-Additionally, the module provides the source of the data, which is through wearable sensors, such as smartwatches.
-The goal of this module is to assist concussion clinicians in monitoring youth concussion patients’ sleep and physical activity at home, addressing the current gap in existing clinical systems that lack this information.
-5. Design Interface Evaluation
-In the previous section, we introduced the potential of AI-RPM technologies to assist clinicians in remotely collecting patient information and supporting mental health-related decision-making.
-We then designed a system that integrates these technologies, focusing on how the collected data is presented and utilized.
-Based on the initial system design, we conducted a design interface evaluation (Savage, 1996) to gather feedback from the six original clinicians who represent our target end-users.
-The goal of this section is to explore concussion clinicians’ perceptions of the system—specifically, whether they find these technologies useful for remote patient monitoring and how the system might better support them in addressing challenges related to mental health decision-making.
+5. Evaluation Study of the Preliminary System Design
+To evaluate the usefulness of our preliminary design, we conducted an evaluation study.
+The goal of the evaluation study is to gather clinicians’ feedback and insights on our preliminary designs with different AI-RPM technologies.
+Specifically, we want to explore (1) whether they find these AI-RPM technologies useful for remotely monitoring mental health-related information of youth concussion patients and (2) how to revise the preliminary system design from data type, visualization, interaction, and usability perspectives to better support clinicians in addressing challenges and decision-making in their workflow.
 5.1. Procedure
-We conducted user reviews with the same participants from our formative study.
-Due to the time constraints of concussion clinicians, each user review was limited to 10 minutes.
-We conducted a Zoom screen-sharing session to present our three modules to participants, introducing the data and functionalities included in each module.
-Based on each module, we then collected verbal feedback from participants regarding data type, visualization, interaction, usability, and potential use cases.
-Our research team documented their responses through note-taking.
-After the session, we collected participants’ feedback and applied axial coding to identify key themes.
-We then identified design modifications based on the themes.
-Next, we iterated on our initial system design following design modifications to come up with the final version of our system design.
-Our goal is to ensure that the final system design shows the potential to effectively support concussion clinicians in addressing the challenges of detecting mental health sequelae.
-The detailed interview protocol for our design interface evaluation is provided in Appendix B.
+We conducted an evaluation study with the same six clinicians from our formative study since they are the target users of the AI-RPM system.
+Due to the significant time constraints faced by concussion clinicians, each evaluation session was limited to 10 minutes with each clinician.
+We conducted a Zoom screen-sharing session to present our preliminary designs to each clinician and collected verbal feedback regarding data type, visualization, interaction, and usability.
+Meanwhile, our research team documented their responses through audio recording and note-taking.
+After the session, two researchers applied axial coding to identify codes and key themes.
+Afterward, we iterated on our preliminary designs based on our findings from this design interface evaluation.
+The detailed interview protocol for our evaluation study is provided in Appendix B.
 5.2. Findings
-In this section, we presented the feedback of concussion clinicians on the initial system design, including the features they found helpful, their concerns and points of confusion, and suggested improvements to the system.
-5.2.1. AI Risk Score Prediction as Supportive Evidence for Clinicians in Patient Communication
-Concussion clinicians (P1, P5, P6) reported that if the AI risk score prediction module provides accurate data, it can be useful in assisting clinicians with treatment planning and making timely adjustments to clinical decisions, such as issuing referrals or a set of follow-up visits.
-”That [the AI risk score prediction] might be something helpful early on for them to say, Hey, this is somebody who probably gonna benefit from seeing another provider earlier rather than later.” (P5)
-Moreover, P1 mentioned that the AI Risk Score Prediction module is straightforward and has the potential to effectively support clinicians in communicating with patients and their parents when concerns arise about patients’ mental health sequelae during clinical visits.
-By directly presenting the risk scores of potential mental health outcomes, youth patients and their families gain a clearer understanding of both the current mental health state and the possibility of potential future mental health sequelae, particularly for youth patients and families who are hesitant to discuss mental health issues with concussion clinicians.
-”I think it’s pretty simple, and if I was able to show that to a family… maybe something as simple as this to show [the patient’] mom… ’I’m concerned about your son… There’s a 50% risk that may develop [mental health sequelae].
+In this section, we present five key findings derived from concussion clinicians’ feedback during the evaluation study.
+5.2.1. Improving Patient-Clinician Communication With AI Risk Score Prediction
+Concussion clinicians (P1, P5, P6) reported that if the AI risk score prediction module provides accurate data, it can be very useful for timely clinical adjustments, such as scheduling follow-up visits or providing referrals.
+“That (the AI risk score prediction) might be something helpful early on for them (concussion clinicians) to say, hey, this is somebody who probably gonna benefit from seeing another provider earlier rather than later.” (P5)
+Moreover, P1 mentioned that the AI risk score prediction module is not only easy to understand but also has the potential to serve as an evidence-based tool to effectively support clinicians in communicating with patients and their parents about mental health conditions during clinical visits.
+By directly presenting the risk scores of the likelihood of developing mental health sequelae, youth patients and their families may gain a clear understanding of the severity of patients’ mental health conditions and be willing to talk about it, particularly for youth patients and families who are hesitant to discuss mental health issues with concussion clinicians.
+“I think it’s pretty simple, and if I was able to show that to a family… maybe something as simple as this to show [the patient’s] mom… ’I’m concerned about your son… there’s a 50 percent risk that [your son] may develop [mental health sequelae].
 We should really be in touch with it.’ Yes, I think this is great.” (P1)
-5.2.2. Explanation of AI Results and Data Sources as Supportive Information for Clinicians’ Understanding
-However, concussion clinicians expressed confusion regarding the data sources and the interpretation of results in certain modules.
-First, they raised concerns about the AI Risk Prediction module.
-When we presented this module, despite the interface providing key contributing data and a brief explanation of data sources, clinicians still found it difficult to understand the meaning of the risk score.
-They were also unsure whether the score was derived from other modules or from self-reported questionnaires administered to patients, such as the Generalized Anxiety Disorder 7-item Scale (GAD-7) and the Patient Health Questionnaire-9 (PHQ-9).
-Additionally, clinicians expressed confusion about the self-reported symptoms module.
-Specifically, P2 and P4 mentioned that they were unclear about how self-reported symptoms were collected.
-”So it’s just my confusion.
+However, P3 found the AI-generated risk score unhelpful because it covered one year, while she only treated youth concussion patients for four weeks.
+If patients did not recover within this period, they would be referred for further treatment.
+5.2.2. Supporting Clinicians’ Understanding with AI-Generated Results and Data Source Transparency
+Concussion clinicians expressed confusion regarding the data sources and AI results in both the AI risk score prediction modules and self-reported symptoms modules.
+When we presented the AI risk score prediction module, P6 sought for a clear explanation of the meaning behind the AI risk score despite the interface already providing key data that contribute to the risk score as well as a brief explanation of the score.
+Additionally, P2 and P4 mentioned that they were unclear about how patient’s mental health symptoms were collected in the self-reported symptoms module.
+Specifically, P2 was unsure whether the symptoms (e.g., suicidality and depression), were from AI risk score prediction modules or from questionnaires (e.g., GAD-7 and PHQ-9) that clinicians gave to their patients for self-assessment.
+However, the symptoms are collected by LLM-powered CAs through check-in conversations with youth patients.
+“So it’s just my confusion.
 So when it says self-reported symptoms that’s based on like a questionnaire that we’ve given them about their depression and anxiety level?
-Or this is what AI predicts that they have a 50 percenty chance of like developing depression based on that?” (P2)
-5.2.3. CAs and Wearables as Additional Valuable Data Sources Outside the Clinic
-Concussion clinicians (P3, P6) expressed optimism about the potential of the self-reported symptoms module.
-They believed that patients might be more willing to share information with a voice assistant than directly with a clinician, making this module a valuable additional data source.
-One clinician mentioned that they would like to quickly review self-reported symptom data before each clinical visit to better understand youth patients’ current conditions.
-At the same time, P6 noted a potential challenge: if patients have a voice-based conversational agent at home, they might place it in a corner and forget about it.
-To address this, they suggested that CAs should initiate daily conversations proactively to collect patient information more effectively.
-”I guess from a perspective of maybe they would tell this device things that they wouldn’t tell me. Possibly I could see that.” (P6)
-Among the three modules in our system design, concussion clinicians (P3, P4, P5) particularly appreciated the sleep and physical activity module.
-They highlighted that such data is currently missing from electronic health record (EHR) systems, yet wearable-generated data on sleep and physical activity could serve as key indicators for monitoring adolescent concussion patients’ mental health and recovery progress.
-P3 believed that this information could help monitor patients’ health on a weekly basis and assess whether they have followed clinical recommendations.
-However, P6 raised a concern that if a patient already owns a smartwatch, wearing a second device for physiological monitoring might be inconvenient.
-Additionally, having two devices could serve as a constant reminder of their injury, potentially reinforcing the perception that they have not yet fully recovered.
-Additionally, P3 suggested that it would be best if patients did not see the data collected by the smartwatch to avoid reminding them of their condition.
-”This one (the Sleep and Physical Activities panel) to me might be the most helpful [one]… I think that’s always hard [to get].” (P5)
-”So I really like the sleep and physical activity, but I would suggest that the patients not see that data that it only goes to the physician.
-Because I don’t want it to cause them anxiety.
-So yeah, I really like that.” (P3)
-5.2.4. Variations in Data Detail and Presentation Types for Clinicians’ Preferences
-Concussion clinicians exhibited varying preferences regarding the type and level of detail in the data presented within the sleep and physical activity module.
-Some clinicians preferred more detailed information than what was currently displayed.
-For instance, clinicians (P2, P3, P5) suggested including details related to napping under the sleep pattern section, such as whether the patient took naps and their duration.
-Additionally, clinicians (P1, P3) emphasized the importance of heart rate in assessing concussion recovery and monitoring mental health sequelae.
+Or this is what AI predicts that they have a 50 percent chance of like developing depression based on that?” (P2)
+5.2.3. Enriching Clinicians’ Insights with Data from LLM-based CAs and Wearables
+Clinicians reported that the data collected from CAs and wearables could be valuable in their workflow.
+Specifically, P6 mentioned that he would have a quick review of data on the self-reported symptoms module collected by CAs before each clinical visit to better understand youth patients’ current health conditions.
+Moreover, concussion clinicians (P3, P6) believed that patients might be more willing to share information with a conversational agent than directly with a clinician, which makes the data collected by CAs a valuable additional data source.
+However, P3 suggested that rather than using CAs to ask patients about their mental health conditions daily, it would be more appropriate to collect the data weekly.
+This suggestion helps avoid constantly reminding patients of their injuries.
+“I guess from a perspective of maybe they (patients) would tell this device things that they wouldn’t tell me. Possibly I could see that.” (P6)
+Moreover, concussion clinicians (P3, P4, P5) particularly appreciated the data in the sleep and physical activities module collected by wearables.
+They highlighted that data such as sleep hours is currently missing from EHR systems, yet wearable-collected data could serve as key indicators for mental health sequelae and concussion recovery progress.
+Additionally, P3 believed that clinicians could use the sleep and activity data to assess whether patients have followed sleep and exercise recommendations, which can be useful for addressing one of their main challenges.
+“This one (the Sleep and Physical Activities panel) to me might be the most helpful [one]… I think that’s always hard [to get].” (P5)
+5.2.4. Diverse Clinician Preferences in Data Details and Presentation Styles
+Concussion clinicians showed varying preferences regarding the type and level of detailed data presented within the sleep and physical activity module.
+Clinicians (P2, P3, P5) suggested incorporating napping-related data into the module, including whether the patient took naps and their duration.
+Additionally, clinicians (P1, P3) emphasized the importance of heart rate in assessing concussion recovery and monitoring mental health conditions.
 Specifically, they expressed interest in viewing daily resting heart rate during sleep over the course of a week.
-They noted that during recovery, a patient’s heart rate tends to be lower, indicating a calm state, and it often correlates with obtaining eight hours of quality sleep.
+P1 noted that a heart rate in the fifties to sixties indicates a calm state in patients.
+A calm state is often associated with obtaining eight hours of quality sleep.
 In contrast, if a patient experiences fragmented or restless sleep, their resting heart rate increases.
-By analyzing these patterns, clinicians could better assess whether a patient is experiencing anxiety.
-”If you (patients) are having a light, irritable kind of sleep, I would expect the resting heart rate to be in the seventies or eighties.
+By analyzing physiological data, clinicians can assess if symptoms are linked to mental health or other non-concussion issues.
+“If someone’s recovering, I would expect it (heart rate) to be low, [meaning] they’re calm……If you (patients) are having a light, irritable kind of sleep, I would expect the resting heart rate to be in the seventies or eighties.
 [It means] maybe they (patients) are anxious.
-Maybe there’s something else going on.
-” (P1)
+Maybe there’s something else going on.” (P1)
 Moreover, in the self-report symptoms module, clinicians (P3, P5) suggested they want to include dietary information, such as appetite or nutrition, as they noted that when youth patients experience depression or high levels of anxiety, both the patients and their parents often report decreased appetite.
-Regarding the symptoms part of the module, instead of focusing solely on whether youth exhibit symptoms of mental health sequelae, concussion clinicians (P3, P5) emphasized that they are more concerned with the duration and severity of these symptoms.
-The severity and duration of symptoms have different impacts on clinicians’ treatment plans and decision-making, which potentially helps clinicians distinguish between general emotional issues and mental health sequelae following a concussion.
-”…persistent or non-improving symptoms. That’s what I look for.”(P5)
-Additionally, P3 expressed a new need for a feature that could indicate how adherence to recommendations impacts recovery rates, suggesting that such insights could enhance clinical decision-making.
-”I don’t know if my patients are following my recommendations and knowing if they are and how that affects their recovery would be really interesting.” (P3)
-5.2.5. Clinicians’ Concerns About Emergency Situation and Time Constraints
-In the self-reported symptoms module, concussion clinicians placed significant emphasis on alerts related to suicidality.
+Importantly, instead of focusing solely on whether youth patients have mental health symptoms, concussion clinicians (P3, P5) were more concerned with the duration and severity of these symptoms.
+Because the severity and duration of symptoms have different impacts on clinicians’ decision-making, such as treatment plans.
+“I don’t think it has to be high level [of mental health symptoms]…… persistent or non-improving symptoms. That’s what I look for.” (P5)
+Additionally, P3 expressed a new need for a feature that could indicate how adherence to sleep and exercise recommendations impacts patients’ recovery, suggesting that such insights could enhance clinical decision-making.
+5.2.5. Concerns of Clinicians in Emergency Situations and Time Constraints
+In the self-reported symptoms module, concussion clinicians placed significant emphasis on alerts related to suicidal tendencies.
 P1 emphasized the need for immediate intervention when suicidal tendencies emerge in youth concussion patients.
-However, they also highlight a critical gap in current clinical practice—the absence of established protocols for managing or intervening in such situations.
-When the user interface flags a potential suicide risk, clinicians report uncertainty about how to respond effectively due to a lack of concrete, actionable guidelines.
-Failure to intervene in a timely manner could result in missed opportunities for intervention and serious legal or ethical consequences.
-Additionally, even when clinicians recognize signs of suicidal ideation, managing emergency interventions without triggering resistance from parents or youth patients remains a highly sensitive and complex challenge.
-Furthermore, P4 reported that they face a dilemma: The youth generation experiences frequent mood swings and may sometimes exaggerate their feelings and symptoms.
-If a youth patient reports self-harm intentions that turn out to be overstated, clinicians worry about overreacting to false alarms.
-Striking the right balance between taking reports seriously and avoiding unnecessary escalations is a persistent concern in their practice.
-Clinicians (P2, P4) generally believe that parents should be brought into the loop when such situations arise.
-”So I think this is when you need to loop the paradigm and say, hey, your child has been identified through our software.
+However, clinicians expressed several challenges.
+Firstly, there is an absence of established protocols for managing such emergency situations.
+Furthermore, P4 mentioned that the youth generation experiences frequent mood swings and may sometimes exaggerate their feelings, which could lead to false alerts.
+However, failure to intervene in patients who have suicidal tendencies in a timely manner could result in a life-or-death consequence.
+To address this, clinicians (P2, P4) believed that youth patients’ parents should be brought into the loop and have a partnership with parents when emergency situations arise.
+“So I think this is when you need to loop the paradigm and say, hey, your child has been identified through our software.
 You know, cause I think, especially since we’re dealing with minors, you’re gonna have to get the parents loop too.” (P4)
-Moreover, clinicians (P2, P5) expressed concerns about their already heavy workloads, stating that they have little time to thoroughly review such detailed data.
-These clinicians noted that having dedicated staff to help filter and process data would greatly reduce their burden.
-Without such support, they often skim through relevant data quickly on the day of the appointment or shortly before, just to gain a basic understanding of the youth concussion patient’s condition.
-”I think that’s gonna that would be really tedious to try to monitor and create extra work for us.”(P5)
-Regarding the newly introduced data and user interface modules, concussion clinicians (P1, P2) expressed concerns about the difficulty of finding time to process the data thoroughly.
-Clinicians already feel overwhelmed by the excessive features within current electronic health record (EHR) systems and their tasks.
-P1 highlighted that processing large amounts of data between tightly scheduled appointments—sometimes up to 20 patients per day—makes it nearly impossible to dedicate sufficient time and energy for in-depth analysis.
-Although monitoring systems can collect vast amounts of valuable data, clinicians often lack the capacity to fully review this information within their workflows.
-Given their heavy workload, the time clinicians can dedicate to each patient is extremely limited—sometimes as little as 15 minutes per appointment.
-As a result, concussion clinicians must rely on brief time windows, such as the moments before a scheduled meeting, to extract key information and quickly assess the condition of youth concussion patients.
-”I would need to look at it on that day of the visit… Honestly, we are quite overwhelmed with the EHR. And all the different features and everything that we’re asked to do with it.”(P1)
-6. Refined System Design
-Based on the design interface evaluation, clinicians acknowledged the value of AI-RPM technologies which can potentially help them remotely collect and monitor patient data.
-In addition, clinicians found that the data, features, and presentation methods in our initial system interface could support their mental health-related decision-making.
-At the same time, they provided valuable feedback and suggestions based on the initial system design.
-In this section, we present design considerations derived from the design interface evaluation and use them to refine our initial system design.
-We then provide a refined system design in this section.
-The goal of presenting this refined system design is to visually illustrate the design considerations more effectively and provide insights for future studies.
-6.1. Design Considerations
-6.1.1. Explanation of Data Sources and Module Relationships
-Through design interface evaluation, we found that while concussion clinicians considered the risk score useful for health decision-making and communication with patients and their families, they also raised concerns.
-Specifically, they were unsure how the score was generated and how it connected to other modules.
-To address this issue, we recommend using visualizations to illustrate the relationships between AI results and other data sources.
-This would enable clinicians to intuitively understand how various factors influence the risk score, potentially enhancing their trust in the system.
-One clinician explicitly mentioned wanting to see the relationship between clinical recommendations and patient recovery.
-So we suggest visualizing this by displaying the correlation between the risk score and exercise recommendation compliance.
-6.1.2. Personalized Time Selection
-One clinician found the AI-generated risk score unhelpful because the initial UI displayed a one-year risk score, whereas she typically treats youth patients for only four weeks.
-If a patient does not recover within this period, they are referred for further evaluation.
-Therefore, we believe clinicians would benefit from viewing AI model risk scores over different time spans to meet their varying needs and align with their workflow.
-To accommodate different clinical needs, we propose adding a time range selector to the module.
-This feature would allow clinicians to customize the risk score display based on their specific decision-making timeframes, offering personalized support for different clinical workflows.
-Furthermore, for the sleep and activities module in Section 5.2.3, a clinician expressed a preference for viewing data on a weekly basis.
-To align with this preference, we recommend that the module default to displaying seven days’ worth of data while also providing the option to explore detailed data for specific dates.
-This would allow clinicians to closely examine anomalies, particularly when irregular patterns appear on certain days.
-6.1.3. Providing Information on Other Stakeholders to Support Collaboration
-Concussion clinicians noted that in the self-reported symptoms module, the most critical yet challenging situation for clinicians is detecting suicidality in patients.
-However, they expressed uncertainty about how to remotely respond if a patient exhibits self-harm tendencies.
-Additionally, clinicians raised concerns about the authenticity of such reports, as youth patients may sometimes exaggerate their symptoms.
-Clinicians emphasized that if an emergency occurs, parents should be involved, as they are considered key stakeholders in managing such situations.
-In response to this concern, we propose integrating an emergency contact module within this UI module, which allows clinicians to contact parents when necessary.
-Since parents may have quicker access to their children and be able to take immediate action, this feature could enhance response efficiency.
-Additionally, clinicians expressed a need to track the number of days symptoms persist, as this directly reflects a patient’s recovery progress and informs mental health-related decisions such as giving a referral.
-We recommend that when CAs engage in conversations with patients and their parents, they also document their observations on the duration of mental health symptoms.
-Then from the design side, we propose adding a data point in the module to display symptom duration.
-In addition to monitoring the patient’s emotional state, clinicians often ask parents about behavioral changes.
-To enhance data comprehensiveness and objectivity, we suggest incorporating parent-VA (virtual assistant) conversation data into this module, offering an additional perspective on the patient’s condition.
-Finally, to reduce confusion and improve the module’s usability for clinicians, we need to clearly define how the data is collected.
-Ensuring transparency in data sources will help clinicians trust and effectively utilize this module in their practice.
-6.1.4. Balancing Additional Data Presentation and Cognitive Load
-In the initial system design, certain key data types that are highly important to clinicians were not covered.
-As section 5.2.4 mentioned, concussion clinicians emphasized the need for heart rate during sleep, nap times, wake-up time, and sleep time rather than just total sleep hours.
-To address this, we recommend displaying these additional data points within the module.
-However, given clinicians’ concerns about data overload—especially as they noted that existing electronic health record (EHR) systems have already increased their workload—we propose a customizable data selection feature.
-Within the system interface, clinicians can personalize their data view based on their specific needs and workload at different times, rather than being presented with all data at once.
-Additionally, we believe it is essential to prioritize the most relevant information in the default display.
-Specifically, the most critical data will be shown by default, while supplementary explanatory details will be secondary and only accessible when needed.
-This design approach aims to reduce cognitive load, ensuring that clinicians can quickly access key insights without being overwhelmed by excessive information.
-By minimizing cognitive burden, this method enables more efficient and personalized analysis.
-6.1.5. Integrating Related Data into a Unified System Interface
-In our formative study, clinicians mentioned using GAD-7 and PHQ-9 to monitor patients’ anxiety and depression.
-Additionally, as Section 5.2.2 mentioned, clinicians noted that they use these questionnaires to patients again during the design interface evaluation, which makes us realize that we could integrate all the information clinicians want to see into the same system interface.
-Based on this, we propose adding a questionnaire results module to the system interface, displaying patients’ completed questionnaires before their clinical visits.
-This integration would allow clinicians to access all relevant data in a single interface before a consultation, saving time, improving data review efficiency, and enabling a faster and more evidence-based assessment of patients’ mental health status.
-Importantly, these questionnaire results differ from the anxiety and depression statuses identified in the self-reported symptoms module.
-They provide a more structured and standardized assessment.
-By combining subjective self-reported symptoms with objective questionnaire scores, clinicians can gain a more comprehensive understanding of patients’ mental health conditions, which would support more confident mental health-related decision-making, even when patients are outside the clinical setting.
-6.2. Visualization of Refined System Design
-In this version, we considered the possibility of seamlessly integrating the system design with existing electronic health record (EHR) systems in the future.
-Therefore, we referenced the design of current EHR systems to ensure consistency.
-To align with the visual and interaction patterns of EHR systems, we incorporated a patient list on the left side of the interface.
-A red dot next to a patient’s name indicates a potential risk of mental health sequelae, making it easier for clinicians to identify at-risk patients at a glance.
-The final system design consists of five modules, which aim to enhance system usability and support clinical decision-making.
-In the following sections, we will explain the data, functionality, potential interactions, and purpose of each module in detail.
-The final system design interface is shown in Fig.
-2.
-Figure 2.
-The final system design is based on feedback from concussion clinicians gathered through design interface evaluation: (A) patient basic information, (B) patient sleep and physical activities, (c) patient self-report symptoms, (D) patient questionnaire result, and (E) AI risk score prediction for mental health sequelae.
-6.2.1. The Basic Patient Information Module
-To better integrate seamlessly with the existing EHR system, we have added a basic patient information section in the upper-left module of our system (Figure 2A).
-This section includes key details such as the date of injury, the number of days since the concussion occurred, the youth patient’s mental health history, and medication usage.
-These pieces of information provide clinicians with essential insights into the patient’s condition, including the time elapsed since the injury and their mental health background.
-In particular, the latter two details—mental health history and psychiatric medication usage—are important for concussion clinicians when detecting patients’ mental health sequelae.
-6.2.2. The Sleep and Physical Activities Module
-In this module (Fig.2 B) concussion clinicians can hover over the question mark icon next to the module title to view its data source, which comes from wearable devices.
-Below the title, there are options to view data from the past week or the most recent day, with the default display set to one week of data.
-The module includes three selectable tabs: Sleep Data, Physical Activities, and Recommendation Compliance.
-The displayed information updates dynamically based on the selected tab.
-In the Sleep Data tab, clinicians can access heart rate, nap duration, sleep and wake-up times, and total sleep hours.
-They can select one or multiple metrics based on their specific needs.
+Moreover, clinicians (P2, P5) expressed that they have limited time to thoroughly review such detailed data due to heavy workloads.
+Specifically, P1 highlighted that processing large amounts of data between tightly scheduled appointments—sometimes up to 20 patients per day—makes it nearly impossible to dedicate sufficient time and energy for in-depth analysis.
+If they have the AI-RPM system, clinicians noted that they might skim through the most relevant data quickly on the day of the appointment or shortly before to gain a basic understanding of their patient’s condition.
+“I think that’s gonna that would be really tedious to try to monitor and create extra work for us.” (P5)
+6. Refined AI-RPM System Design
+Based on the findings in the design interface evaluation session, clinicians acknowledged the value of the system with AI-RPM technologies that can potentially help them remotely monitor patients’ mental health symptoms and support them to make decisions in their workflow.
+At the same time, they provided feedback and suggestions based on the preliminary designs.
+Based on their feedback, we revised our preliminary design and proposed a complete design of clinician-facing AI-RPM system design (Fig.
+1).
+6.1. Reducing Cognitive Burden Through EHR Integration
+To reduce the cognitive burden placed on clinicians by the new system and ensure the system can be seamlessly integrated into clinicians’ workflow, we referenced the style of current EHR systems to ensure consistency.
+We integrated a patient list (Fig.
+1 A), which already exists in the EHR system, into the left side of the interface.
+However, we added a red dot next to a patient’s name to capture clinicians’ attention if AI-RPM technologies detect abnormalities in the patient’s mental health conditions.
+Moreover, we added a basic patient information section, which exists in the current EHR system, in the upper-left module (Fig.
+1 B).
+However, the information here is tailored to concussion clinicians’ needs.
+These pieces of information provide clinicians with essential insights into a concussion patient’s condition.
+In particular, the Individual Psychiatric History and Mental Health Prescription are important for concussion clinicians when detecting youth patients’ mental health sequelae and making decisions.
+6.2. The Sleep and Physical Activities Module
+Based on findings from our evaluation study, concussion clinicians affirmed the value of the data provided in the Sleep and Physical Activities module (Fig.
+2 C) in the preliminary design.
+However, several clinicians requested more information about data sources.
+In response, we designed a question mark next to the module title (Fig.
+1 C).
+Concussion clinicians can hover over the question mark to view the data source.
+Moreover, because clinicians typically prefer a weekly overview of patient data, and some of them also want the ability to see daily metrics (e.g., heart rate), we provided the option to view data from both the last week and the previous 24 hours.
+With this option, concussion clinicians can easily track changes across different time intervals.
+In addition, clinicians requested the monitoring of additional metrics, such as heart rate during sleep and nap duration.
+To incorporate these data points without increasing cognitive load, we organized the data into three tabs: Sleep Data, Physical Activities, and Recommendation Compliance.
+The Sleep Data tab is presented by default, which allows clinicians to immediately access important metrics such as heart rate, nap duration, and sleep and wake-up times.
+They can choose to view one or multiple metrics, depending on their specific needs.
 Different data types are displayed in distinct colors to help clinicians visually distinguish between them more easily.
-When hovering over a data point in the chart, the exact value for that day is displayed, providing detailed insights into the patient’s sleep patterns.
-6.2.3. The Self-Report Symptoms Module
-On the right side of the module (Fig.2 C) title, concussion clinicians can hover over an icon to view the data source, which comes from CAs.
-The module provides two options for selecting a time range: the past week and the last 24 hours.
-Below the time selector, the interface is divided into two sections.
-On the left side, there are two tabs: one displaying VA-patient conversation history and the other showing VA-parent conversation history.
-Clinicians can expand the conversation panel by clicking the button in the top right corner to focus on historical dialogues.
-On the right side, the Symptoms Reported section visualizes symptom-related information.
-If the patient or parent has mentioned any symptoms, a colored indicator dot appears next to the corresponding symptom category, with the color reflecting severity—red for severe symptoms, and orange for moderate symptoms.
-If the patient reports no symptoms, the corresponding dot turns green.
-To the right of each category, the symptom duration is displayed, indicating how long the reported issue has persisted.
-Clinicians can also click AI Summary to view a detailed breakdown of the category, providing additional context to support their assessment.
-At the bottom of the Symptoms Reported section, we provide call buttons for contacting the patient or parent, which ensures that clinicians can quickly intervene in case of an emergency.
-6.2.4. The Questionnaires Result Module
+If clinicians want to delve deeper into a patient’s sleep patterns, they can hover over a data point in the chart, then a more detailed value for that day will be displayed.
+6.3. The Self-Report Symptoms Module
+To maintain consistency with the Sleep and Physical Activities module, we added question marks indicating the data source and provided time selection options in the Self-Report Symptoms module (Fig.
+1 D).
+Moreover, we identified emergent situations (i.e., suicidality) as both critical and challenging for clinicians to monitor, given resource constraints and the risk of false alarms.
+To address this, clinicians emphasized the importance of including parents in managing emergencies.
+Thus, we introduced two sections within this module.
+On the left side, we included the CA-parent conversation history, allowing clinicians to evaluate the patient’s health conditions from family perspectives.
+This conversation history can be expanded by clicking the icon next to the title ”Conversational History”.
+In addition, we added two call buttons at the bottom of the module, which enables clinicians to contact either the patient or a parent to manage emergency.
+Besides, clinicians confirmed the value of different categories (i.e., depression, anxiety) in this module but indicated the need for more information.
+In response, we retained the original categories and introduced new ones—such as nutrition level—based on clinician feedback.
+Moreover,clinicians mentioned about the importance of severity for their treatment plan.
+To meet this need, we placed color-coded dots next to each symptom category (red for severe symptoms, orange for moderate, and green for none).
+To the right of each category, we provided a symptom duration indicator to show how long the reported issue has persisted, which help clinicians form a better understanding of patients health conditions.
+Furthermore, clinicians can click AI Summary to view a detailed breakdown of the category, providing additional context to support their assessment.
+6.4. The Questionnaires Result Module
 In this module (Fig.
-2 D), the title includes an option to view data source information, which comes from patient-administered questionnaires at home.
+1 E), the title includes an option to view data source information, which comes from patient-administered questionnaires at home.
 These questionnaires include GAD-7 for anxiety, PHQ-9 for depression, and PCSS for concussion severity.
-The middle section of the chart presents the scores for each questionnaire along with corresponding explanations to help clinicians interpret the results.
+The middle section of the chart presents the scores for each questionnaire, along with corresponding explanations to help clinicians interpret the results.
 The module also displays a comparison with previous scores, providing clinicians with a sense of the patient’s recovery trend.
 Additionally, clinicians can click ”Show Details” to view a breakdown of individual section scores within each questionnaire for a more in-depth assessment.
-6.2.5. The AI Risk Score Prediction Module
+6.5. The AI Risk Score Prediction Module
 The final module (Fig.
 1 F) presents the AI-generated prediction of a patient’s risk of developing mental health sequelae.
-Concussion clinicians can hover over the question mark icon next to the title to view information about the source of the risk score.
-Based on the current predictive capabilities of the AI model, this module offers four different time range options, with the default set to risk predictions within four weeks.
-Below the time selector, clinicians can view the risk score chart for the selected period and hover over specific time points to examine the risk score at each point.
-At the bottom of the module, we display the features contributing to the risk score, along with the influence of each contributing factor, represented as a percentage.
-Visually, if a feature has a high contribution to the risk score, we provide a visual alert.
-The dot next to the feature title changes color based on the severity of the symptoms associated with that feature, with red indicating severe symptoms and orange indicating moderate severity.
-This severity level is also reflected in the percentage value on the right, using corresponding colors to enhance clarity.
-Additionally, the visual representation of each contributing module is color-matched to its corresponding section in the UI.
-This alignment helps clinicians intuitively understand the connections between different modules and their impact on the AI risk score.
+Given that concussion clinicians primarily manage patient care during the first four weeks following a concussion, we adjusted the default timeframe for risk predictions from one year to four weeks, while still providing other timeframe options for selection.
+Below the imeframe options, the visualization of the risk score chart remains unchanged from the preliminary design, given that clinicians reported it to be easy to interpret.
+However, clinicians expressed an interest in knowing which modules or features influence the AI-generated risk score.
+To highlight the relationship between the AI risk score and other modules, we introduce Features Contributing to the AI Risk Score section at the bottom of the module.
+The section includes contributing features and its corresponding module, along with a contribution rate.
+Moreover, we provide visual feedback on each feature’s contribution to the AI risk score.
+Both the color of the dots next to the feature and the font color of its associated percentage value dynamically change based on the degree of the contribution.
+In this way, clinicians can identify the important metrics quickly.
 7. Discussion
-In this work, we identified multiple challenges that concussion clinicians face when treating youth concussion patients with mental health sequelae, including the difficulties of collecting patients’ information outside the clinics, communicating with patients and their families, as well as making timely clinical decision-making about concussion treatment based on mental health symptoms.
-Further, we proposed multiple AI-RPM technologies accompanied by AI prediction models to support concussion clinicians’ clinical workflow and iteratively refined a clinician-oriented EMR dashboard user interface to consolidate detailed feedback on information representations, interactions, customizability, and multiple responsible AI considerations.
-In this section, we delve into how the aforementioned challenges arise and explore how they can potentially be addressed through AI-RPM technologies.
-Additionally, we discuss the role of AI-RPM systems in supporting clinicians’ decision-making processes.
-Finally, we talk about concerns regarding responsibility in the potential use of AI-RPM systems and identify design opportunities to tackle these challenges.
-7.1. Remote Monitoring of Patient Mental Health-Related Nonclinical Data
-7.1.1. Concussion Clinicians’ Need of Monitoring Patients’ Mental Health Symptoms
-A unique challenge faced by concussion clinicians, which was identified in our work, is that clinicians find it critical but difficult to assess the severity of patients’ mental health symptoms when they are outside the clinic.
-For concussion patients, there is a significant overlap between concussion symptoms and mental health symptoms, and these two illnesses could influence each other’s recovery progress.
-Although concussion clinicians do not possess expertise in the diagnosis and treatment of mental health issues, their clinical decision-making for concussion treatment is closely correlated with patients’ mental health symptoms.
-If patients exhibit severe mental health symptoms, concussion clinicians need to adjust the concussion treatment plan and the scheduling for follow-up visits accordingly and refer the patients to mental health specialists for professional intervention.
+In the following sections, we first discuss the use of AI-RPM technologies in the remote monitoring of patient mental health symptoms.
+Then, we talked about collaborative design in AI-RPM systems for enhancing usability and effectiveness.
+Finally, we conclude with collaborative emergency management with AI-RPM systems.
+Our work contributes to CSCW and HCI research at the intersection of health, RPM, and human-AI decision-making.
+7.1. Unraveling Patients’ Intertwined Mental Health and Concussion Symptoms
 In clinical scenarios, it is not uncommon to have multiple illnesses intertwined with each other.
-For instance, a recent work discovered that cancer oncologists need to be constantly aware of patients’ cardiotoxicity during cancer treatment and collaborate with cardiologists to adjust the cancer treatment plan (Wu et al., 2024).
-Nevertheless, these two scenarios are fundamentally different.
-Based on our formative study, concussion clinicians do not collaborate with mental health specialists, and neither do they diagnose mental health disorders.
-Concussion clinicians, in most of their clinical practices, only need to detect whether patients’ mental health issues surpass a certain level of severity threshold.
-If the mental health issues are severe, concussion clinicians will refer patients to mental health specialists for mental health intervention.
-Otherwise, concussion clinicians will not do the referral but take the patient’s mental health status as a factor in the design of concussion recovery recommendations.
-We believe the integrated design of leveraging AI-RPM technologies to collect patients’ mental health-related data outside the clinic and then developing AI-based predictive models to assess patients’ mental health symptoms could provide actionable insights into concussion clinicians’ current workflow.
+For instance,  Wu et al.
+(2024) discovered that cancer oncologists need to collaborate with cardiologists to adjust the cancer treatment plan and be constantly aware of patients’ cardiotoxicity.
+Nevertheless, the concussion-mental health scenario and the cancer-cardiotoxicity scenario are fundamentally different.
+Concussion clinicians, according to our study, do not collaborate with mental health specialists, and neither do they diagnose mental health disorders.
+In most clinical practices, concussion clinicians assess whether a patient’s mental health symptoms exceed a certain severity threshold during in-person visits.
+Based on this assessment, they adjust the treatment plan accordingly.
+We believe that AI-RPM technologies have the potential to give clinicians insights into the severity of patients’ mental health symptoms.
+First, LLM-based CAs could provide insight into patients’ mental health symptoms, such as the level of anxiety.
+Moreover, AI-based predictive models could help assess the likelihood of developing mental health sequelae, which can offer valuable and actionable insights into patients’ mental health conditions.
 As suggested by Bennett et al.
 (2012); Zhang et al.
-(2024), clinicians expect AI predictive models to provide actionable insights, which is critical to ensure AI models’ usefulness in clinicians’ workflow.
-Future research could further explore the design of AI-RPM technologies with respect to what other modalities of patients’ data could be beneficial for mental health assessment and further develop AI models to integrate such data, along with contextual information like medical histories and behavioral patterns for more reliable predictions with sufficient explainabilities.
-7.1.2. Remote Monitoring of Objective Patient Data
-Our study revealed significant challenges in how concussion clinicians assess mental health issues in youth concussion patients.
-Youth patients often avoid openly discussing mental health concerns due to the stigma of being labeled with mental health disorders.
-Consequently, clinicians are often limited to collecting patient information through mental health assessment questionnaires or direct observation of patients’ behaviors during clinical visits.
-However, if patients are unwilling to communicate openly, the validity of questionnaires can be compromised, and observation during clinical visits can be left as the only viable option.
-However, direction observation is inherently limited from the geographical perspective.
-Once patients leave the clinic, concussion clinicians can hardly track patients’ mental health status in a timely and effective manner.
-Even if patients are willing to faithfully and openly discuss their mental health status, patients’ memories could be fragmented and inaccurate, especially for sleeping patterns.
-We suggest that RPM technologies, particularly wearable devices like smartwatches, could help address the aforementioned limitations by providing clinicians with real-time access to key physiological metrics.
-In particular, these physiological metrics are objective measurements that can faithfully reflect patients’ mental health status.
-Prior research demonstrated the feasibility and advantages of wearable in monitoring objective patient information outside the clinic (Mendel et al., 2024; Frazier et al., 2023; Su et al., 2019; Wu et al., 2024).
-Wearable devices could monitor objective data such as sleep patterns, physical activity levels, and resting heart rate, which are critical signals for tracking and assessing patients’ mental well-being.
-For instance, elevated anxiety levels may be inferred through reduced physical activity or disrupted sleep patterns, while a stable resting heart rate could indicate recovery progress.
-Concussion clinicians could use such data for clinical decision-making, like scheduling timely follow-up visits or initiating referrals to mental health specialists as needed.
-An additional benefit provided by wearable devices is that they can monitor patients’ status non-intrusively to be seamlessly integrated into patients’ daily routines.
-This non-intrusive characteristic minimizes potential cognitive burdens added to the patients, where patients will only need to remember to charge the wearables regularly.
-7.1.3. Remote Monitoring of Subjective Patient Data
-Clinicians in our study emphasized that the severity and duration of symptoms such as anxiety, depression, or even suicidal ideation are critical factors in assessing whether youth concussion patients are experiencing mental health sequelae.
-However, these symptoms are more of subjective experiences that can hardly be captured using wearables because the data can be collected by wearables are standardized objective physiological measurements.
-To address this gap, we propose integrating advanced generative AI with RPM technology to gather patients’ subjective data through text-based or voice-based conversational assistants (CAs).
-These CAs are supported by large language models (LLMs) that could engage in free-form natural language conversations with patients.
-Compared with traditional RPM approaches, natural conversation significantly reduces the requirement of patients’ technical literacy to engage with the devices, enables flexible patient self-reporting, and can be further designed to conduct regular health check-ins through pre-set question lists.
-Recently, a number of works have demonstrated the promising opportunities of LLM-based CA to support various patient populations in healthcare scenarios (Bartle et al., 2023, 2022; Wang et al., 2023a; Simpson et al., 2020; Ma et al., 2024; Yang et al., 2023), for example, to support older adults with chronic diseases (Yang et al., 2024).
-Using AI-driven analytics, key psychological risk factors such as suicidal thoughts could be automatically detected.
-LLMs could then generate a summary report of the patient’s mental health for the day and transmit it remotely to the clinicians’ EMR system.
-By combining subjective data from CAs with objective information collected by wearables, clinicians could gain a more comprehensive understanding of the patients’ mental health in their daily lives.
-Subjective information often provides contextual insights for objective data, which enables clinicians to make more accurate judgments.
-For example, if wearable data shows that a patient had very low physical activity on a given day, the CA might reveal through its conversation that the patient’s mental health was stable and the low activity was due to a family gathering at home.
-This additional context would allow the clinician to assess the situation more accurately, avoiding unnecessary concern over what might otherwise appear to be a worrisome pattern in objective metrics.
-By leveraging the strengths of both data types, our proposed AI-RPM approach could enhance the quality and precision of mental health assessments in youth concussion care.
-7.1.4. Patient Privacy Consideration With AI-RPM Technologies
-Lastly, when designing such an AI-RPM system, it is essential to underscore protecting patient privacy and ensuring data security.
-This issue has been widely discussed in many studies and is recognized as a core challenge in AI system design (Gawankar et al., 2024; Bala et al., 2024; Korobenko et al., 2024).
-However, in our research context, the target population comprises youth patients aged 11–17, whose privacy needs are uniquely sensitive.
-The privacy of youth patients is not only a concern for the patients themselves but is also critical to their parents, who often play a key role in the medical process.
-Parents typically expect to have transparency and informed consent regarding how their child’s data is collected and used (Sisk et al., 2020; Haley et al., 2024).
-Therefore, finding ways to protect patients’ privacy while building and maintaining parental trust in the data collection and usage process is a critical issue that requires careful consideration and balance.
-Future research could explore methods to effectively safeguard youth health data while simultaneously building trust and participation among parents.
-7.2. Multi-Stakeholder Collaboration in the Design of Personalized and Accountable AI-RPM Systems
-7.2.1. Collaboration With Clinicians in Supporting Clinical Workflow
-Our findings highlight that while clinicians are open to using AI-RPM systems, they are concerned about the potential for information overload.
-Clinicians are already overwhelmed by the variety of data and tasks within existing EHR systems.
-Introducing another AI system with additional data could significantly increase their workload, potentially reducing the number of available weekly appointment slots and further straining already limited healthcare resources.
-In our study, clinicians generally believe they do not have the time to process such detailed information in their daily work.
-However, clinicians still expressed willingness to check the data collected by the AI-RPM systems before each clinical visit because they believe such data could support them better prepare for patient consultations.
-To address this concern, AI-RPM systems need to be designed in collaboration with clinicians to prioritize personalized and context-relevant data presentation.
-One approach to achieving this is through layered information presentation, where critical data is prioritized and displayed prominently, while secondary information remains accessible but hidden by default.
-For example, clinicians could be shown a summary of essential metrics, such as daily activity levels, sleep patterns, and key symptom trends, while additional context or raw data could be available on demand (e.g., through interactive elements on the interface).
-Designers must collaborate closely with clinicians to determine which data points should be prioritized and how they should be presented to support efficient clinical decision-making.
-Clinicians in our study also emphasized the importance of defining thresholds for symptom severity, particularly in distinguishing between mild, moderate, and severe symptoms.
-Some concussion clinicians mentioned that they might not be particularly concerned about mild anxiety, but severe or prolonged symptoms would require their attention.
-Without clear definitions, AI systems may either underreport or overemphasize symptoms, leading to delayed interventions or unnecessary concerns.
-Clinicians are the domain experts who can help designers bridge the gap with their knowledge and expertise.
-By working with clinicians to establish these thresholds, AI-RPM systems can more effectively categorize symptoms, ensuring that clinicians are alerted only when necessary.
-Such a collaborative approach will not only reduce cognitive overload but also build trust and confidence in the system’s reliability, which will allow clinicians to rely on the AI system as a supportive tool rather than a burden.
-Ultimately, the goal of engaging concussion clinicians in the collaborative design process is to enable the system for clinicians to access the most relevant patient information quickly, improve the efficiency of their workflows, and make informed decisions without being overwhelmed by data.
-By integrating clinicians’ preferences and needs into the system design, AI-RPM systems can enhance their usability and effectiveness, promoting better human-AI collaboration in clinical practice.
-7.2.2. Collaboration with Parents in Managing Emergency Situations
-In the proposed design section of self-reported symptoms, there is a ”suicidal ideation” alert module which clinicians believe would be of significant benefit to alert clinicians for any unexpected emergency.
-However, clinicians expressed considerable concerns about this feature.
-Clinicians questioned the accuracy of the ”suicidal ideation” alerts, not solely in terms of the technical capabilities of the AI system but more importantly regarding the reliability of the self-reported data.
-In our research context, clinicians pointed out that at the developmental stage, youth patients may often express emotions with exaggeration.
-For example, youth patients might say things like ”I want to hurt myself” or ”I have suicidal thoughts,” which might not necessarily reflect their actual intentions.
-Such statements could result in a false alarm when captured by the AI system and misclassified as a ”high-risk alert.” The implications of such false alarms are multifaceted.
-First, responding to a high-risk alert would require clinicians to dedicate considerable time and resources to contact the patient and address the situation, even if it turns out to be unwarranted.
-False alarms could lead to a waste of a clinician’s valuable time and critical medical resources that could be allocated to other patients.
-Second, if false alarms occur frequently, clinicians could gradually lose trust in the system, which eventually leads to the abandonment of such a system.
-Another concern expressed by clinicians is the ambiguity of responsibility that AI systems introduce in medical decision-making.
-Previous studies have highlighted the phenomenon where clinicians, as final decision-makers, have full responsibility for medical outcomes, while AI systems themselves are not held accountable.
-This issue was also raised in our study.
-Specifically, if the AI system triggers an emergency alert (e.g., indicating a patient might have suicidal ideation), clinicians face multiple constraints in responding.
-These constraints might include a lack of resources or time to address the alert promptly, or situations where clinicians fail to review the AI system’s information in a timely manner.
-Any subsequent negative outcome, such as a life-threatening incident, raises the question of who should be held accountable.
-We believe that in a high-risk and high-uncertainty medical scenario where the patients are young, it’s a key challenge to ensure the factuality of emergency alerts and find a balance between handling alerts and clinicians’ responsibilities throughout the development lifecycle of an AI system.
-Addressing this challenge requires collaboration among a broader set of stakeholders.
-For instance, involving the youth patient’s family in handling emergencies could play a crucial role.
-If the AI system detects an emergency, such as suicidal ideation, the system should notify not only the clinician but also the patient’s family immediately because family members could take appropriate action in cases where the clinician cannot respond promptly.
-For example, family members could calm the patient, take them to a healthcare facility, or contact mental health professionals for intervention.
-Additionally, clinicians who receive such alerts could collaborate with the family by providing timely professional guidance.
-Such notification and collaboration mechanisms would not only enable families to gain a more comprehensive understanding of the patient’s condition but also reduce risks arising from delayed clinical responses.
-Engaging parents in the emergency care-providing process ensures patient safety while alleviating some of the pressure on clinicians who might be constrained by limited time or resources.
-To address ambiguity in responsibility for medical decisions, collaborative efforts involving hospital administrators, legal experts, and researchers are needed to develop clear guidelines on accountability in AI-supported clinical scenarios.
-Stakeholders should clearly define roles and accountability in emergencies triggered by AI systems to address the potential liability concerns highlighted in existing studies (Goodman and Flaxman, 2017).
-An addition benefits would be helping clinicians feel more secure when using AI systems in their clinical workflow.
-Future research should explore mechanisms to promote effective collaboration between clinicians, parents, AI system developers, HCI designers, and other stakeholders.
-The multi-stakeholder collaboration will be essential for designing accountable and trustworthy AI-RPM systems that can handle high-risk scenarios while ensuring patient safety and clinician confidence.
-7.3. AI as a Collaborative Decision-Making Partner For Patient Care Outside the Clinic
-Our study highlights that concussion clinicians found certain AI-supported metrics—such as physical activity, sleep patterns, and nutritional intake—useful.
-The metrics about physical activities serve as critical indicators that clinicians interpret in relation to potential mental health sequelae.
-For instance, disruptions in sleep or drastic reductions in physical activity may signal the need for closer monitoring or further evaluation of youth concussion patients.
-However, some clinicians expressed reluctance to trust AI-generated probabilistic scores unless they were accompanied by high levels of accuracy and interpretability.
-Instead, they prefer using the AI system as a data collection and consolidation tool but leave clinical decisions to human experts.
-This finding emphasizes the importance of designing AI-powered systems that do not attempt to replace clinical judgment but rather complement it by providing structured and reliable data that clinicians need to support their decision-making.
-As identified in previous CSCW studies on human-AI collaboration in high-stakes environments, trust and interpretability are key components influencing whether end-users effectively adopt AI systems in their workflows (Liao et al., 2020; Amershi et al., 2019).
-Our study explores how these principles apply in the specific context of youth concussion care, where clinicians balance multiple sources of uncertainty, patient variability, and time-sensitive decisions with respect to patients’ mental health conditions.
-Our findings resonate with ongoing discussions in the CSCW community regarding the design of collaborative systems that support human decision-making while preserving user agency.
-Existing works emphasize that human-centered AI systems should aim to act as “intelligent assistants” rather than autonomous decision-makers (Yang et al., 2019; Zhang et al., 2020).
-In particular, enabling users to maintain control over critical decisions is critical in clinical settings where over-reliance on AI-generated recommendations could lead to adverse outcomes if predictions are incorrect or poorly contextualized.
-Our work contributes new insights on how AI-RPM systems can be effectively integrated into clinical workflows without introducing additional cognitive burdens or reducing clinician autonomy.
-Specifically, our study extends the understanding of human-AI collaboration by demonstrating how AI can facilitate patient care outside the clinic through continuous monitoring of non-clinical patient data to bridge the long-existing patient care gaps between clinical visits.
-While prior HCI and CSCW research has explored AI’s role in inpatient or acute care settings (Liao et al., 2020), we shift the focus to outpatient care and remote monitoring as an emerging important healthcare area.
-Our study highlights the need for AI system design approaches that prioritize transparency, explainability, and collaboration across multiple stakeholders.
-Our findings suggest that AI-generated insights should be designed as part of an interactive process, where clinicians can not only access key metrics but also contextualize them with their broader understanding of a patient’s condition.
-For instance, future work could investigate interactive dashboards that allow clinicians to adjust or refine AI predictions based on their observations, creating a feedback loop between human expertise and machine learning outputs.
-The human-AI collaboration paradigm should not be a static and standardized format, instead, such collaboration should be seen as dynamic and iterative (Muller and Weisz, 2022; Fragiadakis et al., 2024).
-For the information presentation strategies to support clinicians’ professional workflow, our study highlights the potential for collaborative data presentation strategies that tailor information to the specific needs of different stakeholders, including clinicians, patients, and caregivers.
-As suggested by prior HCI and CSCW studies on personalized information delivery (Amershi et al., 2019), layering information hierarchically and providing context-specific views can reduce information overload while ensuring that users have access to relevant data when needed.
-In the context of AI-RPM systems, we propose to display summary statistics for routine monitoring while offering detailed visualizations or explanations for cases requiring further investigation.
-Our study not only provides practical recommendations for designing AI-RPM systems but also contributes to broader CSCW discussions on human-AI collaboration in high-stake and high-risk domains.
-Future research could explore collaborative design methodologies that involve clinicians, patients, and caregivers throughout the development process, ensuring that AI systems meet the diverse needs of all stakeholders.
-Moreover, addressing trust concerns and ensuring interpretability remain key areas for future exploration (Barnes and Hutson, 2024; Mendel et al., 2024; Prahl and Van Swol, 2017; Dietvorst et al., 2015; Wang et al., 2019).
-The CSCW community could investigate how interactive system features, such as real-time feedback or explanatory interfaces, influence user trust and decision-making in long-term monitoring scenarios.
-Additionally, as AI-RPM systems become more prevalent, interdisciplinary collaborations involving legal, ethical, and technical perspectives will be necessary to establish guidelines for responsible deployment and use.
+(2024), utilizing AI predictive models to provide actionable insights can increase the usefulness of AI in clinicians’ workflow.
+Future research could explore the design of AI-RPM technologies for collecting and distinguishing youth patients’ overlapping symptoms.
+7.2. Augmenting Objective Data With Subjective Experience
+Traditional RPM technologies, such as wearables, have already been used to remotely monitor concussion patients’ objective data (Yang et al., 2020).
+However, collecting objective data alone is not enough in our scenarios.
+As we discussed in the last section, it is critical to monitor the severity of mental health symptoms, such as anxiety and suicidal ideation.
+However, these symptoms are more subjective experiences that can hardly be captured by traditional RPM technologies.
+To address this gap, we propose using LLM-based CAs to gather patients’ anxiety levels and other mental health conditions at home.
+LLM-based CAs could automatically detect key psychological risk factors such as suicidal thoughts from patients’ conversations and generate a summary reportfor the day, which has been demoed in various prior research works in other scenarios (Bartle et al., 2023, 2022; Wang et al., 2023a; Simpson et al., 2020; Ma et al., 2024; Yang et al., 2023).
+We believe that subjective information collected from LLM-powered CAs serves not only as supplementary information but also as contextual information for objective data (i.e., heart rate).
+Prior HCI and CSCW research has highlighted the principle that technology design should incorporate context into data (Dourish, 2004; Yoo et al., 2024).
+By combining subjective information and objective data in AI-RPM systems, clinicians could use the system to assess patients’ situations more accurately and avoid unnecessary concern over what might otherwise appear to be worrisome in objective metrics alone.
+7.3. Collaborative Design in AI-RPM Systems for Enhanced Usability and Effectiveness
+When presenting objective and subjective data in an AI-RPM system, it is important to prioritize the data presentation to the usability of the system.
+Concussion clinicians are already overwhelmed by existing EHR systems, and introducing an AI-RPM system with additional data could further increase their workload and strain limited healthcare resources.
+Nevertheless, clinicians expressed a willingness to briefly review AI-RPM data before each visit to better prepare for patient consultations.
+To provide information without increasing workload, designers need to collaborate with clinicians to prioritize data based on the importance of AI-RPM systems.
+Key data should be upfront while keeping secondary information hidden but accessible.
+Our work offers a refined system design that presents a clear information hierarchy in a visually intuitive manner (Fig.
+1), serving as a reference for future development.
+Moreover, domain knowledge is essential for defining thresholds of mental health symptom severity in AI-RPM system design.
+Without clear definitions of severity, AI-RPM systems may either underreport or overemphasize symptoms, leading to delayed interventions or unnecessary concerns.
+Thus, researcher and designers should closely collaborate with clinicians to establish clear thresholds for the severity of mental health symptoms so that clinicians get alert by the system only when necessary.
+Such a collaborative approach among clinicians, researchers, and designers could enhance the usability and effectiveness of clinician-facing AI-RPM systems.
+In the future, researchers and designers should focus on the key thresholds that influence clinician decisions.
+7.4. Collaborative Emergency Management With AI-RPM Systems
+Beyond the design of AI-RPM systems, it is also important to consider how the systems support clinicians during emergency situations in real-world practice.
+In our study, concussion clinicians view AI as an important collaborative partner to support their decisions, which aligns with precious research in HCI and CSCW communities (Zhang et al., 2024; Hao et al., 2024; Yang et al., 2019; Zhang et al., 2020), However, they expressed two main concerns in emergencies.
+The first concern is the potential false alert of ”suicidal ideation” in the system.
+Youth patients may often experience mood swings and express emotions with exaggeration, which could lead to false alarms.
+As a result, a false alert could waste clinicians’ valuable time and critical medical resources that could be allocated to other patients.
+Moreover, if false alarms occur frequently, clinicians could lose trust in the system, which eventually leads to the abandonment of such a system (Liao et al., 2020; Amershi et al., 2019).
+Another concern is the ambiguity of accountability that AI-RPM systems introduce.
+When an AI-RPM system triggers an emergency alert (e.g., suicidal ideation), clinicians may face time or resource constraints that delay their response.
+Such delays can lead to serious outcomes and raise questions about accountability, which may affect system adoption.
+In high-risk, uncertain cases involving youth, it’s challenging to verify alerts and balance response with accountability.
+We believe that addressing these two challenges requires collaboration among a broader set of stakeholders and clarifying accountability (Goodman and Flaxman, 2017) before implementing the system.
+Previous HCI research has primarily focused on involving clinicians and patients in emergency management with AI systems (Wu et al., 2024; Hao et al., 2024).
+We propose involving the youth patient’s family in the decision-making process alongside clinicians to manage emergencies.
+If suicidal ideation is detected, both the clinician and the patient’s family should be notified.
+Since family members are often nearby, they can respond quickly, while clinicians provide timely professional guidance.
+Such collaboration helps mitigate risks from delayed responses while ensuring shared responsibility in emergencies.
+Future research could explore parent-facing systems that integrate with clinicians’ systems to enhance remote youth patient care.
 8. Limitations and Future Work
-This study identifies the challenges and needs of concussion clinicians when detecting mental health sequelae among youth concussion patients.
-It also presents design considerations and a proposed UI to support clinicians in this process.
-However, the study has several limitations.
-First, the sample size of our participants was limited.
-Six concussion clinicians participated in both the formative study and heuristic evaluation.
-However, recruiting these experts was challenging due to their demanding workloads (Wang et al., 2021; Jin et al., 2020) and the specialized expertise required to diagnose youth concussion patients with mental health sequelae.
-As a result, we conducted our study with highly relevant experts in this specific domain.
-Previous research has also employed a similar number of expert participants in related studies (Cai et al., 2019b; Beede et al., 2020; Jacobs et al., 2021; Yang et al., 2024; Zhang et al., 2024).
-Additionally, our recruited clinicians came from different hospitals and regions across the United States, enhancing sample diversity.
-Future research should expand the participant pool to include a broader range of clinicians across different geographic areas to improve the generalizability of the findings.
-Second, the final UI design was designed based on feedback from concussion clinicians during the heuristic evaluation.
-However, the final design has not yet been evaluated by concussion clinicians to assess its effectiveness and usability.
-Future research should incorporate participatory design methodologies to further refine the UI, ensuring it aligns more closely with clinicians’ workflows and needs.
-Third, our study focused on UI design rather than developing an interactive prototype.
-The primary goal of our UI design was to explore the feasibility and provide insights for further design refinements before actual development, thereby saving resources and time for both researchers and clinicians.
+Our work is not without limitations.
+First, the sample size of our participants was limited, with six concussion clinicians participating in both the formative and evaluation studies.
+Recruiting these experts was particularly challenging due to their demanding workloads (Wang et al., 2021; Jin et al., 2020) and the specialized expertise required to treat youth concussion patients with mental health issues As a result, we recruited six highly relevant experts in this specific domain.
+Despite the limited number, our inductive thematic analysis reached thematic saturation.
+Moreover, previous research employed a similar number of expert participants in related studies (Cai et al., 2019b; Beede et al., 2020; Jacobs et al., 2021; Yang et al., 2024; Zhang et al., 2024), which supports the appropriateness of our sample size.
+Future research should include a broader range of concussion clinicians across locations for greater generalizability of the findings.
+Secondly, the refined AI-RPM system has not yet been further evaluated by concussion clinicians to assess its effectiveness and usability.
+In future research, we plan to incorporate participatory design methodologies (Muller and Kuhn, 1993) to revise the system design, ensuring it aligns more closely with clinicians’ workflows and needs.
+Moreover, our study focused on system design rather than developing an interactive prototype.
+The primary goal of our system design was to explore the feasibility and provide insights for further design refinements before actual development, thereby saving resources and time for both researchers and clinicians.
 Future research should investigate clinicians’ pain points and needs when interacting with a functional system developed.
-Finally, our study scenario specifically examined the detection of mental health sequelae after concussions among youth and explored the potential of AI-RPM in this context.
-Future studies could explore the broader application of AI-RPM in various clinical settings and understand how AI-driven systems can assist clinicians in early detection and intervention across different medical domains.
+Finally, we want to highlight the potential of applying AI-RPM systems in varied healthcare settings.
+Several studies have explored remote monitoring in the healthcare domain.
+Wyche et al.
+(2024) utilized mHealth to track the health conditions of type 1 diabetes among youth.
+Seals et al.
+(2022) leveraged wearables to detect gait impairment, monitor patients’ responses to treatment, and visualize patient data.
+We believe that AI-RPM technologies have the potential to be leveraged and extended to these scenarios.
+Future work should expand AI-RPM technologies to diverse healthcare settings, enabling timely interventions and improving care quality.
 9. Conclusion
-We explored the challenges and needs faced by concussion clinicians related to mental health issues in youth concussion patients and proposed design considerations for the use of AI-RPM technologies in this high-risk, high-stakes context.
-Our study makes two key contributions to the CSCW field.
-First, we identified three core challenges clinicians encounter when collecting and monitoring mental health-related information and making mental health-related decisions: (1) collecting patients’ mental health information outside the clinic; (2) communicating effectively with patients and their family members; and (3) assessing patients’ compliance with clinical recommendations.
-Second, we proposed a series of opportunities for AI-RPM technologies to address these challenges, as well as design considerations for AI-powered systems to collect multimodal, non-clinical data that supports clinicians in making informed mental health-related decisions.
-To our knowledge, this is the first study in the CSCW field to focus on the intersection of concussion clinicians and youth concussion patients in the context of mental health-related decision-making, providing a foundation for designing AI-RPM systems in similar scenarios.
+We conducted a formative study with six concussion clinicians to understand the challenges and needs during their clinical practice with youth concussion patients, particularly with respect to the patient mental health sequelae.
+Then, we derived a suite of design considerations with the use of AI-RPM technologies.
+Here are three design considerations.
+Firstly, we recommend the use of wearables to remotely monitor patients’ sleep and physical activity data, helping assess treatment compliance and inform mental health-related decisions.
+Secondly, we propose LLM-powered CAs to enhance mental health self-reporting, which facilitates private, natural interactions with patients and summarizes key information for clinicians.
+Thirdly, we suggest leveraging AI risk prediction in detecting concealed or worsening mental health sequelae.
+Finally we delivered a clinician-facing system design as our final artifact.
+To our knowledge, this is the first study in the CSCW field to focus on the intersection of health, RPM, and human-AI decision-making, which provides a foundation for designing AI-RPM systems in varies medical scenarios.
 Appendix A APPENDIX: INTERVIEW SCRIPTS FOR FORMATIVE STUDY
@@
 Technology (Expectations and Concerns)
-(1) Do you use any existing tools or features in EHR (e.g., EPIC) in the youth patients’ mental health decision-making? If so, what are they and how do you like them?
-Appendix B APPENDIX: INTERVIEW SCRIPTS FOR DESIGN INTERFACE EVALUATION
+(1) Do you use any existing tools or features in EHR (e.g., EPIC) in the youth patients’ mental health decision-making? If so, what are they, and how do you like them?
+Appendix B APPENDIX: INTERVIEW SCRIPTS FOR EVALUATION STUDY
 (1) This section (Fig.1) shows the probability of mental health sequelae within the next 15 days with a range from 0% to 100%, where the higher probability score signifies a higher probability of experiencing mental health sequelae.